### v1

# Generalized cryptographic semi-regular sequences: A variant of Fröberg conjecture and a simple complexity estimation for Gröbner basis computation

Abstract
In this paper, we study generalized cryptographic semi-regular sequences, which are expected to generic in the space of homogeneous polynomial sequences on which the coordinate rings have Krull dimension one. We provide an upper-bound on the complexity of the Gröbner basis computation for such a sequence.
1 Introduction
In the theory of polynomial rings over fields, Gröbner bases introduced first by Buchberger [8] provide various methods for investigating important properties of an ideal and its associated algebraic variety. Here, a Gröbner basis is defined as a special kind of its generator set, relative to a monomial ordering. Buchberger also presented the first algorithm (Buchberger’s algorithm) for computing Gröbner bases. Lazard’s formulation in [30] and [31] relates the problem of computing Gröbner bases to the problem of reducing so-called Macaulay matrices. FGLM [19] and Gröbner walk [12] were propsed as algorithms that efficiently convert a given Gröbner basis to a Gröbner basis with respect to another monomial ordering; Hilbert driven [37] can be also applied to such a conversion. Faugère’s proposed his innovative algorithms $F_{4}$ [17] and $F_{5}$ [18] allowing one to quite efficiently compute Gröbner bases, and these are nowadays state-of-the-art algorithms. We refer to [16] for a survey on signature-based algorithms including $F_{5}$. Thanks to Gröbner basis conversion algorithms, it is important to estimate the complexity of computing a Gröbner basis with respect to a specific monomial ordering. In this paper, we consider only the case of graded monomial orderings.
In order to estimate the complexity, it suffices to provide an upper-bound on the so-called solving degree: Roughly speaking, the solving degree is the highest degree of polynomials involved in Gröbner basis computation, see [29, Section 2.2] for rigorous definitions. If the ideal is homogeneous, the solving degree can be equal to the maximal degree of the reduced Gröbner basis with respect to a monomial ordering that one considers, and its upper-bounds have been well-studied. In the case of inhomogeneous ideals, the solving degree can be greater than the maximal degree of the reduced Gröbner basis, whence estimating it directly is difficult in general, see [10] and [29] for details. On the other hand, it is well-known (cf. [27, Prop. 4.3.18]) that a Gröbner basis of an inhomogeneous ideal with respect to a graded monomial ordering is easily obtained from a Gröbner basis of the ideal of homogenized generators. Therefore, the homogeneous case is very important, and we focus on that case. In the following, let $K$ be a field, and let $R=K[x_{1},\ldots,x_{n}]$ (resp. $R^{\prime}=K[x_{0},x_{1},\ldots,x_{n}]$) be the polynomial ring of $n$ (resp. $n+1$) variables over $K$. For a possibly inhomogeneous ideal $I$ of $R$ (resp. $R^{\prime}$), we denote by $\mathrm{max.GB.deg}_{\prec}(I)$ the maximal degree of the elements of the reduced Gröbner basis of $I$ with respect to a monomial ordering $\prec$ on the set of monomials in $R$ (resp. $R^{\prime}$), following [9, Definition 7] (it is also denoted by $\deg(I,\prec)$ in the literature of commutative algebra, see e.g., [24, Definition 4]). If $I$ is generated by $F\subset R$, we may write $\mathrm{max.GB.deg}_{\prec}(F)$ in place of it.
For a homogeneous ideal $I\subset R=K[x_{1},\ldots,x_{n}]$ with a generating set $F=\{f_{1},\ldots,f_{m}\}\subset R$, it is well-known (cf. [2, Proposition 1]) that a Gröbner basis of $I$ with respect to a graded monomial ordering up to degree $d$ can be computed in $O(md\binom{n+d-1}{d}^{\omega})$, where $32\leq\omega<$ 3 is the exponent of matrix multiplication. (We can also obtain another upper-bound $O(m\binom{n+d}{d}^{\omega})$ by the proof of [35, Theorem 1.72].) With this fact, we have in the homogeneous case that the complexity of computing the reduced Gröbner basis of $I$ is upper-bounded by $O(mD\binom{n+D-1}{D}^{\omega})$ (or $O(m\binom{n+D}{D}^{\omega})$) for any integer $D$ with $\mathrm{max.GB.deg}_{\prec}(I)\leq D$. Hence, finding a tighter bound $D$ of $\mathrm{max.GB.deg}_{\prec}(I)$ is important for estimating the complexity of Gröbner basis computation in the homogeneous case. The case on which we focus on is the case where the Krull dimension $\mathrm{Krull.dim}(R/I)$ is equal to or less than one; see e.g., [24, Section 1] for a summary of existing results when $\mathrm{Krull.dim}R/I\geq 2$. In the case where $\mathrm{Krull.dim}(R/I)\leq 1$, Lazard’s bound (called also Macaulay’s bound) is well-known, and it is given by $\sum_{j=1}^{\ell}(d_{j}-1)+1$ with $d_{1}\geq d_{2}\geq\cdots\geq d_{m}$ and $\ell:=\max\{m,n\}$, where we set $d_{j}=\deg(f_{j})$ for $1\leq j\leq m$. When $I$ is in quasi stable position (equivalently $R/\langle I,x_{n}\rangle$ is Artinian, see Lemma 2.1.8 below), we can apply Hashemi-Seiler’s inequality
$\mathrm{max.GB.deg}_{\prec}(I)\leq\mathrm{max}\{\mathrm{hilb}(\langle I,x_{n}\rangle),\mathrm{hilb}(I)\}$ (1.0.1)
for the graded reverse lexicographic ordering $\prec$ with $x_{n}\prec\cdots\prec x_{1}$, where $\mathrm{hilb}(J)$ denotes the Hilbert regularity of a homogeneous ideal $J$ of $R$. Note that $\mathrm{hilb}(\langle I,x_{n}\rangle)$ is equal to the degree of regularity ${d}_{\rm reg}(\langle I,x_{n}\rangle)$ since $R/\langle I,x_{n}\rangle$ is Artinian, and $\mathrm{hilb}(I)$ is equal to the generalized degree of regularity $\widetilde{d}_{\rm reg}(I)$. For the definition of $d_{\rm reg}$ and $d_{\rm reg}$, see Remark 2.1.2 below.
On the other hand, in [29], when $F=\{f_{1},\ldots,f_{m}\}\subset R$ with $m\geq n$ is obtained by homogenizing an affine semi-regular sequence in $K[x_{1},\ldots,x_{n-1}]^{m}$, we improved Lazard’s bound on $\mathrm{max.GB.deg}_{\prec}(\langle F\rangle_{R})$ to $\sum_{j=1}^{n}(d_{j}-1)+1$ with $d_{1}\leq\cdots\leq d_{m}$. Under an additional assumption that $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular (see Definition 2.2.7), we also proved in [29] that the right hand side of (1.0.1) is upper-bounded by the degree of regularity for a cryptographic semi-regular sequence of $m$ polynomials of degrees $d_{1},\ldots,d_{m}$ in $n+1$ variables.
In this paper, we shall prove that, if $K$ is large enough, the inequality (1.0.1) always holds after a linear coordinate change. Then, we also estimate the complexity of computing a Gröbner basis for $I$, assuming that $I$ is generated by a generalized cryptographic semi-regular sequence. Such a sequence is expected to be generic (see Definition 2.3.1 for definition), analogously to Fröberg’s conjecture [20]. Our results are summarized into the following theorem and proposition:
Theorem 1.
Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$ (in any characteristic), and let $I$ be a proper homogeneous ideal of $R$. Let $F=\{f_{1},\ldots,f_{m}\}$ be a subset of $R$ generating $I$, and put $d_{j}:=\deg(f_{j})$ (total degree) for $1\leq j\leq m$. Assume that $I$ has finitely many projective zeros over $\overline{K}$, and suppose that the cardinality of $K$ is greater than the number of projective zeros over $\overline{K}$ of $I$. Choose an arbitrary linear form $\ell$ in $R$ such that it does not vanish at any projective zero over $\overline{K}$ of $I$ (such an $\ell$ always exists by our assumption $\mathrm{Krull.dim}(R/I)\leq 1$). Then, the following (1) holds:
  1. (1)
    There exists a linear transformation $\sigma$ of variables $x_{1},\ldots,x_{n}$ over $K$ sending $\ell$ to $x_{n}$ such that
    ${\rm max.GB.deg}_{\prec}(I^{\sigma})\leq\mathrm{max}\{d_{\rm reg}(\langle I,\ell\rangle_{R}),\widetilde{d}_{\rm reg}(I)\}$ (1.0.2)
    for the graded reverse lexicographic ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$, where $I^{\sigma}$ is the image of $I$ by $\sigma$.
Moreover, if $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular, then following (2) and (3) also hold:
  1. (2)
    We have
    $\mathrm{max}\{d_{\rm reg}(\langle I,\ell\rangle_{R}),\widetilde{d}_{\rm reg}(I)\}\leq D^{(n,m)},$ (1.0.3)
    where
    $D^{(n,m)}:=\left\{\!\begin{array}[]{cc}\displaystyle{\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}\right]\right)}+1&(m\geq n),\\ \displaystyle\sum_{i=1}^{n-1}(d_{i}-1)+1&(m=n-1).\end{array}\right.$ (1.0.4)
  2. (3)
    The complexity of computing a Gröbner basis of $I^{\sigma}$ with respect to $\prec$ is upper-bounded by
    $O\left(m\binom{n+D^{(n,m)}-1}{D^{(n,m)}}^{\omega}\right),$ (1.0.5)
    where $32\leq\omega<$ 3 is the exponent of matrix multiplication.
Proposition 1.
Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$ (in any characteristic), and let $\prec$ be the graded reverse lexicographic ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$. Let $I$ be a proper homogeneous ideal of $R$. Assume that the Krull dimension of $R/I$ is $\leq 1$ and $R/\langle I,x_{n}\rangle$ is Artinian. If $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic, then the inequality (1.0.1) becomes an equality. Hence, the inequality (1.0.2) in Theorem 1 (2) also becomes an equality.
Theorem 1 can be viewed as a generalization of [29, Proposition 1 and Corollary 1]. The bound $D_{\rm new}$ given in (1.0.4) coincides with Lazard’s bound when $m\leq n$, while it can be sharper than Lazard’s one when $m>n$. We also note that, the complexity bound (1.0.5) is better than the usual bound $O(mD\binom{n+D-1}{D}^{\omega})$ (see e.g., [2, Proposition 1], [36, Equation (24)]) and than $O(m\binom{n+D}{D}^{\omega})$ (which comes from the proof of [35, Theorem 1.72]) for an arbitrary $D$ with $D\geq\mathrm{max.GB.deg}_{\prec}(F)$.
Our assumption that a homogeneous polynomial sequence is generalized cryptographic semi-regular could be useful to analyze the security of multivariate cryptosystems and so on (cf. [29, Section 4.3]).
Remark 1 (Inhomogeneous case).
By Theorem 1, we can also estimate the complexity of computing a Gröbner basis for an inhomogeneous polynomial sequence defining a zero-dimensional ideal, as in [29, Section 4.3]. Indeed, for a subset $F=\{f_{1},\ldots,f_{m}\}$ of $R=K[x_{1},\ldots,x_{n}]$ containing at least one inhomogeneous polynomial, a Gröbner basis of $\langle F\rangle_{R}$ is obtained by reducing the associated degree-$D$ Macaulay matrix at $D=\mathrm{max.GB.deg}_{\prec^{h}}(I)$ with $I=\langle f_{1}^{h},\ldots,f_{m}^{h}\rangle\subset R^{\prime}:=R[y]$, see [9, Theorem 7]. Here, each $f_{j}^{h}$ denotes the homogenization of $f_{j}$ by an extra variable $y$, and $\prec^{h}$ is the homogenization of $\prec$, see e.g., [27, Section 3.1.1] for definition. The complexity of reducing the Macaulay matrix can be easily estimated as $O(m\binom{n+D}{D}^{\omega})$, without assuming $D=O(n)$. Suppose the following three conditions:
  • The Krull dimension of $R^{\prime}/\langle f_{1}^{h},\ldots,f_{m}^{h},y\rangle$ is zero, namely the Krull dimension of $R/\langle f_{1}^{\rm top},\ldots,f_{m}^{\rm top}\rangle$ is zero, where $f_{j}^{\rm top}:=(f_{j}^{h})|_{y=0}$ for $1\leq j\leq m$. (Equivalently these two graded rings are both Artinian.)
  • The sequence $(f_{1}^{h},\ldots,f_{m}^{h},y)\in(R^{\prime})^{m+1}$ is cryptographic semi-regular, equivalently $(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})\in R^{m}$ is cryptographic semi-regular.
  • The sequence $(f_{1}^{h},\ldots,f_{m}^{h})\in(R^{\prime})^{m+1}$ is generalized cryptographic semi-regular.
Then, by the first assumption, it follows from Theorem 1 (1) that the inequality (1.0.2) holds for $I=\langle F^{h}\rangle_{R^{\prime}}$ and $\ell=y$, where we replace $R$ by $R^{\prime}$ and set $\sigma$ to be the identity. Moreover, under the second and third assumptions, we also have the complexity bound
$O\left(m\binom{n+D^{(n+1,m)}}{D^{(n+1,m)}}^{\omega}\right)$ (1.0.6)
by Theorem 1 (2). Note that this upper-bound is heuristically found in [21].
2 Preliminaries
Let $K$ be a field, and let $\overline{K}$ denote its algebraic closure. Let $n$ be a natural number, and let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over $K$. For $d\geq 0$, we denote by $R_{d}$ the homogeneous component of $R$ of degree $d$. The Krull dimension of a commutative ring $A$ is denoted by $\mathrm{Krull.dim}(A)$. For ideals $I$ and $J$ of $R$, their ideal quotient is defined as $(I:J)=\{f\in R\mid fJ\subset I\}$. The saturation of $I$ with respect to $J$ is $(I:J^{\infty}):=\{f\in R\mid fJ^{s}\subset I\mbox{ for some $s\geq 1$}\}$. In particular, the saturation of a homogeneous ideal $I$ with respect to the maximal homogeneous ideal $\mathfrak{m}=\langle x_{1},\ldots,x_{n}\rangle_{R}$ is denoted by $I^{\rm sat}=(I:\mathfrak{m}^{\infty})$, which is also homogeneous.
2.1 Hilbert regularity
Let $I$ be a homogeneous ideal of $R$. Let ${\rm HP}_{R/I}\in\mathbb{Q}[z]$ denote the Hilbert polynomial of $R/I$, namely there exists a non-negative integer $d_{0}$ such that ${\rm HF}_{R/I}(d)={\rm HP}_{R/I}(d)$ for any $d$ with $d\geq d_{0}$, where $\mathrm{HF}$ is the Hilbert function.
Definition 2.1.1 (Hilbert regularity (or index of regularity [13, Chapter IX, Section 3])).
The smallest $d_{0}$ such that ${\rm HF}_{R/I}(d)={\rm HP}_{R/I}(d)$ for any $d$ with $d\geq d_{0}$ is called the Hilbert regularity of $I$, denoted by ${\rm hilb}(I)$, or is called the index of regularity of $I$, denoted by $i_{\rm reg}(I)$.
Throughout the rest of this subsection, we set $r:=\mathrm{Krull.dim}(R/I)$.
Remark 2.1.2.
It is well-known that the following conditions are all equivalent:
  1. (1)
    The Krull dimension $r$ of $R/I$ is equal to $1$ (resp. 00).
  2. (2)
    The projective algebraic set $V_{\overline{K}}(I)$ is a finite set (resp. an empty set).
  3. (3)
    The Hilbert polynomial ${\rm HP}_{R/I}$ is a positive constant (resp. the zero).
  4. (4)
    There exists a non-negative integer $d_{0}$ such that ${\rm HF}_{R/I}(d)={\rm HF}_{R/I}(d_{0})$ (resp. ${\rm HF}_{R/I}(d)=0$) for any $d$ with $d\geq d_{0}$.
In this case, it is straightforward that
$\mathrm{hilb}(I)$ $=$ $\min\{d_{0}:{\rm HF}_{R/I}(d)={\rm HF}_{R/I}(d_{0})\mbox{ for any$d$ with $d\geq d_{0}$}\}$
$=$ $\min\{d_{0}:\dim_{K}(R/I)_{d}=\dim_{K}(R/I)_{d_{0}}\mbox{ for any$d$ with $d\geq d_{0}$}\}.$
The middle and the right hand sides are called the generalized degree of regularity of $I$ for $r\leq 1$, denoted by $\widetilde{d}_{\rm reg}(I)$ (cf. [29, Sect. 2.3]): When $r=0$, these are equal to the degree of regularity $d_{\rm reg}(I):=\min\{d:R_{d}=I_{d}\}$. Namely, we have
$\mathrm{hilb}(I)=\left\{\begin{array}[]{cll}\widetilde{d}_{\rm reg}(I)&&(r=1),\\ \widetilde{d}_{\rm reg}(I)&=d_{\rm reg}(I)&(r=0).\end{array}\right.$ (2.1.1)
See [7] for bounds on ${d}_{\rm reg}(I)$.
Although we mainly treat the case where $r\leq 1$ in this paper, we distinguish $\mathrm{hil}(I)$ and $\widetilde{d}_{\rm reg}(I)$ for the compatibility with the condition that a sequence of homogeneous polynomials in $R$ is regular (cf. Lemma 2.2.9 below).
Remark 2.1.3.
We also note that ${\rm HS}_{R/I}$ is written as
${\rm HS}_{R/I}(z)=\dfrac{h_{R/I}(z)}{(1-z)^{r}}$
for some polynomial $h_{R/I}(z)\in\mathbb{Z}[z]$, which is called the $h$-polynomial of $R/I$. Then, we have ${\rm hilb}(I)=\deg h_{R/I}-r+1$, see [6, Prop. 4.1.12].
Moreover, it follows from Grothendieck-Serre’s formula that
${\rm hilb}(I)\leq\mathrm{reg}(I)=\mathrm{reg}(R/I)+1,$
where ‘${\rm reg}$’ means Castelnuovo-Mumford regularity, see [15, §20.5] for the definition. The equality holds if $R/I$ is Artinian, i.e., the Krull dimension of $R/I$ is zero.
Definition 2.1.4 (cf. [24, Section 3]).
We say that $I$ is in Noether position with respect to the variables $x_{1},\ldots,x_{n-r}$, where $r=\mathrm{Krull.dim}(R/I)$, if the injective ring extension $K[x_{n-r+1},\ldots,x_{n}]\to R/I$ is integral.
Lemma 2.1.5 ([5, Lemma 4.1]; see also [32, Proposition 5.4]).
Let $I$ be a homogeneous ideal of $R$, and let $\prec$ be the reverse lexicographic order on $R$ with $x_{n}\prec\cdots\prec x_{1}$. Then, the following are equivalent:
  1. (1)
    $I$ is in Noether position with respect to the variables $x_{1},\ldots,x_{n-r}$.
  2. (2)
    For each $i$ with $1\leq i\leq n-r$, there exists $r_{i}\geq 0$ such that $x_{i}^{r_{i}}\in\langle\mathrm{LM}_{\prec}(I)\rangle$.
  3. (3)
    $\mathrm{Krull.dim}(R/\langle I,x_{n-r+1},\ldots,x_{n}\rangle)=0$.
  4. (4)
    $\mathrm{Krull.dim}(R/\langle\langle\mathrm{LM}_{\prec}(I)\rangle,x_{n-r+1},\ldots,x_{n}\rangle)=0$.
In the proof of [5, Lemma 4.1], the authors first proved the equivalence of the conditions (2), (3), and (4), and the implication (1) $\Rightarrow$ (2). The non-trivial part is that (2), (3), and (4) imply (1), which follows from [32, Proposition 5.4].
Here, we also recall the definition introduced by Hashemi [23] that a homogeneous ideal is in strong Noether position.
Definition 2.1.6 ([23, Theorem 2.19]).
A homogeneous ideal $I$ of $R$ is said to be in strong Noether position (SNP) with respect to the ordered variables $x_{1},\ldots,x_{n}$ if we have $I^{\rm sat}=(I:x_{n}^{\infty})$ and
$\langle I,x_{n-i+1},\ldots,x_{n}\rangle^{\rm sat}=(\langle I,x_{n-i+1},\ldots,x_{n}\rangle:x_{n-i}^{\infty})$
for any $i$ with $11\leq i\leq r-$ 1.
For equivalent definitions, see [23, Section 2] for details. Note that $I$ is in SNP if and only if $\langle\mathrm{LM}_{\prec}(I)\rangle$ is in SNP. Hashemi also proved the following inequality:
Theorem 2.1.7 ([23, Proposition 4.7 and Theorem 4.17]).
If $I$ is in SNP, then $\mathrm{reg}(I)$ is equal to
$\overline{\rm hilb}(I):=\max\{\mathrm{hilb}(I)\}\cup\{\mathrm{hilb}(\langle I,x_{n-i+1},\ldots,x_{n}\rangle)\mid 1\leq i\leq{\rm Krull.dim}(R/I)\},$
which is called the stabilized Hilbert regularity of $I$, and moreover
$\mathrm{max.GB.deg}_{\prec}(I)\leq\overline{\rm hilb}(I)$
for the graded lexicographic ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$.
Lemma 2.1.8 ([24, Lemma 23] and [25, Proposition 3.2 (v)]).
If $r=1$, then the following are all equivalent:
  1. (1)
    $I$ is in Noether position with respect to $x_{1},\ldots,x_{n-1}$ (or one of the equivalent conditions given in Lemma 2.1.5 holds).
  2. (2)
    $I$ is in quasi stable position with respect to the ordered variables $x_{1},\ldots,x_{n}$ (see e.g., [24, Definition 6] for definition).
  3. (3)
    $I$ is in strong Noether position.
Proof.
The equivalence of (1) and (2) is proved in [24, Lemma 23], and that of (2) and (3) follows from [25, Proposition 3.2 (v)]. ∎
By Theorem 2.1.7 and Lemma 2.1.8, we obtain the following theorem:
Theorem 2.1.9 ([24, Proof of Theorem 30]).
Assume that $r=1$. If $R/\langle I,x_{n}\rangle$ is Artinian (or one of the equivalent conditions given in Lemma 2.1.8 holds), then we have the inequality (1.0.1) for the graded reverse lexicographic ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$.
2.2 Generalized semi-regular sequences
In this subsection, we recall the definition of cryptographic semi-regular sequences and that of generalized ones. Throughout this subsection, let $f_{1},\ldots,f_{m}\in R$ be homogeneous polynomials of positive degrees $d_{1},\ldots,d_{m}$, and put $I=\langle f_{1},\ldots,f_{m}\rangle$.
Theorem 2.2.1 (cf. [14, Theorem 1]).
For each $d$ with $d\geq\mathrm{max}\{d_{i}:1\leq i\leq m\}$, the following are equivalent:
  1. (1)
    For each $i$ with $1\leq i\leq m$, if a homogeneous polynomial $g\in R$ satisfies $gf_{i}\in\langle f_{1},\ldots,f_{i-1}\rangle_{R}$ and $\mathrm{deg}(gf_{i})<d$, then we have $g\in\langle f_{1},\ldots,f_{i-1}\rangle_{R}$, namely, the multiplication-by-$f_{i}$ map $(R/\langle f_{1},\ldots,f_{i-1}\rangle)_{t-d_{i}}\longrightarrow(R/\langle f_{1},\ldots,f_{i-1}\rangle)_{t}$ is injective for any $t$ with $d_{i}\leq t<d$.
  2. (2)
    We have
    ${\rm HS}_{R/I}(z)\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{d}},$ (2.2.1)
    where $\mathrm{HS}$ denotes the Hilbert series.
  3. (3)
    We have $H_{1}(K_{\bullet}(f_{1},\ldots,f_{m}))_{\leq d-1}=0$, where $K_{\bullet}(f_{1},\ldots,f_{m})$ denotes the Koszul complex on the sequence $(f_{1},\ldots,f_{m})$.
Definition 2.2.2 ([1, Definition 3]; see also [14, Definition 1]).
For each $d$ with $d\geq\mathrm{max}\{d_{i}:1\leq i\leq m\}$, a sequence $(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R\smallsetminus K$ is said to be $d$-regular if the equivalent conditions in Theorem 2.2.1 hold.
Here, we define a cryptographic semi-regular sequence:
Definition 2.2.3 ([1, Definition 5], [3, Definition 5]; see also [14, Section 2]).
A sequence $(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R\smallsetminus K$ is said to be cryptographic semi-regular if it is $d_{\rm reg}(I)$-regular.
The Hilbert series of $R/I$ for a cryptographic semi-regular sequence $(f_{1},\ldots,f_{m})$ is easily computed, by the following Proposition 2.2.4:
Proposition 2.2.4 ([14, Proposition 1 (d)]; see also [3, Proposition 6]).
With the same notation as in Theorem 2.2.1, a sequence $(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R\smallsetminus K$ is cryptograpic semi-regular if and only if
${\rm HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}(z)=\left[\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\right],$ (2.2.2)
where []delimited-[][$\cdot$][ ⋅ ] means truncating a formal power series over $\mathbb{Z}$ after the last consecutive positive coefficient. (Note that, in this case, we have $d_{\rm reg}(I)=\deg({\rm HS}_{R/I}(z))+1$.)
The above ‘standard’ form (2.2.2) of Hilbert series was introduced by Fröberg in his celebrated paper [20]. We will recall his famous conjecture (Conjecture 2.3.3 below) in the next subsection.
Here, we recall the following lemma:
Lemma 2.2.5 ([20, b) on page 121]).
With notation as above, we have
${\rm HS}_{R/I}(z)\geq^{\rm lex}\left[\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\right]$
in the lexicographic sense, i.e., the first coefficient that differs, if any, is greater on the left hand side, see [20, page 118].
Remark 2.2.6.
For any non-negative integer $d$, it follows from Lemma 2.2.5 that
$\left({\rm HS}_{R/I}(z)\bmod{z^{d}}\right)\geq^{\rm lex}\left(\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\bmod{z^{d}}\right).$
To deal with the case where $R/I$ has Krull dimension one, we extend the notion of cryptgraphic semi-regular as follows:
Definition 2.2.7 ([29, Definition 2.3.3]).
A sequence $(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R\smallsetminus K$ is said to be generalized cryptographic semi-regular if it is $\widetilde{d}_{\rm reg}(I)$-regular.
Lemma 2.2.8.
If the sequence $\bm{F}=(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular, then we have
$\widetilde{d}_{\rm reg}(I)\leq{\deg\left(\left[\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\right]\right)}+1.$ (2.2.3)
Proof.
By our assumption that $\bm{F}$ is generalized cryptographic semi-regular, the equality (2.2.1) holds for $d=\widetilde{d}_{\rm reg}(I)$. For any $d^{\prime}$ with $d^{\prime}<\widetilde{d}_{\rm reg}(I)$, we have ${\rm HF}_{R/I}(d^{\prime})>0$, whence the $z^{d^{\prime}}$-coefficient of $\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}$ is also positive. Thus, the inequality (2.2.3) holds. ∎
For a sequence $\bm{F}=(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R\smallsetminus K$, we consider the following four conditions:
$\bm{F}$ is regular, equivalently
${\rm HS}_{R/I}(z)=\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}.$
$\bm{F}$ is cryptographic semi-regular (i.e., $d_{\rm reg}(I)$-regular).
$\bm{F}$ is generalized cryptographic semi-regular (i.e., $\widetilde{d}_{\rm reg}(I)$-regular).
$\bm{F}$ is $\mathrm{hilb}(I)$-regular.
Lemma 2.2.9.
For conditions (A), (B), (C), and (D) as defined above, there is an implication:
$(A)\Rightarrow(B)\Rightarrow(C)\Rightarrow(D),\mbox{ but }(D)\overset{r\geq 2}{\not\Rightarrow}(C)\overset{r=1}{\not\Rightarrow}(B)\overset{m>n}{\not\Rightarrow}(A).$
Namely, (A) and (B) are equivalent for $m\leq n$, (B) and (C) are equivalent for $r\neq 1$, (C) and (D) are equivalent for $r\leq 1$, and (A), (B), (C), and (D) are all equivalent for $m=n$ and $r=0$.
Proof.
It is clear that $(A)\Rightarrow(B)\Rightarrow(C)\Rightarrow(D)$. When $r\leq 1$, it follows from (2.1.1) that $(D)\Rightarrow(C)$ holds. Also when $r\neq 1$, we have $d_{\rm reg}(I)=\widetilde{d}_{\rm reg}(I)$, whence $(C)\Rightarrow(B)$ holds. For $m\leq n$, Proposition 2.2.4 implies $(B)\Rightarrow(C)$. ∎
2.3 Genericness and a variant of Fröberg’s conjecture
Given $m$ non-negative integers $d_{1},\ldots,d_{m}$, let $V(n,m,d_{1},\ldots,d_{m})\!=\!R_{d_{1}}\times\cdots\times R_{d_{m}}$, which is topologically identified with the affine space $\mathbb{A}^{N_{1}}(K)\times\dots\times\mathbb{A}^{N_{m}}(K)$ of dimension $\sum_{j=1}^{m}N_{j}$ with $N_{j}:=\binom{n+d_{j}-1}{d_{j}}$. Here the identification from $R_{d}$ to $\mathbb{A}^{N}$, where $N=\binom{n+d-1}{d}$, is given as follows: Letting $\{t_{1},\ldots,t_{N}\}$ be the set of all monomials in $R$ of degree $d$, each polynomial $f=\sum_{i=1}^{N}c_{i}t_{i}$ in $R_{d}$ is identified with its coefficients vector $(c_{1},\ldots,c_{N})$. We write simply $V$ for $V(n,m,d_{1},\ldots,d_{m})$ when we fix the parameter $(n,m,d_{1},\ldots,d_{m})$.
Definition 2.3.1 (cf. [34, Section 1]).
For given $n$, $m$, and $(d_{1},\ldots,d_{m})$, a property $\mathcal{P}$ for $\bm{F}=(f_{1},\ldots,f_{m})\in V$ is said to be generic (or, we say that a generic sequence in $V$ satisfies $\mathcal{P}$) if it holds on a non-empty Zariski open set in $V$, namely there exists a non-empty Zariski open set $U$ in $V$ such that
$U\subset\{\bm{F}=(f_{1},\ldots,f_{m})\in V:\mbox{$\bm{F}$ satisfies $\mathcal{P}$}\}.$
Remark 2.3.2.
Another definition of genericness is the following (cf. [20, page 119]): For given $n$, $m$, and $(d_{1},\ldots,d_{m})$, a property $\mathcal{P}$ for $\bm{F}=(f_{1},\ldots,f_{m})\in V$ is said to be generic over $K$ if it holds for any $\bm{F}=(f_{1},\ldots,f_{m})\in V$ such that the coefficients of $f_{1},\ldots,f_{m}$ are algebraically independent over the prime field $L$ of $K$, that is, $L=\mathbb{Q}$ or $\mathbb{F}_{p}$ for some prime number $p$. It can be checked that the two definitions of genericness are equivalent.
Here, we recall Fröberg’s conjecture:
Conjecture 2.3.3 (Fröberg’s conjecture [20]).
Assume that $K$ is an infinite field. For given $n$, $m$, and $(d_{1},\ldots,d_{m})$, the property that $\bm{F}=(f_{1},\ldots,f_{m})\in V$ satisfies (2.2.2) (i.e., $\bm{F}$ is cryptographic semi-regular) is generic. In other words, a generic sequence $\bm{F}=(f_{1},\ldots,f_{m})\in V$ is cryptographic semi-regular.
See Pardue’s paper [34] for some conjectures equivalent to Fröberg’s one (Conjecture 2.3.3). He also proved that Moreno-Socías conjecture [33] is stronger than Conjecture 2.3.3, see [34, Theorem 2] for a proof.
Here, as a variant of Conjecture 2.3.3, we raise the following conjecture (cf. [29, Conjecture 4.3.4] as its strict version):
Conjecture 2.3.4.
Assume that $K$ is an infinite field. For given $n$, $m$, and $(d_{1},\ldots,d_{m})$, the property that $\bm{F}=(f_{1},\ldots,f_{m})\in V$ is generalized cryptographic semi-regular is generic.
We shall prove Conjectures 2.3.3 and 2.3.4 are equivalent to each other. For this, we first prove the following lemma:
Lemma 2.3.5.
For $m\geq n$, the property that $R/\langle f_{1},\ldots,f_{m}\rangle$ is Artinian is generic.
Proof.
The case $m=n$ is well-known to hold. Assume $m>n$, fix $n$, $m$, and $(d_{1},\ldots,d_{m})$, and put
$\Lambda^{\rm Art}:=\{\bm{F}=(f_{1},\ldots,f_{m})\in V:\mbox{$R/\langle f_{1},\ldots,f_{m}\rangle$ is Artinian}\}.$
Consider a natural projection
$V\to V_{n}:=\mathbb{A}^{N_{1}}(K)\times\dots\times\mathbb{A}^{N_{n}}(K)\ ;\ (f_{1},\ldots,f_{n},f_{n+1},\ldots,f_{m})\mapsto(f_{1},\ldots,f_{n}).$
Then we have $\pi^{-1}(\Lambda_{n}^{\rm Art})\subset\Lambda^{\rm Art}$, where
$\Lambda_{n}^{\rm Art}:=\{\bm{F}=(f_{1},\ldots,f_{n})\in V:\mbox{$R/\langle f_{1},\ldots,f_{n}\rangle$ is Artinian}\}$
is Zariski-open in $V_{n}$. Since $\pi$ is continuous, $\pi^{-1}(\Lambda_{n}^{\rm Art})$ is Zariski-open in $V$, whence the assertion holds. ∎
Proposition 2.3.6.
Assume that $K$ is an infinite field. For $m\geq n$, our conjecture (Conjecture 2.3.4) is equivalent to Fröberg’s conjecture (Conjecture 2.3.3).
Proof.
Fixed $n$, $m$, and $(d_{1},\ldots,d_{m})$, we set
[…]
so that $\Lambda^{\rm csr}=\Lambda^{\rm gcsr}\cap\Lambda^{\rm Art}$.
Assume that $(f_{1},\ldots,f_{m})\in V$ is generically generalized cryptograhic semi-regular, i.e., there exists a non-empty Zariski-open set $U_{1}\subset V$ such that $U_{1}\subset\Lambda^{\rm gcsr}$. Recall from Lemma 2.3.5 that $(f_{1},\ldots,f_{m})\in V$ generically has the coordinate ring $R/\langle f_{1},\ldots,f_{m}\rangle$ that is Artinian, whence there exists a non-empty Zariski-open set $U_{2}\subset V$ such that $U_{2}\subset\Lambda^{\rm Art}$. Since both $U_{1}$ and $U_{2}$ are dense, we have $\emptyset\neq U_{1}\cap U_{2}\subset\Lambda^{\rm gcsr}\cap\Lambda^{\rm Art}=\Lambda^{\rm csr}$. Hence, $(f_{1},\ldots,f_{m})\in V$ is cryptographic semi-regular on the non-empty Zariski open set $U_{1}\cap U_{2}$ of $V$, as desired.
The converse is obvious by $\Lambda^{\rm csr}\subset\Lambda^{\rm gcsr}$. ∎
Despite of the equivalence between two conjectures, those cannot guarantee the genericness of $\Lambda^{\rm gscr}\cap(V\smallsetminus\Lambda^{\rm Art})$ with respect to the induced topology of $V\smallsetminus\Lambda^{\rm Art}$. Once the genericness holds for the case, we have the following.
Remark 2.3.7.
If the subset $\Lambda^{\rm gcsr}\cap(V\smallsetminus\Lambda^{\rm Art})$ contains a non-empty Zariski-open set in the subspace $V\smallsetminus\Lambda^{\rm Art}$ equipped with the induced topology, then ‘almost’ all sequences $(f_{1},\ldots,f_{m})$ such that $R/\langle f_{1},\ldots,f_{m}\rangle$ has Krull dimension $\geq 1$ are expected to be generalized cryptographic semi-regular.
As the first discussion about the genericness of $\Lambda^{\rm gcsr}\cap(V\smallsetminus\Lambda^{\rm Art})$, we consider the following sub case: Let $Z$ be the non-empty Zariski closed set in $V=V(n,m,d_{1},\ldots,d_{m})$ defined by the condition that the $x_{n}^{d_{i}}$-coefficient of $f_{i}$ is zero for any $i$ with $1\leq i\leq m$. Then, for any element $\bm{F}=(f_{1},\ldots,f_{m})\in Z$, the homogeneous ideal $\langle f_{1},\ldots,f_{m}\rangle_{R}$ has a projective root $(0:\cdots:0:1)$, whence $R/\langle f_{1},\ldots,f_{m}\rangle$ is not Artinian. Namely, we have $Z\subset V\smallsetminus\Lambda^{\rm Art}$.
Lemma 2.3.8.
Let $\bm{G}=(g_{1},\ldots,g_{m})\in Z$, and assume that the other coefficients of $g_{1},\ldots,g_{m}$ are algebraically independent over the prime field of $K$. Then, for any $\bm{F}=(f_{1},\ldots,f_{m})\in Z$, we have the coefficient-wise inequality
${\rm HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}(z)\geq{\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}(z).$
Therefore, if there exists a generalized cryptographic semi-regular sequence $\bm{F}=(f_{1},\ldots,f_{m})$ in $Z$ with $\dim_{K}(R/\langle f_{1},\ldots,f_{m}\rangle)_{D}=1$ for $D=\tilde{d}_{\rm reg}(\langle f_{1},\ldots,f_{m}\rangle_{R})$, then $\bm{G}=(g_{1},\ldots,g_{m})$ is generalized cryptographic semi-regular, i.e.,
${\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}(z)\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{D^{\prime}}}$
for $D^{\prime}=\widetilde{d}_{\rm reg}(\langle g_{1},\ldots,g_{m}\rangle_{R})$, and $\dim_{K}(R/\langle g_{1},\ldots,g_{m}\rangle)_{D^{\prime}}=1$.
Proof.
The first half can be proved in a way similar to the proof of [20, Lemma 1]. To prove the second one, we also use Lemma 2.2.5 (cf. [20]) for a sequence $\bm{H}=(h_{1},\ldots,h_{m})$ of generic forms with the same parameters $n,m,d_{1},\ldots,d_{m}$. Then, we have ${\rm HS}_{R/\langle\bm{G}\rangle}\geq{\rm HS}_{R/\langle\bm{H}\rangle}$ by [20, Lemma 1] and thus,
${\rm HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}\geq{\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}\geq{\rm HS}_{R/\langle h_{1},\ldots,h_{m}\rangle}\geq^{\rm lex}\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}\right],$ (2.3.1)
where $\geq^{\rm lex}$ denotes the lexicographical inequality. As $\bm{F}$ is a generalized cryptographic semi-regular sequence, we also have
${\rm HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{D}}$ (2.3.2)
for $D=\tilde{d}_{\rm reg}(\langle f_{1},\ldots,f_{m}\rangle_{R})$. Combining (2.3.1) and (2.3.2), we have
${\rm HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}\equiv{\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{D}}.$
Meanwhile, it follows from the condition $\dim_{K}(R/\langle f_{1},\ldots,f_{m}\rangle)_{D}=1$ that we obtain $\dim_{K}(R/\langle g_{1},\ldots,g_{m}\rangle)_{D}=1$ and $\tilde{d}_{\rm reg}(\langle g_{1},\ldots,g_{m}\rangle_{R})\leq D$. Hence we have
${\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{D^{\prime}}}$
for $D^{\prime}=\tilde{d}_{\rm reg}(\langle g_{1},\ldots,g_{m}\rangle_{R})$. ∎
From Lemma 2.3.8 and Remark 2.3.2, if there exists a generalized cryptographic semi-regular sequence $\bm{F}=(f_{1},\ldots,f_{m})$ with $\dim_{K}(R/\langle f_{1},\ldots,f_{m}\rangle)_{D}=1$ for $D=\tilde{d}_{\rm reg}(\langle f_{1},\ldots,f_{m}\rangle_{R})$, it follows that $Z\cap\Lambda^{\rm gcsr}$ contains a non-empty Zariski open set in $Z$ equipped with the induced topology, which implies that [29, Conjecture 4.3.4] is true for the parameter $(n,m,d_{1},\ldots,d_{m})$.
3 Proofs of main results
Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$ (in any characteristic), and let $I$ be a proper homogeneous ideal of $R$. Let $F=\{f_{1},\ldots,f_{m}\}$ be a subset of $R$ generating $I$, and put $d_{j}:=\deg(f_{j})$ (total degree) for $1\leq j\leq m$.
For a linear transformation $\sigma$ of variables $x_{1},\ldots,x_{n}$ over $K$ represented by a $n\times n$ matrix $P$ over $K$, we denote by $h^{\sigma}:=h(\bm{x}\cdot P)$ the image of $h\in R$ by $\sigma$, where $\bm{x}:=(x_{1},\ldots,x_{n})$.
3.1 Proofs of Theorem 1 (1) and (2)
Let $\prec$ be the graded reverse lexicographic ordering on $R$ with $x_{n}\prec\cdots\prec x_{1}$. We denote by $N$ the number of its projective zeros over $\overline{K}$ of the homogeneous ideal $I\subset R$. Throughout this subsection, assume that $\mathrm{Krull.dim}(R/I)\leq 1$, i.e., $N<\infty$.
Lemma 3.1.1.
If the cardinality of $K$ is greater than $N$, there exist a non-negative integer $d$ and a linear form $\ell\in R$ with $\ell\notin I$ such that the $K$-linear map $(R/I)_{d-1}\to(R/I)_{d}$ defined by the multiplication by $\ell$ is surjective.
Proof.
The assertion is proved in the proof of (i) $\Rightarrow$ (iv) in [11, Theorem 3.3.4] when $K$ is an infinite field: In fact, it suffices for the proof that one can take a linear form $\ell\in R$ that does not vanish at any projective zero over $\overline{K}$ of $I$. Even in the case where $K$ is an finite field $\mathbb{F}_{q}$, if $q$ is greater than the number $N$ of projective zeros over $\overline{K}$ of $I$, we prove that such a linear form always exists.
For each projective zero $\bm{a}=(a_{1}:\cdots:a_{n})\in\mathbb{P}^{n-1}(\overline{K})$ of $I$, we denote by $S_{\bm{a}}$ the set of all linear forms that vanish at $\bm{a}$, say
$S_{\bm{a}}=\{y_{1}x_{1}+\cdots+y_{n}x_{n}\mid y_{i}\in K\mbox{ and }y_{1}a_{1}+\cdots y_{n}a_{n}=0\},$
which is isomorphic, as $K$-linear spaces, to the following $K$-linear space of $K^{n}$:
$S_{\bm{a}}^{\prime}=\{(y_{1},\ldots,y_{n})\in K^{n}\mid a_{1}y_{1}+\cdots a_{n}y_{n}=0\}.$
The dimension of $S_{\bm{a}}^{\prime}$ as a $K$-linear space is at most $n-1$. Indeed, for an arbitrary non-zero entry $a_{i}$ of $\bm{a}$, the vector with $1$ in the $i$-th entry and 00’s elsewhere does not belong to $S_{\bm{a}}^{\prime}$, so that $S_{\bm{a}}^{\prime}$ is a proper subspace of $K^{n}$.
Here, we have
#$\#\left(\bigcup_{\bm{a}\in V_{\overline{K}}(I)}S_{\bm{a}}\right)\leq\sum_{{\bm{a}\in V_{\overline{K}}(I)}}\#S_{\bm{a}}\leq Nq^{n-1}<q^{n},$
whence a desired linear form over $K$ exists. ∎
Proof of Theorem 1 (1):
Note that, by a choice of $\ell$, the Krull dimension of $R/\langle I,\ell\rangle$ is zero. For $d^{\prime}\geq 1$, the $-\ell$ map $(R/I)_{d^{\prime}-1}\to(R/I)_{d^{\prime}}$ is surjective if and only if $(R/I)_{d^{\prime}}=0$, which follows from the following exact sequence:
${(R/I)_{d^{\prime}-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\times\ell}$${(R/I)_{d^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(R/\langle I,\ell\rangle)_{d^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
When $d^{\prime}=d$, we have $(R/\langle I,\ell\rangle)_{d}=0$, by the surjectivity of the multiplication map $(R/I)_{d^{\prime}-1}\to(R/I)_{d^{\prime}}$ by $\ell$. Thus, it follows that $(R/\langle I,\ell\rangle)_{d^{\prime}}=0$ for any $d^{\prime}\geq d$, from which the $-\ell$ map $(R/I)_{d^{\prime}-1}\to(R/I)_{d^{\prime}}$ is surjective for any $d^{\prime}$ with $d^{\prime}\geq d$. Note that the minimal $d$ for which the map is surjective is equal to $d_{\rm reg}(\langle I,\ell\rangle_{R})$.
We first note that $\widetilde{d}_{\rm reg}(I)=\widetilde{d}_{\rm reg}(I^{\tau})$ and ${d}_{\rm reg}(\langle I,\ell\rangle_{R^{\prime}})={d}_{\rm reg}(\langle I^{\tau},\ell^{\tau}\rangle_{R^{\prime}})$ for any invertible linear transformation $\tau$ of variables. There exists a $k\in\{1,\ldots,n\}$ such that $x_{k}$-coefficient of $\ell$ is not zero. We may also assume that the $x_{k}$-coefficient is $1$, without loss of generality. Let $\sigma_{1}$ be the linear transformation of variables $x_{1},\ldots,x_{n}$ defined by a permutation sending $x_{k}$ to $x_{n}$, and write $\ell^{\sigma_{1}}=a_{0}x_{0}+a_{1}x_{1}+\cdots+a_{n}x_{n}$ for $a_{i}\in K$ with $a_{n}=1$. Then, let $\sigma_{2}$ be a linear transformation of variables $x_{1},\ldots,x_{n}$ represented by the matrix
$P:=\begin{pmatrix}1&&&-a_{1}\\ &\ddots&&\vdots\\ &&1&-a_{n-1}\\ &&&1\\ \end{pmatrix},$
where blank entries are 00’s. Namely, for $h\in R$, we define
$h^{\sigma_{2}}:=h(\bm{x}\cdot P)=h(x_{1},\ldots,x_{n-1},x_{n}-a_{1}x_{1}-\cdots-a_{n-1}x_{n-1})$
with $\bm{x}:=(x_{1},\ldots,x_{n})$. Then, the invertible transformation $\sigma:=\sigma_{2}\circ\sigma_{1}$ of variables $x_{1},\ldots,x_{n}$ sends $\ell$ to $x_{n}$, and we have the equalities $D^{\prime}:=\widetilde{d}_{\rm reg}(I)=\widetilde{d}_{\rm reg}(I^{\sigma})$ and $D:=d_{\rm reg}(\langle I,\ell\rangle_{R})=d_{\rm reg}(\langle I^{\sigma},x_{n}\rangle_{R})$. The assertion follows from Theorem 2.1.9.∎
In the following, we shall prove Theorem 1 (2), for which we first prove the following easy lemma in the case where $m=n-1$:
Lemma 3.1.2.
With notation as above, if $m=n-1$ and if the sequence $(f_{1},\ldots,f_{m},\ell)$ is cryptographic semi-regular (equivalently, it is regular since $m+1=n$), then we have $\widetilde{d}_{\rm reg}(I)=d_{\rm reg}(\langle I,\ell\rangle_{R})-1$, and hence
$\mathrm{max.GB.deg}_{\prec}(I^{\sigma})\leq\mathrm{max}\{d_{\rm reg}(\langle I,\ell\rangle_{R}),\widetilde{d}_{\rm reg}(I)\}=d_{\rm reg}(\langle I,\ell\rangle_{R})=\sum_{j=1}^{n-1}(d_{j}-1)+1.$
Moreover, $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular (equivalently, it is regular since $m<n$).
Proof.
Recall that the the regularity of a sequence of homogeneous polynomials is equivalent to the semi-regularity if the number of polynomials is less than or equal to the number of variables. For the sequence $(f_{1},\ldots,f_{m},\ell)$, the number of polynomials and that of variables are both $n$, whence it is regular. Therefore, $(f_{1},\ldots,f_{m})$ is also regular, and
${\rm HS}_{R/I}(z)={\rm HS}_{R/\langle I,\ell\rangle}(z)\cdot\frac{1}{1-z}={\rm HS}_{R/\langle I,\ell\rangle}(z)\cdot(1+z+z^{2}+\cdots)$
with
${\rm HS}_{R/\langle I,\ell\rangle}(z)=\frac{\prod_{j=1}^{n-1}(1-z^{d_{j}})\cdot(1-z)}{(1-z)^{n}}=\frac{\prod_{j=1}^{n-1}(1-z^{d_{j}})}{(1-z)^{n-1}}=\prod_{j=1}^{n-1}(1+z+\cdots+z^{d_{j}-1}),$
which is a polynomial. Here, we have $\mathrm{HF}_{R/I}(d)=\sum_{i=0}^{d}\mathrm{HF}_{R/\langle I,\ell\rangle}(i)$ for any $d$. Since $\mathrm{HF}_{R/\langle I,\ell\rangle}(d)=0$ (resp. $\mathrm{HF}_{R/\langle I,\ell\rangle}(d)\neq 0$) for $d\geq d_{\rm reg}(\langle I,\ell\rangle_{R})$ (resp. $d\leq d_{\rm reg}(\langle I,\ell\rangle_{R})-1$), we have $\widetilde{d}_{\rm reg}(I)=d_{\rm reg}(\langle I,\ell\rangle_{R})-1$, as desired. ∎
Proof of Theorem 1 (2):
Suppose that $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular. Then, applying Lemma 2.2.8, we have the inequality (2.2.3), the right hand side of which is equal to $D_{\rm new}$ for $m\geq n$. It also follows from Lemma 3.1.2 that $\mathrm{max.GB.deg}_{\prec}(I)$ is upper-bounded by $\sum_{j=1}^{n-1}(d_{j}-1)+1$, which is equal to $D_{\rm new}$ for $m=n-1$, as desired.∎
3.2 Proof of Proposition 1
As in the previous subsection, let $\prec$ be the graded reverse lexicographic ordering on $R=K[x_{1},\ldots,x_{n}]$ with $x_{n}\prec\cdots\prec x_{1}$. We also denote the restriction of $\prec$ on $K[x_{1},\ldots,x_{n-1}]$ by the same notation $\prec$. In the following, we shall prove Proposition 1, namely the inequality (1.0.1) becomes an equality if $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic. Here, a weakly reverse lexicographic ideal is a monomial ideal $J$ such that if $x^{\alpha}$ is one of the minimal generators of $J$ then every monomial of the same degree which preceeds $x^{\alpha}$ must belong to $J$ as well (see also e.g., [34, Section 4] for the definition).
As in [23] and [24] (see also [28] and [29]), we consider the elimination ideal
$I^{\prime}:=\langle I,x_{n}\rangle_{R}\cap K[x_{1},\ldots,x_{n-1}]=\langle(f_{1})|_{x_{n}=0},\ldots,(f_{m})|_{x_{n}=0}\rangle_{K[x_{1},\ldots,x_{n-1}]}.$
Since $R/\langle I,x_{n}\rangle$ and $K[x_{1},\ldots,x_{n-1}]/I^{\prime}$ are isomorphic as graded rings, they have the same Hilbert series, and $d_{\rm reg}(I^{\prime})=d_{\rm reg}(\langle I,x_{n}\rangle_{R})$. Put $D:=d_{\rm reg}(I^{\prime})=d_{\rm reg}(\langle I,x_{n}\rangle_{R})$. By comparing Hilbert series, it can be also easily checked that, the sequence $(f_{1},\ldots,f_{m},x_{n})\in R^{m+1}$ is cryptographic semi-regular if and only if $(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})\in K[x_{1},\ldots,x_{n-1}]^{m}$ is cryptographic semi-regular, equivalently these sequences are both $D$-regular.
Let $G$ (resp. $G^{\prime}$) be the reduced Gröbner basis of $I$ (resp. $I^{\prime}$) with respect to $\prec$. Note that, since $\langle I,x_{n}\rangle_{R}=\langle I^{\prime},x_{n}\rangle_{R}$, the reduced Gröbner basis of $\langle I,x_{n}\rangle_{R}$ is given by $G^{\prime}\cup\{x_{n}\}$, whence $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=\mathrm{max.GB.deg}_{\prec}(\langle I,x_{n}\rangle_{R})$.
To prove Proposition 1, we start with extending some results in [28] to our case.
Lemma 3.2.1 (cf. [28, Corollary 2]).
With notation as above, we assume that $(f_{1},\ldots,f_{m},x_{n})\in R^{m+1}$ is cryptographic semi-regular (i.e., $(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})\in K[x_{1},\ldots,x_{n-1}]^{m}$ is cryptographic semi-regular), namely both the sequences are $D$-regular with $D:=d_{\rm reg}(I^{\prime})=d_{\rm reg}(\langle I,x_{n}\rangle_{R})$. Then, we have
$(\langle\mathrm{LM}_{\prec}(I)\rangle_{R})_{d}=(\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{R})_{d}$ (3.2.1)
for any $d$ with $d<D$.
Proof.
We first prove $(\langle\mathrm{LM}_{\prec}(I)\rangle_{R})_{d}\supset(\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{R})_{d}$ for any non-negative integer $d$. For this, it suffices to show that $\mathrm{LM}_{\prec}(f)\in\mathrm{LM}_{\prec}(I)$ for any $f\in I_{d}^{\prime}\smallsetminus\{0\}$. We may assume that $f$ is homogeneous, and then it is written as $f=g+x_{n}h$ for some $g\in I_{d}$ and $h\in R_{d-1}$. Here, $g$ is not divisible by $x_{n}$, since otherwise $f\notin K[x_{1},\ldots,x_{n-1}]$, a contradiction. Therefore, we have $\mathrm{LM}_{\prec}(f)=\mathrm{LM}_{\prec}(g)$ with $g\in I$, as desired.
Since $(f_{1},\ldots,f_{m},x_{n})$ is $D$-regular, it follows from [14, Proposition 2 a)] (cf. [28, Corollary 1]), its sub-sequence $(f_{1},\ldots,f_{m})$ is also $D$-regular. Hence, we have
${\rm HS}_{R/I}(z)$ $\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})\cdot(1-z)}{(1-z)^{n}}\cdot\frac{1}{1-z}$
$\equiv{\rm HS}_{R/\langle I,x_{n}\rangle}(z)\cdot\frac{1}{1-z}\equiv{\rm HS}_{R/\langle I,x_{n}\rangle}(z)\cdot(1+z+z^{2}+\cdots)$
$\equiv{\rm HS}_{K[x_{1},\ldots,x_{n-1}]/I^{\prime}}(z)\cdot(1+z+z^{2}+\cdots)\pmod{z^{D}}.$
Thus, for any $d$ with $d<D=d_{\rm reg}(I^{\prime})$, one has ${\rm HF}_{R/I}(d)=\sum_{i=0}^{d}\mathrm{HF}_{K[x_{1},\ldots,x_{n-1}]/I^{\prime}}(i)$, so that
$\dim_{K}R_{d}-\dim_{K}I_{d}=\sum_{i=0}^{d}(\dim_{K}K[x_{1},\ldots,x_{n-1}]_{i}-\dim_{K}I^{\prime}_{i}).$ (3.2.2)
Since $(JR)_{d}=x_{n}^{d}J_{0}\oplus x_{n}^{d-1}J_{1}\oplus\cdots\oplus x_{n}J_{d-1}\oplus J_{d}$ for any of $J=I^{\prime}$ and $J=K[x_{1},\ldots,x_{n-1}]$ (note that this holds also for $d\geq D$), the right hand side of (3.2.2) is equal to $\dim_{K}R_{d}-\dim_{K}(I^{\prime}R)_{d}$, so that $\dim_{K}(I^{\prime}R)_{d}=\dim_{K}I_{d}$. Here, it follows from [22, Theorem 5.2.6] that $\dim_{K}(\langle\mathrm{LM}_{\prec}(I)\rangle_{R})_{d}=\dim_{K}(\langle\mathrm{LM}_{\prec}(I^{\prime}R)\rangle_{R})_{d}$. Since $\langle\mathrm{LM}_{\prec}(I^{\prime}R)\rangle_{R}=\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{R}$, we finally obtain (3.2.1). ∎
Lemma 3.2.2 (cf. [28, Lemma 2]).
Under the same setting as in Lemma 3.2.1, for each degree $d<D$, we have
${\rm LM}_{\prec}(G)_{d}={\rm LM}_{\prec}(G^{\prime})_{d}.$ (3.2.3)
Proof.
The equality (3.2.3) can be proved by the induction on $d$, as follows. The case $d=0$ is trivial. Suppose that (3.2.3) holds for any $d$ with $10\leq d<D-$ 1.
Let $t$ be any monomial in ${\rm LM}_{\prec}(G)_{d+1}$, so that $t=\mathrm{LM}_{\prec}(g)$ for some $g\in G$. By $d+1<D$, it follows from Lemma 3.2.1 that
$(\langle\mathrm{LM}_{\prec}(I)\rangle_{R})_{d+1}=(\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{R})_{d+1},$
whence ${\rm LM}_{\prec}(g)$ is divided by ${\rm LM}_{\prec}(g^{\prime})$ for some $g^{\prime}\in G^{\prime}$. Since $G$ is reduced, the induction hypothesis implies that $\mathrm{deg}(g^{\prime})=d+1$, and thus $\mathrm{LM}_{\prec}(g)=\mathrm{LM}_{\prec}(g^{\prime})$. Therefore, we obtain ${\rm LM}_{\prec}(G)_{d+1}\subset{\rm LM}_{\prec}(G^{\prime})_{d+1}$. The same argument shows that ${\rm LM}_{\prec}(G)_{d+1}\supset{\rm LM}_{\prec}(G^{\prime})_{d+1}$, as desired. ∎
Lemma 3.2.3.
Under the same setting as in Lemma 3.2.1, we also suppose that $R/\langle I,x_{n}\rangle$ is Artinian (equivalently $K[x_{1},\ldots,x_{n-1}]/I^{\prime}$ is Artinian), namely $D<\infty$. Furthermore, we assume that $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic. Then we have $\mathrm{max.GB.deg}_{\prec}(I)\geq D$ and $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=D$.
Proof.
Put $d:=\mathrm{max.GB.deg}_{\prec}(I)$. Assume for a contradiction that $d<D$. Since $\mathrm{Krull.dim}(R/\langle I,x_{n}\rangle)=0$, it follows from Lemma 2.1.5 that there is an element $g$ in $G$ with ${\rm LM}_{\prec}(g)=x_{n-1}^{d^{\prime}}$ for some $d^{\prime}\leq d$. Then, any monomial $t\in R_{d^{\prime}}$ with $t\succeq x_{n-1}^{d^{\prime}}$ (namely $t$ is any monomial in $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}$) belongs to $\langle{\rm LM}_{\prec}(I)\rangle_{R}=\langle{\rm LM}_{\prec}(G)\rangle_{R}$ by the weakly reverse lexicographicness. Here, by $d<D$ together with (3.2.3), we have $\langle{\rm LM}_{\prec}(G)\rangle_{R}=\langle{\rm LM}_{\prec}(G^{\prime})_{\leq d}\rangle_{R}$. This implies that $t$ also belongs to $\langle{\rm LM}_{\prec}(G^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$, and hence $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}/I^{\prime}_{d^{\prime}}=0$. Thus, we have $d\geq d^{\prime}\geq d_{\rm reg}(I^{\prime})$, a contradiction.
By the same argument as above, we can show that $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=d_{\rm reg}(I^{\prime})$ as follows. Note that $\mathrm{max.GB.deg}_{\prec}(I^{\prime})\leq d_{\rm reg}(I^{\prime})$ follows from a general fact written in the second paragraph of Remark 3.2.4 below. Thus, we assume to the contrary that $d:=\mathrm{max.GB.deg}_{\prec}(I^{\prime})<D:=d_{\rm reg}(I^{\prime})$. Then, it follows from (3.2.3) that ${\rm LM}_{\prec}(G^{\prime})={\rm LM}_{\prec}(G)_{\leq d}$ and the ideal $\langle{\rm LM}_{\prec}(G^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$ has the weak reverse lexicographicness up to $d$. Since there is an element $g$ in $G^{\prime}$ with ${\rm LM}_{\prec}(g)=x_{n-1}^{d^{\prime}}$ for some $d^{\prime}\leq d$, any monomial in $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}$ belongs to $\langle{\rm LM}_{\prec}(G^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$ and so $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}/I^{\prime}_{d^{\prime}}=0$, which implies $d\geq d^{\prime}\geq D$, a contradiction. ∎
Remark 3.2.4.
In Lemma 3.2.3, when $R/\langle I,x_{n}\rangle$ is Artinian (or equivalently $K[x_{1},\ldots,x_{n-1}]/I^{\prime}$ is Artinian), we can easily prove $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=D$ if $\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$ is a weakly reverse lexicographic ideal, assuming neither that $(f_{1},\ldots,f_{m},x_{n})$ is cryptographic semi-regular nor that $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic.
In general, for any homogeneous ideal $J\subset R$ with $\mathrm{Krull.dim}(R/J)=0$, we have $\mathrm{max.GB.deg}_{\prec}(J)\leq\mathrm{reg}(J)=d_{\rm reg}(J)$ by [4, Corollary 2.5] (see also [23, Proposition 4.15]), and it is straightforward that the equality holds if $\langle\mathrm{LM}_{\prec}(J)\rangle_{R}$ is a weakly reverse lexicographic.
Proof of Proposition 1:
Put $D:=d_{\rm reg}(\langle I,x_{n}\rangle_{R})=d_{\rm reg}(I^{\prime})$ and $D^{\prime}:=\widetilde{d}_{\rm reg}(I)$. From our assumption that $R/\langle I,x_{n}\rangle$ is Artinian, we have the inequality (1.0.1), i.e., $\mathrm{max.GB.deg}(I)\leq\mathrm{max}\{D,D^{\prime}\}$. The case where $\mathrm{Krull.dim}(R/I)=0$ is clear from Remark 3.2.4, so we assume that $\mathrm{Krull.dim}(R/I)=1$. Note that, for the reduced Gröbner basis $G$ of $I$, the elements of $\mathrm{LM}_{\prec}(G)$ are the minimal generators of the monomial ideal $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}=\langle\mathrm{LM}_{\prec}(G)\rangle_{R}$.
First we consider the case where $D^{\prime}\geq D$, so that $\max\{D,D^{\prime}\}=D^{\prime}$. Put $d:=\mathrm{max.GB.deg}_{\prec}(I)$. Then, it follows from Macaulay’s basis theorem (cf. [26, Theorem 1.5.7]) that $R_{d}/I_{d}$ has a $K$-linear basis of the form
$\{t\in R_{d}:\mbox{$t$ is a monomial and }{t\notin\langle{\rm LM}_{\prec}(I)\rangle_{R}=\langle\mathrm{LM}_{\prec}(G)\rangle_{R}}\},$
which is called the standard monomial basis. We set $r:=\dim_{K}R_{d}/I_{d}$, and write this standard monomial basis as $\{t_{1},\ldots,t_{r}\}$. In the following, we prove that $\{t_{1}x_{n}^{s},\ldots,t_{r}x_{n}^{s}\}$ is a $K$-linear basis of $R_{d+s}/I_{d+s}$ for any $s$ with $s\geq 1$, from which we have $\dim_{K}R_{d}/I_{d}=\dim_{K}R_{d+s}/I_{d+s}$ for any positive integer $s$, so that $d\geq D^{\prime}$ and therefore $d=D^{\prime}$.
By Lemma 3.2.3, we have $d\geq D$, and thus it can be shown similarly to Proof of Theorem 1 (1) that the multiplication-by-$x_{n}^{s}$ map from $R_{d}/I_{d}$ to $R_{d+s}/I_{d+s}$ is surjective (since $R/\langle I,x_{n}\rangle$ is Artinian). Therefore $B_{s}:=\{t_{1}x_{n}^{s},\ldots,t_{r}x_{n}^{s}\}$ generates $R_{d+s}/I_{d+s}$. Suppose to contrary that $B_{s}$ is not a basis. In this case, there exists an $i$ such that $t_{i}x_{n}^{s}$ is divisible by $\mathrm{LM}_{\prec}(g)$ for some $g\in G$. Indeed, since the elements of $B_{s}$ are linearly dependent over $K$, we obtain $f:=a_{1}t_{1}x_{n}^{s}+\cdots+a_{r}t_{r}x_{n}^{s}\in I_{d+s}$ for some $(a_{1},\ldots,a_{r})\in K^{r}\smallsetminus\{(0,\ldots,0)\}$. Thus, for the least $i$ with $a_{i}\neq 0$, the leading monomial of $f$ is $t_{i}x_{n}^{s}$, which is divisible by $\mathrm{LM}_{\prec}(g)$ for some $g\in G$. Putting $u_{i}=\mathrm{GCD}(t_{i},\mathrm{LM}_{\prec}(g))$ and $s_{i}=t_{i}/u_{i}$, we have $u_{i}s_{i}x_{n}^{s}=t_{i}x_{n}^{s}$. Since $t_{i}=s_{i}u_{i}$ is not divisible by $\mathrm{LM}_{\prec}(g)$, we can write $u_{i}x_{n}^{s^{\prime}}=\mathrm{LM}_{\prec}(g)$ for some $s^{\prime}\leq s$. (We note that $\deg s_{i}=d-\deg u_{i}\geq\deg{\rm LM}_{\prec}(g)-\deg u_{i}=s^{\prime}$.) Note that $s^{\prime}\geq 1$ since otherwise $t_{i}$ is divisible by $\mathrm{LM}_{\prec}(g)$.
Take an arbitrary monomial $s^{\prime}_{i}$ such that $\deg s^{\prime}_{i}=s^{\prime}$ and $s_{i}^{\prime}$ divides $s_{i}$. Then, by the weakly reverse lexicographicness, as ${\rm LM}_{\prec}(g)=u_{i}x_{n}^{s^{\prime}}\preceq u_{i}s^{\prime}_{i}$, the monomial $s^{\prime}_{i}u_{i}$ should belong to $\langle{\rm LM}_{\prec}(G)\rangle_{R}$. Moreover, since $s^{\prime}_{i}u_{i}$ divides $t_{i}=s_{i}u_{i}$, the monomial $t_{i}=s_{i}u_{i}$ also belongs to $\langle{\rm LM}_{\prec}(G)\rangle_{R}$, which is a contradiction.
Finally, consider the remaining case, i.e., $D>D^{\prime}$, so that $\max\{D,D^{\prime}\}=D$. In this case, it follows from Lemma 3.2.3 that $d\geq D$, whence $d=D$.∎
3.3 Proof of Theorem 1 (3): Complexity estimation
We shall prove (3) of Theorem 1. For this, we estimate the complexity of computing a Gröbner basis of a homogeneous ideal with Macaulay matrices.
Let $F=\{f_{1},\ldots,f_{m}\}\subset R$ be a set of homogeneous polynomials in $R\smallsetminus K$, put $d_{j}:=\deg(f_{j})$ for $1\leq j\leq m$, and let $\prec$ be an arbitrary graded monomial ordering on $R$. The complexity of the Gröbner basis computation for $F$ is estimated as that of computing the reduced row echelon form (RREF) of the degree-$D$ Macaulay matrix, denoted by $M_{\leq D}(F)$, where $D$ is an arbitrary upper-bound on $\mathrm{max.GB.deg}_{\prec}(F)$. Here, the degree-$D$ Macaulay matrix $M_{\leq D}(F)$ (in homogeneous case) is a block matrix of the form
$M_{\leq D}(F)=\begin{pmatrix}M_{D}(F)&&\\ &\ddots&\\ &&M_{d_{0}}(F)\end{pmatrix}$
with $d_{0}:=\min\{\deg(f_{j}):1\leq j\leq m\}$, where each $M_{d}(F)$ is a homogeneous Macaulay matrix defined as follows: For each $d$ with $d\geq d_{0}$, the degree-$d$ homogeneous Macaulay matrix $M_{d}(F)$ of $F$ has columns indexed by the terms of $R_{d}$ sorted, from left to right, according to the chosen order $\prec$. The rows of $M_{d}(F)$ are indexed by the polynomials $m_{i,j}f_{j}$, where $m_{i,j}\in R$ is a term such that $\deg(m_{i,j}f_{j})=d$.
Theorem 3.3.1.
With notation as above, we have the following:
  1. (1)
    [2, Proposition 1] The RREF of $M_{\leq D}(F)$ is computed in
    $O\left(mD\binom{n+D-1}{D}^{\omega}\right).$ (3.3.1)
    Hence, if $\prec$ is a graded order, then the complexity of the Gröbner basis computation for $F$ is upper-bounded by (3.3.2) for $D=\mathrm{max.GB.deg}_{\prec}(F)$.
  2. (2)
    Let $D$ be a non-negative integer such that $D=O(n)$. Then, the RREF of $M_{\leq D}(F)$ is computed in
    $O\left(m\binom{n+D-1}{D}^{\omega}\right).$ (3.3.2)
    Hence, if $\prec$ is a graded order, and if $\mathrm{max.GB.deg}_{\prec}(F)=D=O(n)$, then the complexity of the Gröbner basis computation for $F$ is upper-bounded by (3.3.2).
Proof.
It suffices to estimate the complexity of computing the RREFs of the homogeneous Macaulay matrices $M_{D}(F),\ldots,M_{d_{0}}(F)$. For each $d$ with $d_{0}\leq d\leq D$, the number of rows (resp. columns) in $M_{d}(F)$ is $\sum_{1\leq i\leq m,\ d\geq d_{i}}\binom{n-1+d-d_{i}}{d-d_{i}}$ (resp. $\binom{n-1+d}{d}$). It follows from $d-d_{i}\leq d-1$ that
$\sum_{1\leq i\leq m,\ d\geq d_{i}}\binom{n-1+d-d_{i}}{d-d_{i}}\leq\sum_{i=1}^{m}\binom{n-1+d-1}{d-1}=m\binom{n-1+d-1}{d-1},$
whence the number of rows in $M_{d}(F)$ is upper-bounded by $m\binom{n-1+d-1}{d-1}$. Therefore, the complexity of the row reductions on $M_{D}(F),\ldots,M_{d_{0}}(F)$ is upper-bounded by
$\sum_{d=d_{0}}^{D}m\binom{n-1+d-1}{d-1}\binom{n-1+d}{d}^{\omega-1}$ $\leq m\sum_{d=1}^{D}\binom{n-1+d-1}{d-1}\binom{n-1+d}{d}^{\omega-1}$
$\leq m\sum_{d=1}^{D}\binom{n-1+d}{d}^{\omega},$ (3.3.3)
where we used a fact that the reduced row echelon form of a $k\times\ell$ matrix $A$ over a field $K$ can be computed in $O(k\ell^{\omega-1})$, see Remark 3.3.2 below for its proof. Clearly (3.3.3) is upper-bounded by $mD\binom{n+D-1}{D}^{\omega}$, we obtain the estimation (3.3.1) in (1).
In the following, we shall improve the estimation (3.3.1) in (1) to obtain (3.3.2) in (2), assuming that $D=O(n)$. Here, for each $d$ with $11\leq d\leq D-$ 1, we have
$\binom{n-1+d}{d}=\frac{(n-1+d)!}{(n-1)!d!}=\binom{n-1+d+1}{d+1}\times\frac{d+1}{n+d}.$
Repeating this procedure, we obtain
$\binom{n-1+d+1}{d+1}\times\frac{d+1}{n+d}=\binom{n-1+d+2}{d+2}\times\frac{d+2}{n+d+1}\times\frac{d+1}{n+d}$
$=\cdots=\binom{n-1+D}{D}\times\frac{D}{n+D-1}\times\cdots\times\frac{d+2}{n+d+1}\times\frac{d+1}{n+d}.$
It is clear that $\frac{d^{\prime}+1}{n+d^{\prime}}=1-\frac{n-1}{n+d^{\prime}}$ is monotonically increasing with respect to $d^{\prime}$. Putting $c:=(D-1)/n$, we have $D=cn+1$, so that
$\frac{d^{\prime}+1}{n+d^{\prime}}\leq\frac{D}{n+D-1}=\frac{cn+1}{(c+1)n}=\frac{c+\frac{1}{n}}{c+1}$
for any $d^{\prime}$ with $d^{\prime}\leq D-1$. As we consider $n\to\infty$, we may suppose that
$\frac{d^{\prime}+1}{n+d^{\prime}}\leq\frac{c+\frac{1}{n}}{c+1}<\frac{c+\varepsilon}{c+1}<1$
for some constant $\varepsilon$ with $10<\varepsilon<$ 1. Therefore, we obtain
$\binom{n-1+D}{D}\times\frac{D}{n+D-1}\times\cdots\times\frac{d+2}{n+d+1}\times\frac{d+1}{n+d}<\binom{n-1+D}{D}\left(\frac{c+\varepsilon}{c+1}\right)^{D-d},$
whence
$\sum_{d=1}^{D}\binom{n-1+d}{d}^{\omega}$ $<\binom{n-1+D}{D}^{\omega}\sum_{k=0}^{D-1}\left(\frac{c+\varepsilon}{c+1}\right)^{k\omega}$
$=\binom{n-1+D}{D}^{\omega}\times\frac{1-\left(\frac{c+\varepsilon}{c+1}\right)^{D\omega}}{1-\left(\frac{c+\varepsilon}{c+1}\right)^{\omega}}$
$<\binom{n-1+D}{D}^{\omega}\times\frac{1}{1-\left(\frac{c+\varepsilon}{c+1}\right)^{\omega}},$
where we put $k:=D-d$. It follows from $D=O(n)$ that $c$ is upper-bounded by a constant. By $10<\varepsilon<$ 1, it is clear that $\frac{c+\varepsilon}{c+1}$ is also upper-bounded by a positive constant smaller than $1$, whence
$\frac{1}{1-\left(\frac{c+\varepsilon}{c+1}\right)^{\omega}}=O(1)$
as $n\to\infty$. We have proved the assertion (2). ∎
Remark 3.3.2.
In the proof of Theorem 3.3.1, we used a fact that the reduced row echelon form of a $k\times\ell$ matrix $A$ over a field $K$ can be computed in $O(k\ell^{\omega-1})$, where $k$ can be greater than $\ell$. This fact is proved by considering the following procedures to compute the reduced row echelon form of $A$:
  1. (0)
    Write $k=\ell q+r$ for $q,r\in\mathbb{Z}$ with $0\leq r<\ell$.
  2. (1)
    Choose and remove $2\ell$ rows from $A$, and let $A^{\prime}$ be a $2\ell\times\ell$ matrix whose rows are the chosen $2\ell$ rows.
  3. (2)
    Compute the reduced row echelon form of $A^{\prime}$: Inserting $\ell$ zeros to the end of each row of $A^{\prime}$, construct a $2\ell\times 2\ell$ square matrix, and compute its reduced row echelon form.
  4. (3)
    As a result of (2), we obtain more than or equal to $\ell$ zero row vectors. We add the nonzero (linearly independent) row vectors obtained to $A$.
  5. (4)
    Go back to (1).
Repeating the procedures from (1) to (4) at most $q$ times, we obtain the reduced row echelon form $B$ of the initial $A$. Since the complexity of (2) is $(2\ell)^{\omega}$, the total complexity to obtain $B$ is $q\cdot(2\ell)^{\omega}\leq(k/\ell)\cdot 2^{\omega}\ell^{\omega}=O(k\ell^{\omega-1})$, as desired.
Proof of Theorem 1 (3):
For the graded reverse lexicographic ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$, it follows from Theorem 1 (2) that $D_{\rm new}$ gives an upper bound on $\mathrm{max.GB.deg}_{\prec}(F)$. Therefore, a Gröbner basis of $\langle F\rangle_{R}$ is obtained by computing the RREF of $M_{\leq D}(F)$ for $D=D^{(n,m)}$. Here, the value $D^{(n,m)}$ of (1.0.4) is maximized at $m=n$: In this case, $D^{(n,m)}$ is equal to the Lazard’s bound $\sum_{j=1}^{n}(d_{j}-1)+1$, which is $O(n)$ for fixed $d_{1},\ldots,d_{m}$. Therefore, it follows from Theorem 3.3.1 (2) that the RREF of $M_{\leq D}(F)$ for $D=D^{(n,m)}$ is computed in the complexity (1.0.5), as desired.∎
Acknowledgement
This work was supported by JSPS Grant-in-Aid for Young Scientists 23K12949, JSPS Grant-in-Aid for Scientific Research (C) 21K03377, and JST CREST Grant Number JPMJCR2113.

### v2

@@
-Generalized cryptographic semi-regular sequences: A variant of Fröberg conjecture and a simple complexity estimation for Gröbner basis computation
+Generalization of semi-regular sequences:
 Abstract
-In this paper, we study generalized cryptographic semi-regular sequences, which are expected to generic in the space of homogeneous polynomial sequences on which the coordinate rings have Krull dimension one.
-We provide an upper-bound on the complexity of the Gröbner basis computation for such a sequence.
+Nowadays, the notion of semi-regular sequences, originally proposed by Fröberg, becomes very important not only in Mathematics, but also in Information Science, in particular Cryptology.
+For example, it is highly expected that randomly generated polynomials form a semi-regular sequence, and based on this observation, secure cryptosystems based on polynomial systems can be devised.
+In this paper, we deal with a semi-regular sequence and its extension, named a generalized cryptographic semi-regular sequence, and give precise analysis on the complexity of computing a Gröbner basis of the ideal generated by such a sequence with help of several regularities of the ideal related to Lazard’s bound on maximal Gröbner basis degree and other bounds.
+We also study the genericness of the property that a sequence is semi-regular, and its variants related to Fröberg’s conjecture.
+Moreover, we discuss on the genericness of another important property that the initial ideal is weakly reverse lexicographic, related to Moreno-Socías’ conjecture, and show some criteria to examine whether both Fröberg’s conjecture and Moreno-Socías’ one hold at the same time.
 1 Introduction
@@
 Lazard’s formulation in [38] and [39] relates the problem of computing Gröbner bases to the problem of reducing so-called Macaulay matrices.
-FGLM [19] and Gröbner walk [12] were propsed as algorithms that efficiently convert a given Gröbner basis to a Gröbner basis with respect to another monomial ordering; Hilbert driven [37] can be also applied to such a conversion.
-Faugère’s proposed his innovative algorithms $F_{4}$ [17] and $F_{5}$ [18] allowing one to quite efficiently compute Gröbner bases, and these are nowadays state-of-the-art algorithms.
+FGLM [22] and Gröbner walk [13] were proposed as algorithms that efficiently convert a given Gröbner basis to a Gröbner basis with respect to another monomial ordering; Hilbert driven algorithm [46] can be also applied to such a conversion.
+Faugère proposed his innovative algorithms $F_{4}$ [20] and $F_{5}$ [21] allowing one to quite efficiently compute Gröbner bases, and these are nowadays state-of-the-art algorithms.
 We refer to [19] for a survey on signature-based algorithms including $F_{5}$.
 Thanks to Gröbner basis conversion algorithms, it is important to estimate the complexity of computing a Gröbner basis with respect to a specific monomial ordering.
-In this paper, we consider only the case of graded monomial orderings.
-In order to estimate the complexity, it suffices to provide an upper-bound on the so-called solving degree: Roughly speaking, the solving degree is the highest degree of polynomials involved in Gröbner basis computation, see [29, Section 2.2] for rigorous definitions.
-If the ideal is homogeneous, the solving degree can be equal to the maximal degree of the reduced Gröbner basis with respect to a monomial ordering that one considers, and its upper-bounds have been well-studied.
+(Throughout this paper, all the complexities are measured by the number of arithmetic operations in a field to which all the coefficients of the input polynomials belong.) In this paper, we consider only the case of graded monomial orderings; in particular, we adopt a graded reverse lexicographic ordering.
+In order to estimate the complexity, it suffices to provide an upper-bound on the so-called solving degree: Roughly speaking, the solving degree is the highest degree of polynomials involved in Gröbner basis computation, see [36, Section 2.2] for rigorous definitions.
+If the given ideal is homogeneous, the solving degree can be equal to the maximal degree of the reduced Gröbner basis with respect to a monomial ordering that one considers, and its upper-bounds have been well-studied.
 In the case of inhomogeneous ideals, the solving degree can be greater than the maximal degree of the reduced Gröbner basis, whence estimating it directly is difficult in general, see [11] and [36] for details.
 On the other hand, it is well-known (cf.
-[27, Prop.
+[34, Prop.
 4.3.18]) that a Gröbner basis of an inhomogeneous ideal with respect to a graded monomial ordering is easily obtained from a Gröbner basis of the ideal of homogenized generators.
-Therefore, the homogeneous case is very important, and we focus on that case.
-In the following, let $K$ be a field, and let $R=K[x_{1},\ldots,x_{n}]$ (resp.
-$R^{\prime}=K[x_{0},x_{1},\ldots,x_{n}]$) be the polynomial ring of $n$ (resp.
-$n+1$) variables over $K$.
-For a possibly inhomogeneous ideal $I$ of $R$ (resp.
-$R^{\prime}$), we denote by $\mathrm{max.GB.deg}_{\prec}(I)$ the maximal degree of the elements of the reduced Gröbner basis of $I$ with respect to a monomial ordering $\prec$ on the set of monomials in $R$ (resp.
-$R^{\prime}$), following [9, Definition 7] (it is also denoted by $\deg(I,\prec)$ in the literature of commutative algebra, see e.g., [24, Definition 4]).
+Therefore, the homogeneous case is essential, and we focus on that case.
+In the following, let $K$ be a field, and let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over $K$.
+The term “degree” means total degree, unless otherwise noted.
+Let $R_{d}$ be its homogeneous part of degree $d$ of $R$ (i.e., the $K$-linear space spanned by the monomials in $R$ of degree $d$) for each non-negative integer $d$.
+For an ideal (possibly inhomogeneous) $I$ of $R$, we denote by $\mathrm{max.GB.deg}_{\prec}(I)$ the maximal degree of the elements of the reduced Gröbner basis of $I$ with respect to a monomial ordering $\prec$ on the set $\mathcal{T}$ of monomials in $R$, following [10, Definition 7] (it is also denoted by $\deg(I,\prec)$ in the literature of commutative algebra, see e.g., [31, Definition 4]).
 If $I$ is generated by $F\subset R$, we may write $\mathrm{max.GB.deg}_{\prec}(F)$ in place of it.
-For a homogeneous ideal $I\subset R=K[x_{1},\ldots,x_{n}]$ with a generating set $F=\{f_{1},\ldots,f_{m}\}\subset R$, it is well-known (cf.
-[2, Proposition 1]) that a Gröbner basis of $I$ with respect to a graded monomial ordering up to degree $d$ can be computed in $O(md\binom{n+d-1}{d}^{\omega})$, where $32\leq\omega<$ 3 is the exponent of matrix multiplication.
-(We can also obtain another upper-bound $O(m\binom{n+d}{d}^{\omega})$ by the proof of [35, Theorem 1.72].) With this fact, we have in the homogeneous case that the complexity of computing the reduced Gröbner basis of $I$ is upper-bounded by $O(mD\binom{n+D-1}{D}^{\omega})$ (or $O(m\binom{n+D}{D}^{\omega})$) for any integer $D$ with $\mathrm{max.GB.deg}_{\prec}(I)\leq D$.
-Hence, finding a tighter bound $D$ of $\mathrm{max.GB.deg}_{\prec}(I)$ is important for estimating the complexity of Gröbner basis computation in the homogeneous case.
-The case on which we focus on is the case where the Krull dimension $\mathrm{Krull.dim}(R/I)$ is equal to or less than one; see e.g., [24, Section 1] for a summary of existing results when $\mathrm{Krull.dim}R/I\geq 2$.
-In the case where $\mathrm{Krull.dim}(R/I)\leq 1$, Lazard’s bound (called also Macaulay’s bound) is well-known, and it is given by $\sum_{j=1}^{\ell}(d_{j}-1)+1$ with $d_{1}\geq d_{2}\geq\cdots\geq d_{m}$ and $\ell:=\max\{m,n\}$, where we set $d_{j}=\deg(f_{j})$ for $1\leq j\leq m$.
-When $I$ is in quasi stable position (equivalently $R/\langle I,x_{n}\rangle$ is Artinian, see Lemma 2.1.8 below), we can apply Hashemi-Seiler’s inequality
-$\mathrm{max.GB.deg}_{\prec}(I)\leq\mathrm{max}\{\mathrm{hilb}(\langle I,x_{n}\rangle),\mathrm{hilb}(I)\}$ (1.0.1)
-for the graded reverse lexicographic ordering $\prec$ with $x_{n}\prec\cdots\prec x_{1}$, where $\mathrm{hilb}(J)$ denotes the Hilbert regularity of a homogeneous ideal $J$ of $R$.
-Note that $\mathrm{hilb}(\langle I,x_{n}\rangle)$ is equal to the degree of regularity ${d}_{\rm reg}(\langle I,x_{n}\rangle)$ since $R/\langle I,x_{n}\rangle$ is Artinian, and $\mathrm{hilb}(I)$ is equal to the generalized degree of regularity $\widetilde{d}_{\rm reg}(I)$.
-For the definition of $d_{\rm reg}$ and $d_{\rm reg}$, see Remark 2.1.2 below.
-On the other hand, in [29], when $F=\{f_{1},\ldots,f_{m}\}\subset R$ with $m\geq n$ is obtained by homogenizing an affine semi-regular sequence in $K[x_{1},\ldots,x_{n-1}]^{m}$, we improved Lazard’s bound on $\mathrm{max.GB.deg}_{\prec}(\langle F\rangle_{R})$ to $\sum_{j=1}^{n}(d_{j}-1)+1$ with $d_{1}\leq\cdots\leq d_{m}$.
-Under an additional assumption that $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular (see Definition 2.2.7), we also proved in [29] that the right hand side of (1.0.1) is upper-bounded by the degree of regularity for a cryptographic semi-regular sequence of $m$ polynomials of degrees $d_{1},\ldots,d_{m}$ in $n+1$ variables.
-In this paper, we shall prove that, if $K$ is large enough, the inequality (1.0.1) always holds after a linear coordinate change.
-Then, we also estimate the complexity of computing a Gröbner basis for $I$, assuming that $I$ is generated by a generalized cryptographic semi-regular sequence.
-Such a sequence is expected to be generic (see Definition 2.3.1 for definition), analogously to Fröberg’s conjecture [20].
-Our results are summarized into the following theorem and proposition:
-Theorem 1.
-Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$ (in any characteristic), and let $I$ be a proper homogeneous ideal of $R$.
-Let $F=\{f_{1},\ldots,f_{m}\}$ be a subset of $R$ generating $I$, and put $d_{j}:=\deg(f_{j})$ (total degree) for $1\leq j\leq m$.
-Assume that $I$ has finitely many projective zeros over $\overline{K}$, and suppose that the cardinality of $K$ is greater than the number of projective zeros over $\overline{K}$ of $I$.
-Choose an arbitrary linear form $\ell$ in $R$ such that it does not vanish at any projective zero over $\overline{K}$ of $I$ (such an $\ell$ always exists by our assumption $\mathrm{Krull.dim}(R/I)\leq 1$).
-Then, the following (1) holds:
-  1. (1)
-    There exists a linear transformation $\sigma$ of variables $x_{1},\ldots,x_{n}$ over $K$ sending $\ell$ to $x_{n}$ such that
-    ${\rm max.GB.deg}_{\prec}(I^{\sigma})\leq\mathrm{max}\{d_{\rm reg}(\langle I,\ell\rangle_{R}),\widetilde{d}_{\rm reg}(I)\}$ (1.0.2)
-    for the graded reverse lexicographic ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$, where $I^{\sigma}$ is the image of $I$ by $\sigma$.
-Moreover, if $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular, then following (2) and (3) also hold:
+Lazard’s bound and Hashemi-Seiler’s inequality
+For a homogeneous ideal $I\subset R=K[x_{1},\ldots,x_{n}]$ with a generating set $F=\{f_{1},\ldots,f_{m}\}\subset R$, it is well-known (see e.g., [3, Proposition 1], [45, Equation (24)]) that a Gröbner basis of $I$ with respect to a graded monomial ordering up to degree $d$ can be computed in $O(md\binom{n+d-1}{d}^{\omega})$, where $32\leq\omega<$ 3 is the exponent of matrix multiplication.
+(We can also obtain another upper-bound $O(m\binom{n+d}{d}^{\omega})$ by the proof of [44, Theorem 1.72].) From this fact, it follows in the homogeneous case that the complexity of computing the reduced Gröbner basis of $I$ is upper-bounded by $O(mD\binom{n+D-1}{D}^{\omega})$ (or $O(m\binom{n+D}{D}^{\omega})$) for any integer $D$ with $\mathrm{max.GB.deg}_{\prec}(I)\leq D$.
+Hence, finding a tighter bound $D$ of $\mathrm{max.GB.deg}_{\prec}(I)$ is important for estimating shapely the complexity of Gröbner basis computation in the homogeneous case.
+The case on which we focus is the case where the Krull dimension of $R/I$, denoted by $\mathrm{Krull.dim}(R/I)$, is equal to or less than one; see e.g., [31, Section 1] for a summary of existing results when $\mathrm{Krull.dim}(R/I)\geq 2$.
+In the case where $\mathrm{Krull.dim}(R/I)\leq 1$, Lazard’s bound (called also Macaulay’s bound) is well-known.
+Specifically, if $K$ is large enough, then there exists a (suitable) linear coordinate change $\sigma$ such that $R/(I^{\sigma}+\langle x_{n}\rangle)$ is Artinian (equivalently $I$ is in quasi stable position, see Lemma 2.1.8 below).
+From this, we assume without loss of generality that $R/(I+\langle x_{n}\rangle)$ is Artinian.
+In this case, putting $d_{j}=\deg(f_{j})$ for $1\leq j\leq m$ and supposing that $d_{1}\geq d_{2}\geq\cdots\geq d_{m}$ (descending order), it follows that
+$\mathrm{max.GB.deg}_{\prec}(I)\leq\sum_{j=1}^{\min\{m,n\}}(d_{j}-1)+1$ (1.1)
+for the graded reverse lexicographical ordering $\prec$ with $x_{n}\prec\cdots\prec x_{1}$, see [39, Théorème 3.3] and [38, Theorem 2].
+Recently, Hashemi-Seiler [31] re-proved Lazard’s bound by using the inequality
+$\mathrm{max.GB.deg}_{\prec}(I)\leq\mathrm{max}\{\mathrm{hilb}(I+\langle x_{n}\rangle),\mathrm{hilb}(I)\},$ (1.2)
+where $\mathrm{hilb}(J)$ denotes the Hilbert regularity of a homogeneous ideal $J$ of $R$.
+Note that $\mathrm{hilb}(I+\langle x_{n}\rangle)$ is equal to the degree of regularity ${d}_{\rm reg}(I+\langle x_{n}\rangle)$ since $R/(I+\langle x_{n}\rangle)$ is Artinian, and $\mathrm{hilb}(I)$ is equal to the generalized degree of regularity $\widetilde{d}_{\rm reg}(I)$.
+For the definition of $d_{\rm reg}$ and $\widetilde{d}_{\rm reg}$, see Remark 2.1.2 below.
+Remark 1.
+Note that the condition $\mathrm{Krull.dim}(R/I)\leq 1$ does not necessarily imply $d_{\rm reg}(I+\langle x_{n}\rangle)=\mathrm{hilb}(I+\langle x_{n}\rangle)<\infty$, without considering any linear coordinate change $\sigma$.
+For example, putting $F:=\{f_{1}:=x_{1}x_{2}-1,f_{2}:=x_{2}-1\}\subset K[x_{1},x_{2}]$ (inhomogeneous), it is obvious that $\mathrm{Krull.dim}(R[y]/I)=1$ for $I=\langle F^{h}\rangle$ but $d_{\rm reg}(I+\langle y\rangle)=d_{\rm reg}(\langle F^{\rm top}\rangle)=\infty$, where $F^{h}$ is the homogenization of $F$ by an extra variable $y$, and where $F^{\rm top}$ is the homogeneous part of highest degree, see Remark 3 below.
+In this case, $F^{h}=\{x_{1}x_{2}-y^{2},x_{2}-y\}$ and $F^{\rm top}=\{x_{1}x_{2},x_{2}\}$.
+On the other hand, if $K$ is a finite field of order $q$ and if we would compute a zero of $\langle F\rangle$ over $K$, we often add the set of so-called field equations $x_{i}^{q}-x_{i}$ to an inhomogeneous $F$, say $F_{1}:=F\cup\{x_{i}^{q}-x_{i}:1\leq i\leq n\}$, so that we always have $d_{\rm reg}(\langle F_{1}^{h}\rangle+\langle y\rangle)=d_{\rm reg}(\langle F_{1}^{\rm top}\rangle)<\infty$.
+Our results and conjectures
+In this paper, we focus on several cases that are expected to be generic in some sense, and obtain upper-bounds on $\mathrm{max.GB.deg}_{\prec}(I)$ and some results related to Fröberg’s conjecture [23] and Moreno-Socías’ one [41], in those cases.
+For example, we will assume that $R/(I+\langle x_{n}\rangle)$ has the Hilbert series associated to generic forms, as conjectured by Fröberg [23] (see Conjecture 2.3.7 below).
+Note that a homogeneous polynomial sequence is called cryptographic semi-regular (cf.
+[8]) if its associated coordinate ring has the Hilbert series conjectured by Fröberg, see Definition 2.2.3 and Proposition 2.2.4 below for details.
+In the literature, a cryptographic semi-regular sequence plays a very important role in analyzing Gröbner basis computation, both in theory and in practical applications to cryptography, see e.g., Bardet et al’s works [1], [2], [3], [4], Gorla et al.’s works [8], [10], [11], and our ones [35], [36].
+In our earlier work [36], when $m\geq n$, we improved Lazard’s bound on $\mathrm{max.GB.deg}_{\prec}(F)$ to $\sum_{j=1}^{n}(d_{j}-1)+1$ with $d_{1}\leq\cdots\leq d_{m}$ (ascending order), supposing that $(f_{1}|_{x_{n}=0},\ldots,f_{i}|_{x_{n}=0})$ is cryptographic semi-regular on $K[x_{1},\ldots,x_{n-1}]$ for each $1\leq i\leq m$.
+In [36], we also defined a generalized cryptographic semi-regular sequence (see Definition 2.2.8 below for the definition and Remark 2.2.9 for a construction), as an extension of cryptographic semi-regular sequence to the case of Krull dimension one.
+As the first result of this paper, we prove Theorem 1.1, which provides an upper-bound on $\mathrm{max.GB.deg}_{\prec}(I)$ for a generalized cryptographic semi-regular sequence and a complexity bound on computing a Gröbner basis of $I$.
+Theorem 1.1.
+Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$ (in any characteristic), and $\prec$ the graded reverse lexicographical ordering with $x_{n}\prec\cdots\prec x_{1}$.
+Let $I$ be a homogeneous ideal of $R$ generated by homogeneous polynomials $f_{1},\ldots,f_{m}$ in $R\smallsetminus K$ such that $\mathrm{Krull.dim}(R/I)\leq 1$.
+Assume that $(f_{1}|_{x_{n}=0},\ldots,f_{m}|_{x_{n}=0})$ is cryptographic semi-regular on $K[x_{1},\ldots,x_{n-1}]$ (equivalently $(f_{1},\ldots,f_{m},x_{n})$ is cryptographic semi-regular), and that $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular, Then the following (1) and (2) hold:
   1. (1)
     We have
-    $\mathrm{max}\{d_{\rm reg}(\langle I,\ell\rangle_{R}),\widetilde{d}_{\rm reg}(I)\}\leq D^{(n,m)},$ (1.0.3)
+    $\mathrm{max.GB.deg}_{\prec}(I)\leq\mathrm{max}\{d_{\rm reg}(I+\langle x_{n}\rangle),\widetilde{d}_{\rm reg}(I)\}\leq D^{(n,m)},$ (1.3)
     where
-    $D^{(n,m)}:=\left\{\!\begin{array}[]{cc}\displaystyle{\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}\right]\right)}+1&(m\geq n),\\ \displaystyle\sum_{i=1}^{n-1}(d_{i}-1)+1&(m=n-1).\end{array}\right.$ (1.0.4)
+    $D^{(n,m)}:=\left\{\!\begin{array}[]{cc}\displaystyle{\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}\right]\right)}+1&(m\geq n),\\ \displaystyle\sum_{i=1}^{n-1}(d_{i}-1)+1&(m=n-1).\end{array}\right.$ (1.4)
   2. (2)
-    The complexity of computing a Gröbner basis of $I^{\sigma}$ with respect to $\prec$ is upper-bounded by
-    $O\left(m\binom{n+D^{(n,m)}-1}{D^{(n,m)}}^{\omega}\right),$ (1.0.5)
-    where $32\leq\omega<$ 3 is the exponent of matrix multiplication.
-Proposition 1.
-Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$ (in any characteristic), and let $\prec$ be the graded reverse lexicographic ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$.
-Let $I$ be a proper homogeneous ideal of $R$.
-Assume that the Krull dimension of $R/I$ is $\leq 1$ and $R/\langle I,x_{n}\rangle$ is Artinian.
-If $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic, then the inequality (1.0.1) becomes an equality.
-Hence, the inequality (1.0.2) in Theorem 1 (2) also becomes an equality.
-Theorem 1 can be viewed as a generalization of [29, Proposition 1 and Corollary 1].
-The bound $D_{\rm new}$ given in (1.0.4) coincides with Lazard’s bound when $m\leq n$, while it can be sharper than Lazard’s one when $m>n$.
-We also note that, the complexity bound (1.0.5) is better than the usual bound $O(mD\binom{n+D-1}{D}^{\omega})$ (see e.g., [2, Proposition 1], [36, Equation (24)]) and than $O(m\binom{n+D}{D}^{\omega})$ (which comes from the proof of [35, Theorem 1.72]) for an arbitrary $D$ with $D\geq\mathrm{max.GB.deg}_{\prec}(F)$.
-Our assumption that a homogeneous polynomial sequence is generalized cryptographic semi-regular could be useful to analyze the security of multivariate cryptosystems and so on (cf. [29, Section 4.3]).
+    The complexity of computing a Gröbner basis of $I$ with respect to $\prec$ is upper-bounded by
+    $O\left(m\binom{n+D^{(n,m)}-1}{D^{(n,m)}}^{\omega}\right)$ (1.5)
+    arithmetic operations in $K$, where $32\leq\omega<$ 3 is the exponent of matrix multiplication.
+The bound $D^{(n,m)}$ in (1.3) is optimal for a homogeneous ideal $I$ satisfying the assumptions of Theorem 1.1, see Example 3.1.4 below.
+The proof of (2) in Theorem 1.1 is based on a well-known fact that the a Gröbner basis of $I$ is obtained by computing the reduced row echelon form of the degree-$D$ Macaulay matrix for any $D$ with $D\geq\mathrm{max.GB.deg}_{\prec}(I)$, see Section 3.1 for details.
+Our assumption that a homogeneous polynomial sequence is generalized cryptographic semi-regular could be useful to analyze the security of multivariate cryptosystems and so on (cf.
+[36, Section 4.3]).
+The bound $D^{(n,m)}$ given in (1.4) coincides with Lazard’s bound when $m\in\{n-1,n\}$, while it can be sharper than Lazard’s one when $m>n$.
+We also note that, the complexity bound (1.5) is slightly better than the usual bound $O(mD\binom{n+D-1}{D}^{\omega})$ (see e.g., [3, Proposition 1], [45, Equation (24)]) and than $O(m\binom{n+D}{D}^{\omega})$ (which comes from the proof of [44, Theorem 1.72]) for an arbitrary $D$ with $D\geq\mathrm{max.GB.deg}_{\prec}(F)$.
+As the second result, we shall prove that Hashemi-Seiler’s inequality (1.2) is as possible as tight under some assumptions, one of which is that an initial ideal is weakly reverse lexicographic.
+Here, a weakly reverse lexicographic ideal is a monomial ideal $J$ such that if $x^{\alpha}$ is one of the minimal generators of $J$ then every monomial of the same degree which preceeds $x^{\alpha}$ must belong to $J$ as well.
+Moreno-Socías [41] conjectured that, for the graded reverse lexicographical ordering, the initial ideal of a generic ideal is a weakly reverse lexicographic.
+Theorem 1.2.
+Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$ (in any characteristic), and let $\prec$ be the graded reverse lexicographical ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$.
+Let $I$ be a proper homogeneous ideal of $R$ with $\mathrm{Krull.dim}(R/I)\leq 1$ and $\mathrm{Krull.dim}(R/(I+\langle x_{n}\rangle))=0$.
+If $(f_{1}|_{x_{n}=0},\ldots,f_{m}|_{x_{n}=0})$ is cryptographic semi-regular on $K[x_{1},\ldots,x_{n-1}]$ (equivalently $(f_{1},\ldots,f_{m},x_{n})$ is cryptographic semi-regular) and if the initial ideal ${\rm in}_{\prec}(I)=\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic, then the inequality (1.2) becomes an equality.
+After proving Theorems 1.1 and 1.2, in Section 4, we also study the genericness of our assumptions used in the theorems, based on the theory of parametric Gröbner bases.
+For this, in Section 2.3, we formulate the genericness of cryptographic semi-regular sequences in terms of parametric Gröbner bases.
+Then, in Section 4, some analogous results for generalized cryptographic semi-regular sequences will be proved, and the following conjectures will be raised as variants of Fröberg’s conjecture:
+Conjecture 1.3 (Conjecture 4.1.10; see also [36, Conjecture 4.3.4]).
+Assume that $K$ is an infinite field, and we endow $V_{n,m,(d_{1},\ldots,d_{m})}:=R_{d_{1}}\times\cdots\times R_{d_{m}}$ naturally with the Zariski topology.
+For any $n$, $m$, and $(d_{1},\ldots,d_{m})$ and for any $\bm{p}\in\mathbb{P}^{n-1}(K)$, the property that $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ is generalized cryptographic semi-regular is generic on $Z_{\bm{p}}:=\{\bm{H}\in V_{n,m,(d_{1},\ldots,d_{m})}:\bm{p}\in V_{\overline{K}}(\bm{H})\}$, namely there exists a non-empty Zariski-open set of $Z_{\bm{p}}$ with respect to the induced topology on which the property holds.
+Conjecture 1.4 (Conjecture 4.1.11).
+Assume that $K$ is an infinite field.
+For any $n$, $m$, and $(d_{1},\ldots,d_{m})$, the property that $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ is generalized cryptographic semi-regular is generic on $V_{n,m,(d_{1},\ldots,d_{m})}^{\dim\geq 1}:=\{\bm{H}\in V_{n,m,(d_{1},\ldots,d_{m})}:\mathrm{Krull.dim}(R/\langle\bm{H}\rangle)\geq 1\}$, namely there exists a non-empty Zariski-open set of $V_{n,m,(d_{1},\ldots,d_{m})}^{\dim\geq 1}$ with respect to the induced topology on which the property holds.
+We also show a new criterion to examine whether Fröberg’s conjecture and Moreno-Socías’ one hold at the same time, for a fixed parameter set $(n,m,d_{1},\ldots,d_{m})$ (although these conjectures are set for any parameter set, see Remark 2 for details):
+Theorem 1.5 (Theorem 4.2.1).
+Assume that $K$ is an infinite field, and let $n,m,d_{1},\ldots,d_{m}$ be positive integers.
+If there exists a cryptographic semi-regular sequence $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ such that ${\rm in}_{\prec}(\langle\bm{F}\rangle)$ is weakly reverse lexicographic, then there also exists a non-empty Zariski-open set $U$ of $V_{n,m,(d_{1},\ldots,d_{m})}$ such that $\bm{H}$ is cryptographic semi-regular and ${\rm in}_{\prec}(\langle\bm{H}\rangle)$ is weakly reverse lexicographic for any $\bm{H}\in U$.
+Namely, Fröberg’s conjecture and Moreno-Socías’ one both hold for the parameter set $(n,m,d_{1},\ldots,d_{m})$.
+Remark 2.
+Pardue proved in his seminal paper [43] that Fröberg’s conjecture is equivalent to his several conjectures, and that Moreno-Socías’ conjecture implies all of them.
+Note that these equivalence and implication are intended to mean that each conjecture holds for any choice of parameters $(n,m,d_{1},\ldots,d_{m})$.
+More precisely, we see from the proof of his theorem [43, Theorem 2] that (1) and (2) are equivalent to each other and that (3) implies (1) and (2), over an infinite field:
+  1. (1)
+    (Fröberg’s conjecture) The property that $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ is cryptographic semi-regular is generic (in terms of Definition 2.3.1) for any $(n,m,d_{1},\ldots,d_{m})$.
+  2. (2)
+    (Pardue’s conjecture [43, Conjecture B]) The property that $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ is semi-regular is generic for any $(n,m,d_{1},\ldots,d_{m})$.
+  3. (3)
+    (Moreno-Socías’ conjecture [41]) The property that the initial ideal of the homogeneous ideal generated by $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ with respect to a graded reverse lexicographical ordering is weakly reverse lexicographic is generic for any $(n,m,d_{1},\ldots,d_{m})$.
+Namely, even if the property in (3) is shown to be generic for a fixed parameter set $(n,m,d_{1},\ldots,d_{m})$, it is unclear that the property in (1) (or (2)) is generic for the same parameter set (this also applies to the case of (1) $\Rightarrow$ (2)).
+Analogously to Theorem 1.5, we finally prove the following theorem, which can also examine Conjecture 1.3:
+Theorem 1.6 (Proposition 4.2.3 and Proposition 4.2.4).
+Under the same setting as in Theorem 1.5, assume that there exists a generalized cryptographic semi-regular sequence $\bm{F}$ in $V_{n,m,(d_{1},\ldots,d_{m})}$ with $\mathrm{Krull.dim}(R/\langle\bm{F}\rangle)=1$ such that $\dim_{K}(R/\langle\bm{F}\rangle)_{D}=1$ for $D=\widetilde{d}_{\rm reg}(\langle\bm{F}\rangle)$ and ${\rm in}(\langle\bm{F}\rangle)$ is weakly reverse lexicographic.
+Then, for any $\bm{p}\in\mathbb{P}^{n-1}(K)$, there exists a non-empty Zariski-open set $U_{\bm{p}}$ of $Z_{\bm{p}}:=\{\bm{H}\in V_{n,m,(d_{1},\ldots,d_{m})}:\bm{p}\in V_{\overline{K}}(\bm{H})\}$ with respect to the induced topology such that $\bm{H}$ is generalized cryptographic semi-regular and ${\rm in}(\langle\bm{H}\rangle)$ is weakly reverse lexicographic for any $\bm{H}$ in $U_{\bm{p}}$.
+To the end of Introduction, we remark an application of Theorem 1.1 to the inhomogeneous case.
 Remark 3 (Inhomogeneous case).
-By Theorem 1, we can also estimate the complexity of computing a Gröbner basis for an inhomogeneous polynomial sequence defining a zero-dimensional ideal, as in [29, Section 4.3].
-Indeed, for a subset $F=\{f_{1},\ldots,f_{m}\}$ of $R=K[x_{1},\ldots,x_{n}]$ containing at least one inhomogeneous polynomial, a Gröbner basis of $\langle F\rangle_{R}$ is obtained by reducing the associated degree-$D$ Macaulay matrix at $D=\mathrm{max.GB.deg}_{\prec^{h}}(I)$ with $I=\langle f_{1}^{h},\ldots,f_{m}^{h}\rangle\subset R^{\prime}:=R[y]$, see [9, Theorem 7].
-Here, each $f_{j}^{h}$ denotes the homogenization of $f_{j}$ by an extra variable $y$, and $\prec^{h}$ is the homogenization of $\prec$, see e.g., [27, Section 3.1.1] for definition.
+By Theorem 1.1, we can also estimate the complexity of computing a Gröbner basis for an inhomogeneous polynomial sequence defining a zero-dimensional ideal, as in [36, Section 4.3].
+Indeed, for a subset $F=\{f_{1},\ldots,f_{m}\}$ of $R=K[x_{1},\ldots,x_{n}]$ containing at least one inhomogeneous polynomial, a Gröbner basis of $\langle F\rangle_{R}$ is obtained by reducing the associated degree-$D$ Macaulay matrix at $D=\mathrm{max.GB.deg}_{\prec^{h}}(I)$ with $I=\langle f_{1}^{h},\ldots,f_{m}^{h}\rangle\subset R^{\prime}:=R[y]$, see [10, Theorem 7].
+Here, each $f_{j}^{h}$ denotes the homogenization of $f_{j}$ by an extra variable $y$, and $\prec^{h}$ is the homogenization of $\prec$, see e.g., [34, Section 3.1.1] for definition.
 The complexity of reducing the Macaulay matrix can be easily estimated as $O(m\binom{n+D}{D}^{\omega})$, without assuming $D=O(n)$.
@@
   • The sequence $(f_{1}^{h},\ldots,f_{m}^{h},y)\in(R^{\prime})^{m+1}$ is cryptographic semi-regular, equivalently $(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})\in R^{m}$ is cryptographic semi-regular.
-  • The sequence $(f_{1}^{h},\ldots,f_{m}^{h})\in(R^{\prime})^{m+1}$ is generalized cryptographic semi-regular.
-Then, by the first assumption, it follows from Theorem 1 (1) that the inequality (1.0.2) holds for $I=\langle F^{h}\rangle_{R^{\prime}}$ and $\ell=y$, where we replace $R$ by $R^{\prime}$ and set $\sigma$ to be the identity.
-Moreover, under the second and third assumptions, we also have the complexity bound
-$O\left(m\binom{n+D^{(n+1,m)}}{D^{(n+1,m)}}^{\omega}\right)$ (1.0.6)
-by Theorem 1 (2). Note that this upper-bound is heuristically found in [21].
+  • The sequence $(f_{1}^{h},\ldots,f_{m}^{h})\in(R^{\prime})^{m}$ is generalized cryptographic semi-regular.
+Then, it follows from Theorem 1.1 that $\max\{d_{\rm reg}(\langle F^{\rm top}\rangle),\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\}\leq D^{(n+1,m)}$ with $F^{\rm top}:=\{f_{1}^{\rm top},\ldots,f_{m}^{\rm top}\}$ and we have the complexity bound
+$O\left(m\binom{n+D^{(n+1,m)}}{D^{(n+1,m)}}^{\omega}\right).$ (1.6)
+Note that this upper-bound is heuristically found in [28].
 2 Preliminaries
@@
 Let $n$ be a natural number, and let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over $K$.
-For $d\geq 0$, we denote by $R_{d}$ the homogeneous component of $R$ of degree $d$.
+We denote by $\mathcal{T}$ the set of monomials in $\{x_{1},\ldots,x_{n}\}$.
+As our convention, as graded structure objects, for a subset $S$ of $R$ and a non-negative integer $d$, we also denote by $S_{d}$ the subset consisting of all elements of degree $d$ in $S$.
+As an exception, if $S$ is an ideal of $R$, it is allowed that $S_{d}$ contains 00.
 The Krull dimension of a commutative ring $A$ is denoted by $\mathrm{Krull.dim}(A)$.
@@
 2.1 Hilbert regularity
-Let $I$ be a homogeneous ideal of $R$.
+Let $I$ be a non-zero homogeneous ideal of $R$.
 Let ${\rm HP}_{R/I}\in\mathbb{Q}[z]$ denote the Hilbert polynomial of $R/I$, namely there exists a non-negative integer $d_{0}$ such that ${\rm HF}_{R/I}(d)={\rm HP}_{R/I}(d)$ for any $d$ with $d\geq d_{0}$, where $\mathrm{HF}$ is the Hilbert function.
-Definition 2.1.1 (Hilbert regularity (or index of regularity [13, Chapter IX, Section 3])).
+Definition 2.1.1 (Hilbert regularity (or index of regularity [14, Chapter IX, Section 3])).
 The smallest $d_{0}$ such that ${\rm HF}_{R/I}(d)={\rm HP}_{R/I}(d)$ for any $d$ with $d\geq d_{0}$ is called the Hilbert regularity of $I$, denoted by ${\rm hilb}(I)$, or is called the index of regularity of $I$, denoted by $i_{\rm reg}(I)$.
@@
   4. (4)
-    There exists a non-negative integer $d_{0}$ such that ${\rm HF}_{R/I}(d)={\rm HF}_{R/I}(d_{0})$ (resp. ${\rm HF}_{R/I}(d)=0$) for any $d$ with $d\geq d_{0}$.
+    There exists a non-negative integer $d_{0}$ such that ${\rm HF}_{R/I}(d)={\rm HF}_{R/I}(d_{0})>0$ (resp. ${\rm HF}_{R/I}(d)=0$) for any $d$ with $d\geq d_{0}$.
 In this case, it is straightforward that
@@
 The middle and the right hand sides are called the generalized degree of regularity of $I$ for $r\leq 1$, denoted by $\widetilde{d}_{\rm reg}(I)$ (cf.
-[29, Sect.
+[36, Sect.
 2.3]): When $r=0$, these are equal to the degree of regularity $d_{\rm reg}(I):=\min\{d:R_{d}=I_{d}\}$.
@@
 Remark 2.1.3.
-We also note that ${\rm HS}_{R/I}$ is written as
-${\rm HS}_{R/I}(z)=\dfrac{h_{R/I}(z)}{(1-z)^{r}}$
-for some polynomial $h_{R/I}(z)\in\mathbb{Z}[z]$, which is called the $h$-polynomial of $R/I$. Then, we have ${\rm hilb}(I)=\deg h_{R/I}-r+1$, see [6, Prop. 4.1.12].
+Note also that the Hilbert series ${\rm HS}_{R/I}$ is written as ${\rm HS}_{R/I}(z)=h_{R/I}(z)/{(1-z)^{r}}$ for some polynomial $h_{R/I}(z)\in\mathbb{Z}[z]$, which is called the $h$-polynomial of $R/I$.
+Then, we have ${\rm hilb}(I)=\deg h_{R/I}-r+1$, see [7, Prop.
+4.1.12].
 Moreover, it follows from Grothendieck-Serre’s formula that
 ${\rm hilb}(I)\leq\mathrm{reg}(I)=\mathrm{reg}(R/I)+1,$
-where ‘${\rm reg}$’ means Castelnuovo-Mumford regularity, see [15, §20.5] for the definition. The equality holds if $R/I$ is Artinian, i.e., the Krull dimension of $R/I$ is zero.
-Definition 2.1.4 (cf. [24, Section 3]).
+where ‘${\rm reg}$’ means Castelnuovo-Mumford regularity, see [17, §20.5] for the definition.
+The equality holds if $R/I$ is Artinian, i.e., the Krull dimension of $R/I$ is zero, see e.g., [18, Corollary 4.4].
+Definition 2.1.4 (cf. [31, Section 3]).
 We say that $I$ is in Noether position with respect to the variables $x_{1},\ldots,x_{n-r}$, where $r=\mathrm{Krull.dim}(R/I)$, if the injective ring extension $K[x_{n-r+1},\ldots,x_{n}]\to R/I$ is integral.
-Lemma 2.1.5 ([5, Lemma 4.1]; see also [32, Proposition 5.4]).
-Let $I$ be a homogeneous ideal of $R$, and let $\prec$ be the reverse lexicographic order on $R$ with $x_{n}\prec\cdots\prec x_{1}$. Then, the following are equivalent:
+Lemma 2.1.5 ([6, Lemma 4.1]; see also [40, Proposition 5.4]).
+Let $I$ be a homogeneous ideal of $R$, and let $\prec$ be the graded reverse lexicographical ordering on $R$ with $x_{n}\prec\cdots\prec x_{1}$. Then, the following are equivalent:
   1. (1)
@@
   3. (3)
-    $\mathrm{Krull.dim}(R/\langle I,x_{n-r+1},\ldots,x_{n}\rangle)=0$.
+    $\mathrm{Krull.dim}(R/(I+\langle x_{n-r+1},\ldots,x_{n}\rangle))=0$.
   4. (4)
-    $\mathrm{Krull.dim}(R/\langle\langle\mathrm{LM}_{\prec}(I)\rangle,x_{n-r+1},\ldots,x_{n}\rangle)=0$.
-In the proof of [5, Lemma 4.1], the authors first proved the equivalence of the conditions (2), (3), and (4), and the implication (1) $\Rightarrow$ (2).
-The non-trivial part is that (2), (3), and (4) imply (1), which follows from [32, Proposition 5.4].
+    $\mathrm{Krull.dim}(R/(\langle\mathrm{LM}_{\prec}(I)\rangle+\langle x_{n-r+1},\ldots,x_{n}\rangle))=0$.
+In the proof of [6, Lemma 4.1], the authors first proved the equivalence of the conditions (2), (3), and (4), and the implication (1) $\Rightarrow$ (2).
+The non-trivial part is that (2), (3), and (4) imply (1), which follows from [40, Proposition 5.4].
 Here, we also recall the definition introduced by Hashemi [30] that a homogeneous ideal is in strong Noether position.
-Definition 2.1.6 ([23, Theorem 2.19]).
+Definition 2.1.6 ([30, Definition 2.4 and Theorem 2.19]).
 A homogeneous ideal $I$ of $R$ is said to be in strong Noether position (SNP) with respect to the ordered variables $x_{1},\ldots,x_{n}$ if we have $I^{\rm sat}=(I:x_{n}^{\infty})$ and
-$\langle I,x_{n-i+1},\ldots,x_{n}\rangle^{\rm sat}=(\langle I,x_{n-i+1},\ldots,x_{n}\rangle:x_{n-i}^{\infty})$
+${(I+\langle x_{n-i+1},\ldots,x_{n}\rangle)^{\rm sat}}=({(I+\langle x_{n-i+1},\ldots,x_{n}\rangle)}:x_{n-i}^{\infty})$
 for any $i$ with $11\leq i\leq r-$ 1.
-For equivalent definitions, see [23, Section 2] for details. Note that $I$ is in SNP if and only if $\langle\mathrm{LM}_{\prec}(I)\rangle$ is in SNP. Hashemi also proved the following inequality:
-Theorem 2.1.7 ([23, Proposition 4.7 and Theorem 4.17]).
+For equivalent definitions, see [30, Section 2] for details.
+Note that $I$ is in SNP if and only if $\langle\mathrm{LM}_{\prec}(I)\rangle$ is in SNP (cf.
+[30, Proposition 2.11]).
+Hashemi also proved the following inequality:
+Theorem 2.1.7 ([30, Proposition 4.7 and Theorem 4.17]).
 If $I$ is in SNP, then $\mathrm{reg}(I)$ is equal to
@@
 $\mathrm{max.GB.deg}_{\prec}(I)\leq\overline{\rm hilb}(I)$
-for the graded lexicographic ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$.
-Lemma 2.1.8 ([24, Lemma 23] and [25, Proposition 3.2 (v)]).
+for the graded reverse lexicographical ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$.
+Lemma 2.1.8 ([31, Lemma 23] and [32, Proposition 3.2 (v)]).
 If $r=1$, then the following are all equivalent:
@@
   2. (2)
-    $I$ is in quasi stable position with respect to the ordered variables $x_{1},\ldots,x_{n}$ (see e.g., [24, Definition 6] for definition).
+    $I$ is in quasi stable position with respect to the ordered variables $x_{1},\ldots,x_{n}$ (see e.g., [31, Definition 6] for definition).
   3. (3)
@@
 Proof.
-The equivalence of (1) and (2) is proved in [24, Lemma 23], and that of (2) and (3) follows from [25, Proposition 3.2 (v)]. ∎
+The equivalence of (1) and (2) is proved in [31, Lemma 23], and that of (2) and (3) follows from [32, Proposition 3.2 (v)]. ∎
 By Theorem 2.1.7 and Lemma 2.1.8, we obtain the following theorem:
-Theorem 2.1.9 ([24, Proof of Theorem 30]).
+Theorem 2.1.9 ([31, Proof of Theorem 30]).
 Assume that $r=1$.
-If $R/\langle I,x_{n}\rangle$ is Artinian (or one of the equivalent conditions given in Lemma 2.1.8 holds), then we have the inequality (1.0.1) for the graded reverse lexicographic ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$.
+If $R/(I+\langle x_{n}\rangle)$ is Artinian (or one of the equivalent conditions given in Lemma 2.1.8 holds), then we have the inequality (1.2) for the graded reverse lexicographical ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$.
+Here, we recall a fact that, if $K$ contains sufficiently many elements, the assumption $\mathrm{Krull.dim}(R/(I+\langle x_{n}\rangle))=0$ in Theorem 2.1.9 is satisfied after a suitable linear changer of coordinates.
+In the rest of this paper, we explain this fact in details, since we could not find any appropriate reference.
+Although it may be an easy consequence of the theory of hyperplane sections, there exists a hyperplane defined over $K$ that does not pass through any of given finite points in a projective space over $\overline{K}$ if $K$ is large enough.
+Here, we prove a somehow detailed argument:
+Lemma 2.1.10.
+Let $\bm{p}_{1},\ldots,\bm{p}_{N}$ be mutually distinct points in $\mathbb{P}^{n-1}(\overline{K})$.
+If the cardinality of $K$ is greater than $N$, then there exists a linear form $\ell\in R$ that does not vanish at any of $\bm{p}_{1},\ldots,\bm{p}_{N}$.
+Proof.
+For each point $\bm{p}_{k}=(p_{1,k}:\cdots:p_{n,k})\in\mathbb{P}^{n-1}(\overline{K})$, we denote by $S_{k}$ the set of all linear forms that vanish at $\bm{p}_{k}$, say
+$S_{k}=\{y_{1}x_{1}+\cdots+y_{n}x_{n}\mid y_{i}\in K\mbox{ and }y_{1}p_{1,k}+\cdots{+}y_{n}p_{n,k}=0\},$
+which is isomorphic, as $K$-linear spaces, to the following $K$-linear space of $K^{n}$:
+$S_{k}^{\prime}=\{(y_{1},\ldots,y_{n})\in K^{n}\mid{p_{1,k}y_{1}+\cdots{+}p_{n,k}y_{n}=0\}}.$
+The dimension of $S_{k}^{\prime}$ as a $K$-linear space is at most $n-1$.
+Indeed, for an arbitrary non-zero entry $p_{i,k}$ of $\bm{p}_{k}$, the vector with $1$ in the $i$-th entry and 00’s elsewhere does not belong to $S_{k}^{\prime}$, so that $S_{k}^{\prime}$ is a proper subspace of $K^{n}$.
+When $K$ is an infinite field, it follows from the irreducibility of an affine space over $K$ that $\bigcup_{k=1}^{N}S^{\prime}_{k}$ cannot cover $K^{n}$.
+When $K$ is a finite field of order $q$, we have
+#${\color[rgb]{0,0,0}\#\left(\bigcup_{k=1}^{N}S_{k}^{\prime}\right)\leq\sum_{k=1}^{N}\#S_{k}^{\prime}\leq N\cdot q^{n-1}<q^{n}},$
+whence a desired linear form over $K$ exists. ∎
+Lemma 2.1.11.
+With notation as above, assume that $\mathrm{Krull.dim}(R/I)\leq 1$.
+If the cardinality of $K$ is greater than #$\#V_{\overline{K}}(I)$, then there exists a linear form $\ell\in R$ such that $\mathrm{Krull.dim}(R/(I+\langle\ell\rangle))=0$.
+Moreover, there also exists a suitable linear coordinate change $\sigma$ sending $\ell$ to $x_{n}$ such that $\mathrm{Krull.dim}(R/(I^{\sigma}+\langle x_{n}\rangle))=0$.
+Proof.
+The first part follows from Lemma 2.1.10.
+There exists a $k\in\{1,\ldots,n\}$ such that the $x_{k}$-coefficient of $\ell$ is not zero.
+We may also assume that the $x_{k}$-coefficient is $1$, without loss of generality.
+Let $\sigma_{1}$ be the linear transformation of variables $x_{1},\ldots,x_{n}$ defined by a permutation exchanging $x_{k}$ and $x_{n}$, and write the image of $\ell$ by $\sigma_{1}$ as $\ell^{\sigma_{1}}=a_{0}x_{0}+a_{1}x_{1}+\cdots+a_{n}x_{n}$ for $a_{i}\in K$ with $a_{n}=1$.
+Then, let $\sigma_{2}$ be a linear transformation of variables $x_{1},\ldots,x_{n}$ defined by
+$h^{\sigma_{2}}:=h(x_{1},\ldots,x_{n-1},x_{n}-a_{1}x_{1}-\cdots-a_{n-1}x_{n-1})$
+for $h\in R$.
+Then, the invertible transformation $\sigma:=\sigma_{2}\circ\sigma_{1}$ of variables $x_{1},\ldots,x_{n}$ sends $\ell$ to $x_{n}$, so that $\mathrm{Krull.dim}(R/(I^{\sigma}+\langle x_{n}\rangle))=0$.
+∎
+Remark 2.1.12.
+In Lemma 2.1.11, we have an exact sequence
+${(R/I)_{d^{\prime}-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\times\ell}$${(R/I)_{d^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(R/(I+\langle\ell\rangle))_{d^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
+for each $d^{\prime}\geq 1$.
+Therefore, the minimal $d$ for which the $-\ell$ map $(R/I)_{d^{\prime}-1}\to(R/I)_{d^{\prime}}$ is surjective for any $d^{\prime}\geq d$ is equal to $d_{\rm reg}(I+\langle\ell\rangle)$.
 2.2 Generalized semi-regular sequences
@@
 Throughout this subsection, let $f_{1},\ldots,f_{m}\in R$ be homogeneous polynomials of positive degrees $d_{1},\ldots,d_{m}$, and put $I=\langle f_{1},\ldots,f_{m}\rangle$.
-Theorem 2.2.1 (cf. [14, Theorem 1]).
+Theorem 2.2.1 (cf. [15, Theorem 1]).
 For each $d$ with $d\geq\mathrm{max}\{d_{i}:1\leq i\leq m\}$, the following are equivalent:
@@
     We have $H_{1}(K_{\bullet}(f_{1},\ldots,f_{m}))_{\leq d-1}=0$, where $K_{\bullet}(f_{1},\ldots,f_{m})$ denotes the Koszul complex on the sequence $(f_{1},\ldots,f_{m})$.
-Definition 2.2.2 ([1, Definition 3]; see also [14, Definition 1]).
+Definition 2.2.2 ([2, Definition 3]; see also [15, Definition 1]).
 For each $d$ with $d\geq\mathrm{max}\{d_{i}:1\leq i\leq m\}$, a sequence $(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R\smallsetminus K$ is said to be $d$-regular if the equivalent conditions in Theorem 2.2.1 hold.
 Here, we define a cryptographic semi-regular sequence:
-Definition 2.2.3 ([1, Definition 5], [3, Definition 5]; see also [14, Section 2]).
+Definition 2.2.3 ([2, Definition 5], [4, Definition 5]; see also [15, Section 2]).
 A sequence $(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R\smallsetminus K$ is said to be cryptographic semi-regular if it is $d_{\rm reg}(I)$-regular.
 The Hilbert series of $R/I$ for a cryptographic semi-regular sequence $(f_{1},\ldots,f_{m})$ is easily computed, by the following Proposition 2.2.4:
-Proposition 2.2.4 ([14, Proposition 1 (d)]; see also [3, Proposition 6]).
-With the same notation as in Theorem 2.2.1, a sequence $(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R\smallsetminus K$ is cryptograpic semi-regular if and only if
+Proposition 2.2.4 ([15, Proposition 1 (d)]; see also [4, Proposition 6]).
+With the same notation as in Theorem 2.2.1, a sequence $(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R\smallsetminus K$ is cryptographic semi-regular if and only if
 ${\rm HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}(z)=\left[\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\right],$ (2.2.2)
@@
 The above ‘standard’ form (2.2.2) of Hilbert series was introduced by Fröberg in his celebrated paper [23]. We will recall his famous conjecture (Conjecture 2.3.7 below) in the next subsection.
-Here, we recall the following lemma:
-Lemma 2.2.5 ([20, b) on page 121]).
+Remark 2.2.5.
+In [25, Definition 1.3], Fröberg-Hollman already defined a sufficiently generic ideal as one generated by homogeneous polynomials $f_{1},\ldots,f_{m}\in R$ satisfying (2.2.2).
+Namely, a cryptographic semi-regular sequence is nothing but a sequence of homogeneous polynomials generating a sufficiently generic ideal.
+Here, we recall the following lemma by Fröberg:
+Lemma 2.2.6 ([23, b) on page 121]).
 With notation as above, we have
 ${\rm HS}_{R/I}(z)\geq^{\rm lex}\left[\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\right]$
-in the lexicographic sense, i.e., the first coefficient that differs, if any, is greater on the left hand side, see [20, page 118].
+in the lexicographic sense, i.e., the first coefficient that differs, if any, is greater on the left hand side, see [23, page 118].
 Remark 2.2.7.
@@
 $\left({\rm HS}_{R/I}(z)\bmod{z^{d}}\right)\geq^{\rm lex}\left(\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\bmod{z^{d}}\right).$
-To deal with the case where $R/I$ has Krull dimension one, we extend the notion of cryptgraphic semi-regular as follows:
-Definition 2.2.7 ([29, Definition 2.3.3]).
+To deal with the case where $R/I$ has Krull dimension one, we extend the notion of cryptographic semi-regular as follows:
+Definition 2.2.8 ([36, Definition 2.3.3]).
 A sequence $(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R\smallsetminus K$ is said to be generalized cryptographic semi-regular if it is $\widetilde{d}_{\rm reg}(I)$-regular.
+Remark 2.2.9 (cf. [36, Section 4.1]).
+A generalized cryptographic semi-regular (but neither regular nor semi-regular) sequence is easily generated.
+For example, choosing coefficients of $g_{1},\ldots,g_{m}$ with $m\geq n$ uniformly at random from a finite field $K=\mathbb{F}_{q}$ and choosing an arbitrary point $\bm{p}=(p_{1},\ldots,p_{n-1},p_{n})\in\mathbb{A}^{n}(K)$ with $p_{n}=1$, then it is expected that $(f_{1},\ldots,f_{m})$ with $f_{j}:=g_{j}(x_{1},\ldots,x_{n})-g_{j}(p_{1},\ldots,p_{n})x_{n}^{d_{j}}$ is generalized cryptographic semi-regular (but neither regular nor semi-regular).
+Without loss of generality, we may take $\bm{o}=(0:\cdots:0:1)$ as $\bm{p}$, see Section 4 below for details.
+Example 2.2.10.
+Let $K=\mathbb{F}_{31}$, $n=3$, and $R=K[x_{1},x_{2},x_{3}]$. Here we consider the case where $(d_{1},d_{2},d_{3})=(3,2,2)$ with $m=3$, and let
+$f_{1}$ $=$ $27x_{1}^{3}+5x_{1}^{2}x_{2}+6x_{1}x_{2}^{2}+20x_{2}^{3}+29x_{1}x_{2}x_{3}+25x_{2}^{2}x_{3}+3x_{1}x_{3}^{2}+26x_{2}x_{3}^{2},$
+$f_{2}$ $=$ $2x_{1}^{3}+29x_{1}^{2}x_{2}+25x_{1}x_{2}^{2}+3x_{2}^{3}+25x_{1}^{2}x_{3}+20x_{1}x_{2}x_{3}+25x_{2}^{2}x_{3}+19x_{1}x_{3}^{2}+24x_{3}x_{3}^{2},$
+$f_{3}$ $=$ $12x_{1}^{2}+12x_{1}x_{2}+23x_{2}^{2}+x_{1}x_{3}+10x_{2}x_{3}.$
+Putting $I=\langle f_{1},f_{2},f_{3}\rangle$, one can verify that
+${\rm HS}_{R/I}(z)=1+3z+5z^{2}+5z^{3}+3z^{4}+z^{5}+z^{6}+\cdots$
+with $\mathrm{Krull.dim}(R/I)=1$, $D:=\mathrm{hilb}(I)=5$, and $\dim_{K}(R/I)_{D}=1$, so that $(f_{1},f_{2},f_{3})$ is $D$-regular (in fact $D+1$-regular).
+Namely $(f_{1},f_{2},f_{3})$ is generalized cryptographic semi-regular, but is neither regular nor cryptographic semi-regular.
 Lemma 2.2.11.
@@
 By our assumption that $\bm{F}$ is generalized cryptographic semi-regular, the equality (2.2.1) holds for $d=\widetilde{d}_{\rm reg}(I)$.
-For any $d^{\prime}$ with $d^{\prime}<\widetilde{d}_{\rm reg}(I)$, we have ${\rm HF}_{R/I}(d^{\prime})>0$, whence the $z^{d^{\prime}}$-coefficient of $\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}$ is also positive.
+For any $d^{\prime}$ with $d^{\prime}<\widetilde{d}_{\rm reg}(I)$, we have ${\rm HF}_{R/I}(d^{\prime})>0$, whence the $z^{d^{\prime}}$-coefficient of $(\prod_{j=1}^{m}(1-z^{d_{j}}))/(1-z)^{n}$ is also positive.
 Thus, the inequality (2.2.3) holds.
@@
 For a sequence $\bm{F}=(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R\smallsetminus K$, we consider the following four conditions:
-$\bm{F}$ is regular, equivalently
-${\rm HS}_{R/I}(z)=\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}.$
+$\bm{F}$ is regular, equivalently ${\rm HS}_{R/I}(z)=(\prod_{j=1}^{m}(1-z^{d_{j}}))/(1-z)^{n}$.
 $\bm{F}$ is cryptographic semi-regular (i.e., $d_{\rm reg}(I)$-regular).
@@
 For $m\leq n$, Proposition 2.2.4 implies $(B)\Rightarrow(C)$.
+On the other hand, it is clear that (B) does not imply (A) if $m>n$.
+The sequence $(f_{1},f_{2},f_{3})$ provided in Example 2.2.10 gives a counterexample for the implication (C) $\Rightarrow$ (B) when $r=1$.
+As for (D) $\Rightarrow$ (C), here is a counterexample with $r=2$: Let $(g_{1},g_{2}):=(x^{3},xy)\in\mathbb{Q}[x,y,z]^{2}$ and put $I:=\langle g_{1},g_{2}\rangle$.
+Then one can verify that $\mathrm{HS}_{R/I}(z)=1+3z+5z^{2}+6z^{3}+7z^{4}+\cdots$ with $\mathrm{Krull.dim}(R/I)=2$ and $D:=\mathrm{hilb}(I)=2$, whence $(g_{1},g_{2})$ is $D$-regular but not regular.
 ∎
-2.3 Genericness and a variant of Fröberg’s conjecture
-Given $m$ non-negative integers $d_{1},\ldots,d_{m}$, let $V(n,m,d_{1},\ldots,d_{m})\!=\!R_{d_{1}}\times\cdots\times R_{d_{m}}$, which is topologically identified with the affine space $\mathbb{A}^{N_{1}}(K)\times\dots\times\mathbb{A}^{N_{m}}(K)$ of dimension $\sum_{j=1}^{m}N_{j}$ with $N_{j}:=\binom{n+d_{j}-1}{d_{j}}$.
-Here the identification from $R_{d}$ to $\mathbb{A}^{N}$, where $N=\binom{n+d-1}{d}$, is given as follows: Letting $\{t_{1},\ldots,t_{N}\}$ be the set of all monomials in $R$ of degree $d$, each polynomial $f=\sum_{i=1}^{N}c_{i}t_{i}$ in $R_{d}$ is identified with its coefficients vector $(c_{1},\ldots,c_{N})$.
-We write simply $V$ for $V(n,m,d_{1},\ldots,d_{m})$ when we fix the parameter $(n,m,d_{1},\ldots,d_{m})$.
-Definition 2.3.1 (cf. [34, Section 1]).
-For given $n$, $m$, and $(d_{1},\ldots,d_{m})$, a property $\mathcal{P}$ for $\bm{F}=(f_{1},\ldots,f_{m})\in V$ is said to be generic (or, we say that a generic sequence in $V$ satisfies $\mathcal{P}$) if it holds on a non-empty Zariski open set in $V$, namely there exists a non-empty Zariski open set $U$ in $V$ such that
-$U\subset\{\bm{F}=(f_{1},\ldots,f_{m})\in V:\mbox{$\bm{F}$ satisfies $\mathcal{P}$}\}.$
+2.3 Genericness and Fröberg’s conjecture
+Given $m$ positive integers $d_{1},\ldots,d_{m}$, let ${V_{n,m,(d_{1},\ldots,d_{m})}}:=R_{d_{1}}\times\cdots\times R_{d_{m}}$, which is topologically identified with the affine space $\mathbb{A}^{N_{1}}(K)\times\dots\times\mathbb{A}^{N_{m}}(K)=\mathbb{A}^{N}(K)$ of dimension $N:=\sum_{j=1}^{m}N_{j}$ with $N_{j}:=\binom{n+d_{j}-1}{d_{j}}$.
+Here, we identify a form in $R_{d_{j}}$ with a point in $\mathbb{A}^{N_{j}}(K)$ by taking the coefficients.
+We may write $K^{N}$ simply for $\mathbb{A}^{N}(K)$.
+We also correspond a point $\bm{F}=(f_{1},\ldots,f_{m})\in V_{n,m,(d_{1},\ldots,d_{m})}$ to an ideal $\langle\bm{F}\rangle_{R}:=\langle f_{1},\ldots,f_{m}\rangle_{R}$, where the subscript $R$ may be omitted if the meaning is clear from the context.
+Following Fröberg’s definition (cf.
+[23], [24], [27]), we say that a form of $R_{d}$ is generic if all monomials of degree $d$ occur with nonzero coefficient and if all the coefficients are algebraically independent over the prime field of $K$.
+We call $\bm{F}=(f_{1},\ldots,f_{m})\in V_{n,m,(d_{1},\ldots,d_{m})}$ a generic sequence if all $f_{i}$’s are generic forms with coefficients that are mutually algebraic independent over the prime field.
+If $K$ does not contain $N$ elements that are algebraically independent over the prime field of $K$ (for example $K$ is an algebraic extension of the prime field), we may replace $K$ by its arbitrary extension field $L$ containing such $N$ elements.
+When we consider a property of Hilbert series, this makes a sense since every Hilbert series is not changed by extending the field of coefficients.
+On the other hand, the genericness of a property for $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ is defined as follows:
+Definition 2.3.1 (cf. [43, Section 1]).
+For given $n$, $m$, and $(d_{1},\ldots,d_{m})$, a property $\mathcal{P}$ for $\bm{F}=(f_{1},\ldots,f_{m})\in V_{n,m,(d_{1},\ldots,d_{m})}$ is said to be generic if it holds on a non-empty Zariski open set in $V_{n,m,(d_{1},\ldots,d_{m})}$, namely there exists a non-empty Zariski open set $U$ in $V_{n,m,(d_{1},\ldots,d_{m})}$ such that
+$U\subset\{\bm{F}=(f_{1},\ldots,f_{m})\in V_{n,m,(d_{1},\ldots,d_{m})}:\mbox{$\bm{F}$ satisfies $\mathcal{P}$}\}.$
+Here, we recall Fröberg’s important lemma for the minimality of the Hilbert series on generic sequences.
+Note that the characteristic of $K$ is assumed to be 00 in the original paper [23], but it can be easily proved that the assertion of the lemma holds for any characteristic.
+We here state the lemma in a slightly modified (but equivalent) form for the convenience of our use.
+Lemma 2.3.2 ([23, Lemma 1]).
+With notation as above, let $\Omega$ be an extension field of $K$, and let $\mathcal{C}:=\{c_{j,t}:1\leq j\leq m,\ t\in\mathcal{T}_{d_{j}}\}$ be a set of elements in $\Omega$ that are algebraically independent over the prime field $K_{0}$ of $K$.
+Here $\mathcal{T}_{d}$ is the set of monomials of degree $d$ in $\{x_{1},\ldots,x_{n}\}$.
+Let $L:=K(\mathcal{C})$, let $R_{L}:=R\otimes_{K}L=L[x_{1},\ldots,x_{n}]$, and put $g_{j}:=\sum_{t\in\mathcal{T}_{d_{j}}}c_{j,t}t\in(R_{L})_{d_{j}}$ for each $1\leq j\leq m$.
+Namely $\bm{G}=(g_{1},\ldots,g_{m})$ is a generic sequence.
+Then, for any $\bm{F}=(f_{1},\ldots,f_{m})\in V_{n,m,(d_{1},\ldots,d_{m})}$, we have the coefficient-wise inequality
+$\mathrm{HS}_{R/\langle f_{1},\ldots,f_{m}\rangle_{R}}(z)=\mathrm{HS}_{R_{L}/\langle f_{1},\ldots,f_{m}\rangle_{R_{L}}}\geq\mathrm{HS}_{R_{L}/\langle g_{1},\ldots,g_{m}\rangle_{R_{L}}}(z).$ (2.3.1)
 Remark 2.3.3.
-Another definition of genericness is the following (cf.
-[20, page 119]): For given $n$, $m$, and $(d_{1},\ldots,d_{m})$, a property $\mathcal{P}$ for $\bm{F}=(f_{1},\ldots,f_{m})\in V$ is said to be generic over $K$ if it holds for any $\bm{F}=(f_{1},\ldots,f_{m})\in V$ such that the coefficients of $f_{1},\ldots,f_{m}$ are algebraically independent over the prime field $L$ of $K$, that is, $L=\mathbb{Q}$ or $\mathbb{F}_{p}$ for some prime number $p$.
-It can be checked that the two definitions of genericness are equivalent.
+Since every Hilbert series is not changed by extending the field of coefficients, the Hilbert series of the quotient by an ideal generated by a generic sequence depends only on $\mathrm{char}(K)$, $n$, $m$, and $(d_{1},\ldots,d_{m})$, see e.g., [16, Lemma and Definition 4].
+Thus, in discussing the Hilbert series, we may consider the ideal generated by $\bm{G}$ over $K_{0}(\mathcal{C})[X]$ instead of that over $K(\mathcal{C})[X]$.
+Because, the reduced Gröbner basis of $\langle\bm{G}\rangle_{R_{L}}$ coincides with that of $\langle\bm{G}\rangle_{K_{0}(\mathcal{C})[X]}$, since the well-known Buchberger’s algorithm can be performed over a field, over which the generating set is defined.
+Hence the minimal generating set of ${\rm in}(\langle\bm{G}\rangle_{R_{L}})$ coincides with that of ${\rm in}(\langle\bm{G}\rangle_{K_{0}(\mathcal{C})[X]})$.
+This also implies that, for computing the reduced Gröbner basis of $\langle\bm{G}\rangle_{R_{L}}$, we may assume that all elements of $\mathcal{C}$ are algebraically independent over $K$ and compute it over $L=K(\mathcal{C})$.
+If $K$ is an infinite field, Fröberg-Löfwall proved in [26, Theorem 1] that there exists a non-empty Zariski-open subset $U_{\rm min}\subset V_{n,m,(d_{1},\ldots,d_{m})}$ on which the quotients by the corresponding ideals have the same Hilbert series, which is minimal in the coefficient-wise sense.
+Note that, even if $K$ contains the set $\mathcal{C}$ of $N$ algebraically independent elements, $\bm{F}\in U_{\rm min}$ not necessarily implies the genericness of $\bm{F}$.
+On the other hand, the Hilbert series of $R/\langle\bm{F}\rangle$ for $\bm{F}\in U_{\rm min}$ is equal to that of $R_{L}/\langle\bm{G}\rangle$ for a generic sequence $\bm{G}$ over $L$ as in Lemma 2.3.2.
+The existence of the non-empty Zariski open set $U_{\rm min}$ can be easily verified by considering the stability of the reduced Gröbner basis in the context of so-called a comprehensive Gröbner system for a parametric ideal.
+A result by Nabeshima in [42] on “generic Gröbner basis of a parametric ideal” can be translated directly to our problem where the generic sequence $\bm{G}$ is given from $K[\mathcal{C}][x_{1},\ldots,x_{n}]$ and $\mathcal{C}$ is the set of coefficient parameters.
+Here we assume that all elements of $\mathcal{C}$ are algebraically independent over $K$ and hence they are treated as independent indeterminates.
+(See Remark 2.3.3.) In the below, we show a useful result for our purpose.
+We let $\prec_{1}$ be a block ordering with $X:=\{x_{1},\ldots,x_{n}\}\succ\succ\mathcal{C}$ such that its restriction on $X$ coincides with the given ordering $\prec$.
+Let $\bm{H}$ be a finite set of $K[\mathcal{C}][X]$ and ${\rm GB}(\bm{H})$ the reduced Gröbner basis of the ideal $\langle\bm{H}\rangle_{L[x]}$ with respect to the restriction $\prec$ of $\prec_{1}$ on the monomials in $X$, where $L=K(\mathcal{C})$.
+Moreover, let $\widetilde{\rm GB}(\bm{H}):=\{{\rm dlcm}(h)h:h\in{\rm GB}(\bm{H})\}$ and $\ell(\bm{H})=\prod_{h^{\prime}\in\widetilde{\rm GB}(\bm{H})}{\rm LC}_{\prec}(h^{\prime})$ in $K[\mathcal{C}][X]$, where ${\rm dlcm}(h)$ denotes the least common multiple of all denominators of coefficients in $K(\mathcal{C})$ of $h$ and each $h^{\prime}\in\widetilde{\rm GB}(\bm{H})$ is considered as an element in $K[\mathcal{C}][X]$, that is, a polynomial in $X$ over $K[\mathcal{C}]$.
+For $\bm{a}\in\mathbb{A}^{N}(K)$, we denote by $\pi_{\bm{a}}$ the substitution homomorphism from $K[\mathcal{C}][X]$ to $R=K[X]$, that is, $\pi_{\bm{a}}(h(\mathcal{C},X))=h(\bm{a},X)$ for $h\in K[\mathcal{C},X]$.
+For $h\in L[X]$, when the denominator of each coefficient of $h$ does not vanish at $\bm{a}$, the image $\pi_{\bm{a}}(h)$ is defined naturally.
+Theorem 2.3.4 ([42, Theorem 3]).
+With notation as above, consider $\bm{H}$ and $\widetilde{\rm GB}(\bm{H})$ to be subsets of $K[\mathcal{C}\cup X]$, and let $S(\bm{H})$ be the reduced Gröbner basis of the ideal quotient $\langle\bm{H}\rangle_{K[\mathcal{C}\cup X]}:\langle\widetilde{\rm GB}(\bm{H})\rangle_{K[\mathcal{C}\cup X]}$ with respect to $\prec_{1}$.
+Then, we have $S(\bm{H})\cap K[\mathcal{C}]\not=\emptyset$, and for any $\bm{a}$ in $\mathbb{A}^{N}(\overline{K})\smallsetminus(V_{\overline{K}}(S(\bm{H})\cap K[\mathcal{C}])\cup V_{\overline{K}}(\ell(\bm{H})))$, the set $\pi_{\bm{a}}({\rm GB}(\bm{H}))$ is the reduced Gröbner basis of $\langle\pi_{\bm{a}}(\bm{H})\rangle_{K[X]}$ with respect to $\prec$.
+Now we return to our case where $\bm{H}=\bm{G}$.
+Let $\mathcal{O}=\mathbb{A}^{N}(K)\smallsetminus(V_{K}(S(\bm{G})\cap K[\mathcal{C}])\cup V_{K}(\ell(\bm{G})))$.
+Then $\mathcal{O}$ is Zariski-open in $\mathbb{A}^{N}(K)$.
+Moreover, $\mathcal{O}$ is not empty, since $\langle S(\bm{G})\cap K[C]\rangle\cap\langle\ell(\bm{G})\rangle$ is a non-zero ideal and since $K$ is an infinite field.
+It is obvious that for any $\bm{a}\in\mathcal{O}$, $\pi_{\bm{a}}({\rm GB}(\bm{G}))$ is the reduced Gröbner basis of $\langle\pi_{\bm{a}}(\bm{G})\rangle_{R}$ with respect to $\prec$.
+Moreover, for such $\bm{a}$, since ${\rm LC}_{\prec}(\pi_{\bm{a}}(h))=1$ for any $h\in{\rm GB}(\bm{G})$, we have ${\rm LM}_{\prec}(\pi_{\bm{a}}({\rm GB}(\bm{G})))={\rm LM}_{\prec}({\rm GB}(\bm{G}))$.
+Corollary 2.3.5.
+Under the assumption in Theorem 2.3.4, there is a non-empty Zariski open set $\mathcal{O}$ in $K^{N}$ such that, for any $\bm{a}$ in $\mathcal{O}$, the set of leading monomials of the reduced Gröbner basis of $\langle\pi_{\bm{a}}(\bm{G})\rangle_{R}$ coincides with that of $\langle\bm{G}\rangle_{R_{L}}$.
+For each ideal $I$ of $K[X]$, the Hilbert series ${\rm HS}_{R/I}(z)$ is determined by its initial ideal and so by the leading monomials of its (reduced) Gröbner basis.
+Thus, Corollary 2.3.5 implies that ${\rm HS}_{R/\langle\pi_{\bm{a}}(\bm{G})\rangle}(z)$ coincides with ${\rm HS}_{R_{L}/\langle\bm{G}\rangle}(z)$ for any $\bm{a}\in\mathcal{O}$.
+Considering the corresponding set $U$ of $\mathcal{O}$ in $V_{n,m,(d_{1},\ldots,d_{m})}$, we have the following proposition and a conjecture.
+Here, we remark that the following proposition may be a well-known fact.
+However, the authors could not find any reference for its proof, and therefore they would write a proof here for the readers’ convenience; the idea of the proof will be used to obtain an analogous result in Section 4 below.
+Proposition 2.3.6.
+Assume that $K$ is an infinite field. Then, for given $n$, $m$, and $(d_{1},\ldots,d_{m})$, the following conditions (1), (2), (3), and (4) are equivalent to each other:
+  1. (1)
+    There exists an $\bm{F}=(f_{1},\ldots,f_{m})\in V_{n,m,(d_{1},\ldots,d_{m})}$ satisfying (2.2.2), i.e., $\bm{F}$ is cryptographic semi-regular (in other words, the ideal $\langle\bm{F}\rangle$ is sufficiently generic).
+  2. (2)
+    For the non-empty Zariski open set $U_{\rm min}$ defined in the paragraph just after Remark 2.3.3, any $\bm{F}=(f_{1},\ldots,f_{m})\in U_{\rm min}$ satisfies (2.2.2).
+  3. (3)
+    The property that $\bm{F}=(f_{1},\ldots,f_{m})\in V_{n,m,(d_{1},\ldots,d_{m})}$ satisfies (2.2.2) is generic (in terms of Definition 2.3.1).
+  4. (4)
+    For $\mathcal{C}$, $L$, $R_{L}$, and $\bm{G}=(g_{1},\ldots,g_{m})$ as in Lemma 2.3.2, we have
+    $\mathrm{HS}_{R_{L}/\langle g_{1},\ldots,g_{m}\rangle}(z)=\left[\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\right].$ (2.3.2)
+    Namely, the generic sequence $\bm{G}$ is cryptographic semi-regular (in other words, the ideal $\langle\bm{G}\rangle_{R_{L}}$ is sufficiently generic).
+Proof.
+The implication (2) $\Rightarrow$ (1) is obvious.
+To see (1) $\Rightarrow$ (4), it follows from Lemma 2.2.6 and Lemma 2.3.2 that
+$\mathrm{HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}(z)\geq\mathrm{HS}_{R_{L}/\langle g_{1},\ldots,g_{m}\rangle}(z)\geq^{\rm lex}\left[\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\right].$ (2.3.3)
+By our assumption that $\mathrm{HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}(z)$ satisfies (2.2.2), we have (2.3.2).
+The implication (4) $\Rightarrow$ (3) follows from Corollary 2.3.5.
+The implication (3) $\Rightarrow$ (2) follows from the irreducibility of $V_{n,m,(d_{1},\ldots,d_{m})}$. ∎
 Here, we recall Fröberg’s conjecture:
 Conjecture 2.3.7 (Fröberg’s conjecture [23]).
-Assume that $K$ is an infinite field.
-For given $n$, $m$, and $(d_{1},\ldots,d_{m})$, the property that $\bm{F}=(f_{1},\ldots,f_{m})\in V$ satisfies (2.2.2) (i.e., $\bm{F}$ is cryptographic semi-regular) is generic.
-In other words, a generic sequence $\bm{F}=(f_{1},\ldots,f_{m})\in V$ is cryptographic semi-regular.
+Assume that $K$ is an infinite field. For any $n$, $m$, and $(d_{1},\ldots,d_{m})$, the equivalent conditions (1), (2), (3), and (4) in Proposition 2.3.6 hold.
 See Pardue’s paper [43] for some conjectures equivalent to Fröberg’s one (Conjecture 2.3.7).
-He also proved that Moreno-Socías conjecture [33] is stronger than Conjecture 2.3.3, see [34, Theorem 2] for a proof.
-Here, as a variant of Conjecture 2.3.3, we raise the following conjecture (cf. [29, Conjecture 4.3.4] as its strict version):
-Conjecture 2.3.4.
-Assume that $K$ is an infinite field. For given $n$, $m$, and $(d_{1},\ldots,d_{m})$, the property that $\bm{F}=(f_{1},\ldots,f_{m})\in V$ is generalized cryptographic semi-regular is generic.
-We shall prove Conjectures 2.3.3 and 2.3.4 are equivalent to each other. For this, we first prove the following lemma:
+He also proved that Moreno-Socías conjecture [41] is stronger than Conjecture 2.3.7, see [43, Theorem 2] for a proof.
+3 Proofs of Theorems 1.1 and 1.2
+As in the previous sections, let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$ (in any characteristic).
+Let $I$ be a proper homogeneous ideal of $R$, and let $F=\{f_{1},\ldots,f_{m}\}$ be a homogeneous subset of $R$ generating $I$.
+For each $1\leq j\leq m$, put $d_{j}:=\deg(f_{j})$ (total degree).
+3.1 Proof of Theorem 1.1
+We first prove Theorem 1.1 (1), which can be easily proved by Lemma 2.2.11 together with the following easy lemma in the case where $m=n-1$:
 Lemma 3.1.1.
-For $m\geq n$, the property that $R/\langle f_{1},\ldots,f_{m}\rangle$ is Artinian is generic.
-Proof.
-The case $m=n$ is well-known to hold. Assume $m>n$, fix $n$, $m$, and $(d_{1},\ldots,d_{m})$, and put
-$\Lambda^{\rm Art}:=\{\bm{F}=(f_{1},\ldots,f_{m})\in V:\mbox{$R/\langle f_{1},\ldots,f_{m}\rangle$ is Artinian}\}.$
-Consider a natural projection
-$V\to V_{n}:=\mathbb{A}^{N_{1}}(K)\times\dots\times\mathbb{A}^{N_{n}}(K)\ ;\ (f_{1},\ldots,f_{n},f_{n+1},\ldots,f_{m})\mapsto(f_{1},\ldots,f_{n}).$
-Then we have $\pi^{-1}(\Lambda_{n}^{\rm Art})\subset\Lambda^{\rm Art}$, where
-$\Lambda_{n}^{\rm Art}:=\{\bm{F}=(f_{1},\ldots,f_{n})\in V:\mbox{$R/\langle f_{1},\ldots,f_{n}\rangle$ is Artinian}\}$
-is Zariski-open in $V_{n}$. Since $\pi$ is continuous, $\pi^{-1}(\Lambda_{n}^{\rm Art})$ is Zariski-open in $V$, whence the assertion holds. ∎
-Proposition 2.3.6.
-Assume that $K$ is an infinite field. For $m\geq n$, our conjecture (Conjecture 2.3.4) is equivalent to Fröberg’s conjecture (Conjecture 2.3.3).
-Proof.
-Fixed $n$, $m$, and $(d_{1},\ldots,d_{m})$, we set
-[…]
-[…]
-so that $\Lambda^{\rm csr}=\Lambda^{\rm gcsr}\cap\Lambda^{\rm Art}$.
-Assume that $(f_{1},\ldots,f_{m})\in V$ is generically generalized cryptograhic semi-regular, i.e., there exists a non-empty Zariski-open set $U_{1}\subset V$ such that $U_{1}\subset\Lambda^{\rm gcsr}$.
-Recall from Lemma 2.3.5 that $(f_{1},\ldots,f_{m})\in V$ generically has the coordinate ring $R/\langle f_{1},\ldots,f_{m}\rangle$ that is Artinian, whence there exists a non-empty Zariski-open set $U_{2}\subset V$ such that $U_{2}\subset\Lambda^{\rm Art}$.
-Since both $U_{1}$ and $U_{2}$ are dense, we have $\emptyset\neq U_{1}\cap U_{2}\subset\Lambda^{\rm gcsr}\cap\Lambda^{\rm Art}=\Lambda^{\rm csr}$.
-Hence, $(f_{1},\ldots,f_{m})\in V$ is cryptographic semi-regular on the non-empty Zariski open set $U_{1}\cap U_{2}$ of $V$, as desired.
-The converse is obvious by $\Lambda^{\rm csr}\subset\Lambda^{\rm gcsr}$. ∎
-Despite of the equivalence between two conjectures, those cannot guarantee the genericness of $\Lambda^{\rm gscr}\cap(V\smallsetminus\Lambda^{\rm Art})$ with respect to the induced topology of $V\smallsetminus\Lambda^{\rm Art}$.
-Once the genericness holds for the case, we have the following.
-Remark 2.3.7.
-If the subset $\Lambda^{\rm gcsr}\cap(V\smallsetminus\Lambda^{\rm Art})$ contains a non-empty Zariski-open set in the subspace $V\smallsetminus\Lambda^{\rm Art}$ equipped with the induced topology, then ‘almost’ all sequences $(f_{1},\ldots,f_{m})$ such that $R/\langle f_{1},\ldots,f_{m}\rangle$ has Krull dimension $\geq 1$ are expected to be generalized cryptographic semi-regular.
-As the first discussion about the genericness of $\Lambda^{\rm gcsr}\cap(V\smallsetminus\Lambda^{\rm Art})$, we consider the following sub case: Let $Z$ be the non-empty Zariski closed set in $V=V(n,m,d_{1},\ldots,d_{m})$ defined by the condition that the $x_{n}^{d_{i}}$-coefficient of $f_{i}$ is zero for any $i$ with $1\leq i\leq m$.
-Then, for any element $\bm{F}=(f_{1},\ldots,f_{m})\in Z$, the homogeneous ideal $\langle f_{1},\ldots,f_{m}\rangle_{R}$ has a projective root $(0:\cdots:0:1)$, whence $R/\langle f_{1},\ldots,f_{m}\rangle$ is not Artinian.
-Namely, we have $Z\subset V\smallsetminus\Lambda^{\rm Art}$.
-Lemma 2.3.8.
-Let $\bm{G}=(g_{1},\ldots,g_{m})\in Z$, and assume that the other coefficients of $g_{1},\ldots,g_{m}$ are algebraically independent over the prime field of $K$.
-Then, for any $\bm{F}=(f_{1},\ldots,f_{m})\in Z$, we have the coefficient-wise inequality
-${\rm HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}(z)\geq{\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}(z).$
-Therefore, if there exists a generalized cryptographic semi-regular sequence $\bm{F}=(f_{1},\ldots,f_{m})$ in $Z$ with $\dim_{K}(R/\langle f_{1},\ldots,f_{m}\rangle)_{D}=1$ for $D=\tilde{d}_{\rm reg}(\langle f_{1},\ldots,f_{m}\rangle_{R})$, then $\bm{G}=(g_{1},\ldots,g_{m})$ is generalized cryptographic semi-regular, i.e.,
-${\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}(z)\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{D^{\prime}}}$
-for $D^{\prime}=\widetilde{d}_{\rm reg}(\langle g_{1},\ldots,g_{m}\rangle_{R})$, and $\dim_{K}(R/\langle g_{1},\ldots,g_{m}\rangle)_{D^{\prime}}=1$.
-Proof.
-The first half can be proved in a way similar to the proof of [20, Lemma 1].
-To prove the second one, we also use Lemma 2.2.5 (cf.
-[20]) for a sequence $\bm{H}=(h_{1},\ldots,h_{m})$ of generic forms with the same parameters $n,m,d_{1},\ldots,d_{m}$.
-Then, we have ${\rm HS}_{R/\langle\bm{G}\rangle}\geq{\rm HS}_{R/\langle\bm{H}\rangle}$ by [20, Lemma 1] and thus,
-${\rm HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}\geq{\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}\geq{\rm HS}_{R/\langle h_{1},\ldots,h_{m}\rangle}\geq^{\rm lex}\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}\right],$ (2.3.1)
-where $\geq^{\rm lex}$ denotes the lexicographical inequality. As $\bm{F}$ is a generalized cryptographic semi-regular sequence, we also have
-${\rm HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{D}}$ (2.3.2)
-for $D=\tilde{d}_{\rm reg}(\langle f_{1},\ldots,f_{m}\rangle_{R})$. Combining (2.3.1) and (2.3.2), we have
-${\rm HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}\equiv{\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{D}}.$
-Meanwhile, it follows from the condition $\dim_{K}(R/\langle f_{1},\ldots,f_{m}\rangle)_{D}=1$ that we obtain $\dim_{K}(R/\langle g_{1},\ldots,g_{m}\rangle)_{D}=1$ and $\tilde{d}_{\rm reg}(\langle g_{1},\ldots,g_{m}\rangle_{R})\leq D$.
-Hence we have
-${\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{D^{\prime}}}$
-for $D^{\prime}=\tilde{d}_{\rm reg}(\langle g_{1},\ldots,g_{m}\rangle_{R})$. ∎
-From Lemma 2.3.8 and Remark 2.3.2, if there exists a generalized cryptographic semi-regular sequence $\bm{F}=(f_{1},\ldots,f_{m})$ with $\dim_{K}(R/\langle f_{1},\ldots,f_{m}\rangle)_{D}=1$ for $D=\tilde{d}_{\rm reg}(\langle f_{1},\ldots,f_{m}\rangle_{R})$, it follows that $Z\cap\Lambda^{\rm gcsr}$ contains a non-empty Zariski open set in $Z$ equipped with the induced topology, which implies that [29, Conjecture 4.3.4] is true for the parameter $(n,m,d_{1},\ldots,d_{m})$.
-3 Proofs of main results
-Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$ (in any characteristic), and let $I$ be a proper homogeneous ideal of $R$.
-Let $F=\{f_{1},\ldots,f_{m}\}$ be a subset of $R$ generating $I$, and put $d_{j}:=\deg(f_{j})$ (total degree) for $1\leq j\leq m$.
-For a linear transformation $\sigma$ of variables $x_{1},\ldots,x_{n}$ over $K$ represented by a $n\times n$ matrix $P$ over $K$, we denote by $h^{\sigma}:=h(\bm{x}\cdot P)$ the image of $h\in R$ by $\sigma$, where $\bm{x}:=(x_{1},\ldots,x_{n})$.
-3.1 Proofs of Theorem 1 (1) and (2)
-Let $\prec$ be the graded reverse lexicographic ordering on $R$ with $x_{n}\prec\cdots\prec x_{1}$.
-We denote by $N$ the number of its projective zeros over $\overline{K}$ of the homogeneous ideal $I\subset R$.
-Throughout this subsection, assume that $\mathrm{Krull.dim}(R/I)\leq 1$, i.e., $N<\infty$.
-Lemma 3.1.1.
-If the cardinality of $K$ is greater than $N$, there exist a non-negative integer $d$ and a linear form $\ell\in R$ with $\ell\notin I$ such that the $K$-linear map $(R/I)_{d-1}\to(R/I)_{d}$ defined by the multiplication by $\ell$ is surjective.
-Proof.
-The assertion is proved in the proof of (i) $\Rightarrow$ (iv) in [11, Theorem 3.3.4] when $K$ is an infinite field: In fact, it suffices for the proof that one can take a linear form $\ell\in R$ that does not vanish at any projective zero over $\overline{K}$ of $I$.
-Even in the case where $K$ is an finite field $\mathbb{F}_{q}$, if $q$ is greater than the number $N$ of projective zeros over $\overline{K}$ of $I$, we prove that such a linear form always exists.
-For each projective zero $\bm{a}=(a_{1}:\cdots:a_{n})\in\mathbb{P}^{n-1}(\overline{K})$ of $I$, we denote by $S_{\bm{a}}$ the set of all linear forms that vanish at $\bm{a}$, say
-$S_{\bm{a}}=\{y_{1}x_{1}+\cdots+y_{n}x_{n}\mid y_{i}\in K\mbox{ and }y_{1}a_{1}+\cdots y_{n}a_{n}=0\},$
-which is isomorphic, as $K$-linear spaces, to the following $K$-linear space of $K^{n}$:
-$S_{\bm{a}}^{\prime}=\{(y_{1},\ldots,y_{n})\in K^{n}\mid a_{1}y_{1}+\cdots a_{n}y_{n}=0\}.$
-The dimension of $S_{\bm{a}}^{\prime}$ as a $K$-linear space is at most $n-1$.
-Indeed, for an arbitrary non-zero entry $a_{i}$ of $\bm{a}$, the vector with $1$ in the $i$-th entry and 00’s elsewhere does not belong to $S_{\bm{a}}^{\prime}$, so that $S_{\bm{a}}^{\prime}$ is a proper subspace of $K^{n}$.
-Here, we have
-#$\#\left(\bigcup_{\bm{a}\in V_{\overline{K}}(I)}S_{\bm{a}}\right)\leq\sum_{{\bm{a}\in V_{\overline{K}}(I)}}\#S_{\bm{a}}\leq Nq^{n-1}<q^{n},$
-whence a desired linear form over $K$ exists. ∎
-Proof of Theorem 1 (1):
-Note that, by a choice of $\ell$, the Krull dimension of $R/\langle I,\ell\rangle$ is zero.
-For $d^{\prime}\geq 1$, the $-\ell$ map $(R/I)_{d^{\prime}-1}\to(R/I)_{d^{\prime}}$ is surjective if and only if $(R/I)_{d^{\prime}}=0$, which follows from the following exact sequence:
-${(R/I)_{d^{\prime}-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\times\ell}$${(R/I)_{d^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(R/\langle I,\ell\rangle)_{d^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
-When $d^{\prime}=d$, we have $(R/\langle I,\ell\rangle)_{d}=0$, by the surjectivity of the multiplication map $(R/I)_{d^{\prime}-1}\to(R/I)_{d^{\prime}}$ by $\ell$.
-Thus, it follows that $(R/\langle I,\ell\rangle)_{d^{\prime}}=0$ for any $d^{\prime}\geq d$, from which the $-\ell$ map $(R/I)_{d^{\prime}-1}\to(R/I)_{d^{\prime}}$ is surjective for any $d^{\prime}$ with $d^{\prime}\geq d$.
-Note that the minimal $d$ for which the map is surjective is equal to $d_{\rm reg}(\langle I,\ell\rangle_{R})$.
-We first note that $\widetilde{d}_{\rm reg}(I)=\widetilde{d}_{\rm reg}(I^{\tau})$ and ${d}_{\rm reg}(\langle I,\ell\rangle_{R^{\prime}})={d}_{\rm reg}(\langle I^{\tau},\ell^{\tau}\rangle_{R^{\prime}})$ for any invertible linear transformation $\tau$ of variables.
-There exists a $k\in\{1,\ldots,n\}$ such that $x_{k}$-coefficient of $\ell$ is not zero.
-We may also assume that the $x_{k}$-coefficient is $1$, without loss of generality.
-Let $\sigma_{1}$ be the linear transformation of variables $x_{1},\ldots,x_{n}$ defined by a permutation sending $x_{k}$ to $x_{n}$, and write $\ell^{\sigma_{1}}=a_{0}x_{0}+a_{1}x_{1}+\cdots+a_{n}x_{n}$ for $a_{i}\in K$ with $a_{n}=1$.
-Then, let $\sigma_{2}$ be a linear transformation of variables $x_{1},\ldots,x_{n}$ represented by the matrix
-$P:=\begin{pmatrix}1&&&-a_{1}\\ &\ddots&&\vdots\\ &&1&-a_{n-1}\\ &&&1\\ \end{pmatrix},$
-where blank entries are 00’s. Namely, for $h\in R$, we define
-$h^{\sigma_{2}}:=h(\bm{x}\cdot P)=h(x_{1},\ldots,x_{n-1},x_{n}-a_{1}x_{1}-\cdots-a_{n-1}x_{n-1})$
-with $\bm{x}:=(x_{1},\ldots,x_{n})$.
-Then, the invertible transformation $\sigma:=\sigma_{2}\circ\sigma_{1}$ of variables $x_{1},\ldots,x_{n}$ sends $\ell$ to $x_{n}$, and we have the equalities $D^{\prime}:=\widetilde{d}_{\rm reg}(I)=\widetilde{d}_{\rm reg}(I^{\sigma})$ and $D:=d_{\rm reg}(\langle I,\ell\rangle_{R})=d_{\rm reg}(\langle I^{\sigma},x_{n}\rangle_{R})$.
-The assertion follows from Theorem 2.1.9.∎
-In the following, we shall prove Theorem 1 (2), for which we first prove the following easy lemma in the case where $m=n-1$:
-Lemma 3.1.2.
-With notation as above, if $m=n-1$ and if the sequence $(f_{1},\ldots,f_{m},\ell)$ is cryptographic semi-regular (equivalently, it is regular since $m+1=n$), then we have $\widetilde{d}_{\rm reg}(I)=d_{\rm reg}(\langle I,\ell\rangle_{R})-1$, and hence
-$\mathrm{max.GB.deg}_{\prec}(I^{\sigma})\leq\mathrm{max}\{d_{\rm reg}(\langle I,\ell\rangle_{R}),\widetilde{d}_{\rm reg}(I)\}=d_{\rm reg}(\langle I,\ell\rangle_{R})=\sum_{j=1}^{n-1}(d_{j}-1)+1.$
+With notation as above, if $m=n-1$ and if the sequence $(f_{1},\ldots,f_{m},x_{n})$ is cryptographic semi-regular (equivalently, it is regular since $m+1=n$), then we have $\widetilde{d}_{\rm reg}(I)=d_{\rm reg}(I+\langle x_{n}\rangle_{R})-1$, and hence
+$\mathrm{max.GB.deg}_{\prec}(I)\leq\mathrm{max}\{d_{\rm reg}(I+\langle x_{n}\rangle),\widetilde{d}_{\rm reg}(I)\}=d_{\rm reg}(I+\langle x_{n}\rangle)=\sum_{j=1}^{n-1}(d_{j}-1)+1$
+for the graded reverse lexicographical ordering $\prec$ with $x_{n}\prec\cdots\prec x_{1}$.
 Moreover, $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular (equivalently, it is regular since $m<n$).
 Proof.
-Recall that the the regularity of a sequence of homogeneous polynomials is equivalent to the semi-regularity if the number of polynomials is less than or equal to the number of variables.
-For the sequence $(f_{1},\ldots,f_{m},\ell)$, the number of polynomials and that of variables are both $n$, whence it is regular.
-Therefore, $(f_{1},\ldots,f_{m})$ is also regular, and
-${\rm HS}_{R/I}(z)={\rm HS}_{R/\langle I,\ell\rangle}(z)\cdot\frac{1}{1-z}={\rm HS}_{R/\langle I,\ell\rangle}(z)\cdot(1+z+z^{2}+\cdots)$
+Since $(f_{1},\ldots,f_{m},x_{n})$ is regular, $(f_{1},\ldots,f_{m})$ is also regular, so that
+${\rm HS}_{R/I}(z)={\rm HS}_{R/(I+\langle x_{n}\rangle)}(z)\cdot\frac{1}{1-z}={\rm HS}_{R/(I+\langle x_{n}\rangle)}(z)\cdot(1+z+z^{2}+\cdots)$
 with
-${\rm HS}_{R/\langle I,\ell\rangle}(z)=\frac{\prod_{j=1}^{n-1}(1-z^{d_{j}})\cdot(1-z)}{(1-z)^{n}}=\frac{\prod_{j=1}^{n-1}(1-z^{d_{j}})}{(1-z)^{n-1}}=\prod_{j=1}^{n-1}(1+z+\cdots+z^{d_{j}-1}),$
-which is a polynomial.
-Here, we have $\mathrm{HF}_{R/I}(d)=\sum_{i=0}^{d}\mathrm{HF}_{R/\langle I,\ell\rangle}(i)$ for any $d$.
-Since $\mathrm{HF}_{R/\langle I,\ell\rangle}(d)=0$ (resp.
-$\mathrm{HF}_{R/\langle I,\ell\rangle}(d)\neq 0$) for $d\geq d_{\rm reg}(\langle I,\ell\rangle_{R})$ (resp.
-$d\leq d_{\rm reg}(\langle I,\ell\rangle_{R})-1$), we have $\widetilde{d}_{\rm reg}(I)=d_{\rm reg}(\langle I,\ell\rangle_{R})-1$, as desired.
-∎
-Proof of Theorem 1 (2):
-Suppose that $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular.
-Then, applying Lemma 2.2.8, we have the inequality (2.2.3), the right hand side of which is equal to $D_{\rm new}$ for $m\geq n$.
-It also follows from Lemma 3.1.2 that $\mathrm{max.GB.deg}_{\prec}(I)$ is upper-bounded by $\sum_{j=1}^{n-1}(d_{j}-1)+1$, which is equal to $D_{\rm new}$ for $m=n-1$, as desired.∎
-3.2 Proof of Proposition 1
-As in the previous subsection, let $\prec$ be the graded reverse lexicographic ordering on $R=K[x_{1},\ldots,x_{n}]$ with $x_{n}\prec\cdots\prec x_{1}$.
-We also denote the restriction of $\prec$ on $K[x_{1},\ldots,x_{n-1}]$ by the same notation $\prec$.
-In the following, we shall prove Proposition 1, namely the inequality (1.0.1) becomes an equality if $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic.
-Here, a weakly reverse lexicographic ideal is a monomial ideal $J$ such that if $x^{\alpha}$ is one of the minimal generators of $J$ then every monomial of the same degree which preceeds $x^{\alpha}$ must belong to $J$ as well (see also e.g., [34, Section 4] for the definition).
-As in [23] and [24] (see also [28] and [29]), we consider the elimination ideal
-$I^{\prime}:=\langle I,x_{n}\rangle_{R}\cap K[x_{1},\ldots,x_{n-1}]=\langle(f_{1})|_{x_{n}=0},\ldots,(f_{m})|_{x_{n}=0}\rangle_{K[x_{1},\ldots,x_{n-1}]}.$
-Since $R/\langle I,x_{n}\rangle$ and $K[x_{1},\ldots,x_{n-1}]/I^{\prime}$ are isomorphic as graded rings, they have the same Hilbert series, and $d_{\rm reg}(I^{\prime})=d_{\rm reg}(\langle I,x_{n}\rangle_{R})$.
-Put $D:=d_{\rm reg}(I^{\prime})=d_{\rm reg}(\langle I,x_{n}\rangle_{R})$.
-By comparing Hilbert series, it can be also easily checked that, the sequence $(f_{1},\ldots,f_{m},x_{n})\in R^{m+1}$ is cryptographic semi-regular if and only if $(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})\in K[x_{1},\ldots,x_{n-1}]^{m}$ is cryptographic semi-regular, equivalently these sequences are both $D$-regular.
-Let $G$ (resp.
-$G^{\prime}$) be the reduced Gröbner basis of $I$ (resp.
-$I^{\prime}$) with respect to $\prec$.
-Note that, since $\langle I,x_{n}\rangle_{R}=\langle I^{\prime},x_{n}\rangle_{R}$, the reduced Gröbner basis of $\langle I,x_{n}\rangle_{R}$ is given by $G^{\prime}\cup\{x_{n}\}$, whence $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=\mathrm{max.GB.deg}_{\prec}(\langle I,x_{n}\rangle_{R})$.
-To prove Proposition 1, we start with extending some results in [28] to our case.
-Lemma 3.2.1 (cf. [28, Corollary 2]).
-With notation as above, we assume that $(f_{1},\ldots,f_{m},x_{n})\in R^{m+1}$ is cryptographic semi-regular (i.e., $(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})\in K[x_{1},\ldots,x_{n-1}]^{m}$ is cryptographic semi-regular), namely both the sequences are $D$-regular with $D:=d_{\rm reg}(I^{\prime})=d_{\rm reg}(\langle I,x_{n}\rangle_{R})$.
-Then, we have
-$(\langle\mathrm{LM}_{\prec}(I)\rangle_{R})_{d}=(\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{R})_{d}$ (3.2.1)
-for any $d$ with $d<D$.
-Proof.
-We first prove $(\langle\mathrm{LM}_{\prec}(I)\rangle_{R})_{d}\supset(\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{R})_{d}$ for any non-negative integer $d$.
-For this, it suffices to show that $\mathrm{LM}_{\prec}(f)\in\mathrm{LM}_{\prec}(I)$ for any $f\in I_{d}^{\prime}\smallsetminus\{0\}$.
-We may assume that $f$ is homogeneous, and then it is written as $f=g+x_{n}h$ for some $g\in I_{d}$ and $h\in R_{d-1}$.
-Here, $g$ is not divisible by $x_{n}$, since otherwise $f\notin K[x_{1},\ldots,x_{n-1}]$, a contradiction.
-Therefore, we have $\mathrm{LM}_{\prec}(f)=\mathrm{LM}_{\prec}(g)$ with $g\in I$, as desired.
-Since $(f_{1},\ldots,f_{m},x_{n})$ is $D$-regular, it follows from [14, Proposition 2 a)] (cf. [28, Corollary 1]), its sub-sequence $(f_{1},\ldots,f_{m})$ is also $D$-regular. Hence, we have
-${\rm HS}_{R/I}(z)$ $\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})\cdot(1-z)}{(1-z)^{n}}\cdot\frac{1}{1-z}$
-$\equiv{\rm HS}_{R/\langle I,x_{n}\rangle}(z)\cdot\frac{1}{1-z}\equiv{\rm HS}_{R/\langle I,x_{n}\rangle}(z)\cdot(1+z+z^{2}+\cdots)$
-$\equiv{\rm HS}_{K[x_{1},\ldots,x_{n-1}]/I^{\prime}}(z)\cdot(1+z+z^{2}+\cdots)\pmod{z^{D}}.$
-Thus, for any $d$ with $d<D=d_{\rm reg}(I^{\prime})$, one has ${\rm HF}_{R/I}(d)=\sum_{i=0}^{d}\mathrm{HF}_{K[x_{1},\ldots,x_{n-1}]/I^{\prime}}(i)$, so that
-$\dim_{K}R_{d}-\dim_{K}I_{d}=\sum_{i=0}^{d}(\dim_{K}K[x_{1},\ldots,x_{n-1}]_{i}-\dim_{K}I^{\prime}_{i}).$ (3.2.2)
-Since $(JR)_{d}=x_{n}^{d}J_{0}\oplus x_{n}^{d-1}J_{1}\oplus\cdots\oplus x_{n}J_{d-1}\oplus J_{d}$ for any of $J=I^{\prime}$ and $J=K[x_{1},\ldots,x_{n-1}]$ (note that this holds also for $d\geq D$), the right hand side of (3.2.2) is equal to $\dim_{K}R_{d}-\dim_{K}(I^{\prime}R)_{d}$, so that $\dim_{K}(I^{\prime}R)_{d}=\dim_{K}I_{d}$.
-Here, it follows from [22, Theorem 5.2.6] that $\dim_{K}(\langle\mathrm{LM}_{\prec}(I)\rangle_{R})_{d}=\dim_{K}(\langle\mathrm{LM}_{\prec}(I^{\prime}R)\rangle_{R})_{d}$.
-Since $\langle\mathrm{LM}_{\prec}(I^{\prime}R)\rangle_{R}=\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{R}$, we finally obtain (3.2.1).
-∎
-Lemma 3.2.2 (cf. [28, Lemma 2]).
-Under the same setting as in Lemma 3.2.1, for each degree $d<D$, we have
-${\rm LM}_{\prec}(G)_{d}={\rm LM}_{\prec}(G^{\prime})_{d}.$ (3.2.3)
-Proof.
-The equality (3.2.3) can be proved by the induction on $d$, as follows. The case $d=0$ is trivial. Suppose that (3.2.3) holds for any $d$ with $10\leq d<D-$ 1.
-Let $t$ be any monomial in ${\rm LM}_{\prec}(G)_{d+1}$, so that $t=\mathrm{LM}_{\prec}(g)$ for some $g\in G$. By $d+1<D$, it follows from Lemma 3.2.1 that
-$(\langle\mathrm{LM}_{\prec}(I)\rangle_{R})_{d+1}=(\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{R})_{d+1},$
-whence ${\rm LM}_{\prec}(g)$ is divided by ${\rm LM}_{\prec}(g^{\prime})$ for some $g^{\prime}\in G^{\prime}$.
-Since $G$ is reduced, the induction hypothesis implies that $\mathrm{deg}(g^{\prime})=d+1$, and thus $\mathrm{LM}_{\prec}(g)=\mathrm{LM}_{\prec}(g^{\prime})$.
-Therefore, we obtain ${\rm LM}_{\prec}(G)_{d+1}\subset{\rm LM}_{\prec}(G^{\prime})_{d+1}$.
-The same argument shows that ${\rm LM}_{\prec}(G)_{d+1}\supset{\rm LM}_{\prec}(G^{\prime})_{d+1}$, as desired.
-∎
-Lemma 3.2.3.
-Under the same setting as in Lemma 3.2.1, we also suppose that $R/\langle I,x_{n}\rangle$ is Artinian (equivalently $K[x_{1},\ldots,x_{n-1}]/I^{\prime}$ is Artinian), namely $D<\infty$.
-Furthermore, we assume that $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic.
-Then we have $\mathrm{max.GB.deg}_{\prec}(I)\geq D$ and $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=D$.
-Proof.
-Put $d:=\mathrm{max.GB.deg}_{\prec}(I)$.
-Assume for a contradiction that $d<D$.
-Since $\mathrm{Krull.dim}(R/\langle I,x_{n}\rangle)=0$, it follows from Lemma 2.1.5 that there is an element $g$ in $G$ with ${\rm LM}_{\prec}(g)=x_{n-1}^{d^{\prime}}$ for some $d^{\prime}\leq d$.
-Then, any monomial $t\in R_{d^{\prime}}$ with $t\succeq x_{n-1}^{d^{\prime}}$ (namely $t$ is any monomial in $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}$) belongs to $\langle{\rm LM}_{\prec}(I)\rangle_{R}=\langle{\rm LM}_{\prec}(G)\rangle_{R}$ by the weakly reverse lexicographicness.
-Here, by $d<D$ together with (3.2.3), we have $\langle{\rm LM}_{\prec}(G)\rangle_{R}=\langle{\rm LM}_{\prec}(G^{\prime})_{\leq d}\rangle_{R}$.
-This implies that $t$ also belongs to $\langle{\rm LM}_{\prec}(G^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$, and hence $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}/I^{\prime}_{d^{\prime}}=0$.
-Thus, we have $d\geq d^{\prime}\geq d_{\rm reg}(I^{\prime})$, a contradiction.
-By the same argument as above, we can show that $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=d_{\rm reg}(I^{\prime})$ as follows.
-Note that $\mathrm{max.GB.deg}_{\prec}(I^{\prime})\leq d_{\rm reg}(I^{\prime})$ follows from a general fact written in the second paragraph of Remark 3.2.4 below.
-Thus, we assume to the contrary that $d:=\mathrm{max.GB.deg}_{\prec}(I^{\prime})<D:=d_{\rm reg}(I^{\prime})$.
-Then, it follows from (3.2.3) that ${\rm LM}_{\prec}(G^{\prime})={\rm LM}_{\prec}(G)_{\leq d}$ and the ideal $\langle{\rm LM}_{\prec}(G^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$ has the weak reverse lexicographicness up to $d$.
-Since there is an element $g$ in $G^{\prime}$ with ${\rm LM}_{\prec}(g)=x_{n-1}^{d^{\prime}}$ for some $d^{\prime}\leq d$, any monomial in $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}$ belongs to $\langle{\rm LM}_{\prec}(G^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$ and so $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}/I^{\prime}_{d^{\prime}}=0$, which implies $d\geq d^{\prime}\geq D$, a contradiction.
-∎
-Remark 3.2.4.
-In Lemma 3.2.3, when $R/\langle I,x_{n}\rangle$ is Artinian (or equivalently $K[x_{1},\ldots,x_{n-1}]/I^{\prime}$ is Artinian), we can easily prove $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=D$ if $\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$ is a weakly reverse lexicographic ideal, assuming neither that $(f_{1},\ldots,f_{m},x_{n})$ is cryptographic semi-regular nor that $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic.
-In general, for any homogeneous ideal $J\subset R$ with $\mathrm{Krull.dim}(R/J)=0$, we have $\mathrm{max.GB.deg}_{\prec}(J)\leq\mathrm{reg}(J)=d_{\rm reg}(J)$ by [4, Corollary 2.5] (see also [23, Proposition 4.15]), and it is straightforward that the equality holds if $\langle\mathrm{LM}_{\prec}(J)\rangle_{R}$ is a weakly reverse lexicographic.
-Proof of Proposition 1:
-Put $D:=d_{\rm reg}(\langle I,x_{n}\rangle_{R})=d_{\rm reg}(I^{\prime})$ and $D^{\prime}:=\widetilde{d}_{\rm reg}(I)$.
-From our assumption that $R/\langle I,x_{n}\rangle$ is Artinian, we have the inequality (1.0.1), i.e., $\mathrm{max.GB.deg}(I)\leq\mathrm{max}\{D,D^{\prime}\}$.
-The case where $\mathrm{Krull.dim}(R/I)=0$ is clear from Remark 3.2.4, so we assume that $\mathrm{Krull.dim}(R/I)=1$.
-Note that, for the reduced Gröbner basis $G$ of $I$, the elements of $\mathrm{LM}_{\prec}(G)$ are the minimal generators of the monomial ideal $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}=\langle\mathrm{LM}_{\prec}(G)\rangle_{R}$.
-First we consider the case where $D^{\prime}\geq D$, so that $\max\{D,D^{\prime}\}=D^{\prime}$.
-Put $d:=\mathrm{max.GB.deg}_{\prec}(I)$.
-Then, it follows from Macaulay’s basis theorem (cf.
-[26, Theorem 1.5.7]) that $R_{d}/I_{d}$ has a $K$-linear basis of the form
-$\{t\in R_{d}:\mbox{$t$ is a monomial and }{t\notin\langle{\rm LM}_{\prec}(I)\rangle_{R}=\langle\mathrm{LM}_{\prec}(G)\rangle_{R}}\},$
-which is called the standard monomial basis.
-We set $r:=\dim_{K}R_{d}/I_{d}$, and write this standard monomial basis as $\{t_{1},\ldots,t_{r}\}$.
-In the following, we prove that $\{t_{1}x_{n}^{s},\ldots,t_{r}x_{n}^{s}\}$ is a $K$-linear basis of $R_{d+s}/I_{d+s}$ for any $s$ with $s\geq 1$, from which we have $\dim_{K}R_{d}/I_{d}=\dim_{K}R_{d+s}/I_{d+s}$ for any positive integer $s$, so that $d\geq D^{\prime}$ and therefore $d=D^{\prime}$.
-By Lemma 3.2.3, we have $d\geq D$, and thus it can be shown similarly to Proof of Theorem 1 (1) that the multiplication-by-$x_{n}^{s}$ map from $R_{d}/I_{d}$ to $R_{d+s}/I_{d+s}$ is surjective (since $R/\langle I,x_{n}\rangle$ is Artinian).
-Therefore $B_{s}:=\{t_{1}x_{n}^{s},\ldots,t_{r}x_{n}^{s}\}$ generates $R_{d+s}/I_{d+s}$.
-Suppose to contrary that $B_{s}$ is not a basis.
-In this case, there exists an $i$ such that $t_{i}x_{n}^{s}$ is divisible by $\mathrm{LM}_{\prec}(g)$ for some $g\in G$.
-Indeed, since the elements of $B_{s}$ are linearly dependent over $K$, we obtain $f:=a_{1}t_{1}x_{n}^{s}+\cdots+a_{r}t_{r}x_{n}^{s}\in I_{d+s}$ for some $(a_{1},\ldots,a_{r})\in K^{r}\smallsetminus\{(0,\ldots,0)\}$.
-Thus, for the least $i$ with $a_{i}\neq 0$, the leading monomial of $f$ is $t_{i}x_{n}^{s}$, which is divisible by $\mathrm{LM}_{\prec}(g)$ for some $g\in G$.
-Putting $u_{i}=\mathrm{GCD}(t_{i},\mathrm{LM}_{\prec}(g))$ and $s_{i}=t_{i}/u_{i}$, we have $u_{i}s_{i}x_{n}^{s}=t_{i}x_{n}^{s}$.
-Since $t_{i}=s_{i}u_{i}$ is not divisible by $\mathrm{LM}_{\prec}(g)$, we can write $u_{i}x_{n}^{s^{\prime}}=\mathrm{LM}_{\prec}(g)$ for some $s^{\prime}\leq s$.
-(We note that $\deg s_{i}=d-\deg u_{i}\geq\deg{\rm LM}_{\prec}(g)-\deg u_{i}=s^{\prime}$.) Note that $s^{\prime}\geq 1$ since otherwise $t_{i}$ is divisible by $\mathrm{LM}_{\prec}(g)$.
-Take an arbitrary monomial $s^{\prime}_{i}$ such that $\deg s^{\prime}_{i}=s^{\prime}$ and $s_{i}^{\prime}$ divides $s_{i}$.
-Then, by the weakly reverse lexicographicness, as ${\rm LM}_{\prec}(g)=u_{i}x_{n}^{s^{\prime}}\preceq u_{i}s^{\prime}_{i}$, the monomial $s^{\prime}_{i}u_{i}$ should belong to $\langle{\rm LM}_{\prec}(G)\rangle_{R}$.
-Moreover, since $s^{\prime}_{i}u_{i}$ divides $t_{i}=s_{i}u_{i}$, the monomial $t_{i}=s_{i}u_{i}$ also belongs to $\langle{\rm LM}_{\prec}(G)\rangle_{R}$, which is a contradiction.
-Finally, consider the remaining case, i.e., $D>D^{\prime}$, so that $\max\{D,D^{\prime}\}=D$. In this case, it follows from Lemma 3.2.3 that $d\geq D$, whence $d=D$.∎
-3.3 Proof of Theorem 1 (3): Complexity estimation
-We shall prove (3) of Theorem 1. For this, we estimate the complexity of computing a Gröbner basis of a homogeneous ideal with Macaulay matrices.
-Let $F=\{f_{1},\ldots,f_{m}\}\subset R$ be a set of homogeneous polynomials in $R\smallsetminus K$, put $d_{j}:=\deg(f_{j})$ for $1\leq j\leq m$, and let $\prec$ be an arbitrary graded monomial ordering on $R$.
-The complexity of the Gröbner basis computation for $F$ is estimated as that of computing the reduced row echelon form (RREF) of the degree-$D$ Macaulay matrix, denoted by $M_{\leq D}(F)$, where $D$ is an arbitrary upper-bound on $\mathrm{max.GB.deg}_{\prec}(F)$.
+${\rm HS}_{R/(I+\langle x_{n}\rangle)}(z)=\frac{\prod_{j=1}^{n-1}(1-z^{d_{j}})\cdot(1-z)}{(1-z)^{n}}=\frac{\prod_{j=1}^{n-1}(1-z^{d_{j}})}{(1-z)^{n-1}}=\prod_{j=1}^{n-1}\left(\sum_{k=0}^{d_{j}-1}z^{k}\right),$
+which is a polynomial. It is clear that $\widetilde{d}_{\rm reg}(I)=d_{\rm reg}(I+\langle x_{n}\rangle_{R})-1$, as desired. ∎
+Proof of Theorem 1.1 (1): If $m\geq n$, our assumption that $(f_{1},\ldots,f_{m},x_{n})$ is cryptographic semi-regular implies
+${\color[rgb]{0,0,0}d_{\rm reg}(I+\langle x_{n}\rangle)=\deg\left(\left[\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n-1}}\right]\right)+1,}$
+whose right hand side is upper-bounded by $D^{(n,m)}$ for $m\geq n$.
+On the other hand, since $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular, it follows from Lemma 2.2.11 that we have the inequality (2.2.3), the right hand side of which is nothing but $D^{(n,m)}$ for $m\geq n$.
+Therefore, we obtain (1.3).
+If $m=n-1$, it follows from Lemma 3.1.1 that $\mathrm{max}\{d_{\rm reg}(I+\langle x_{n}\rangle),\widetilde{d}_{\rm reg}(I)\}$ is upper-bounded by $\sum_{j=1}^{n-1}(d_{j}-1)+1$, which is equal to $D^{(n,m)}$ for $m=n-1$.∎
+Next, we shall prove (2) of Theorem 1.1. For this, we estimate the complexity of computing a Gröbner basis of a homogeneous ideal with Macaulay matrices.
+Let $\prec$ be an arbitrary graded monomial ordering on $R$.
+It is well-known that the complexity of the Gröbner basis computation for the homogeneous set $F$ is estimated as that of computing the reduced row echelon form (RREF) of the degree-$D$ Macaulay matrix, denoted by $M_{\leq D}(F)$, where $D$ is an arbitrary upper-bound on $\mathrm{max.GB.deg}_{\prec}(F)$.
 Here, the degree-$D$ Macaulay matrix $M_{\leq D}(F)$ (in homogeneous case) is a block matrix of the form
 $M_{\leq D}(F)=\begin{pmatrix}M_{D}(F)&&\\ &\ddots&\\ &&M_{d_{0}}(F)\end{pmatrix}$
-with $d_{0}:=\min\{\deg(f_{j}):1\leq j\leq m\}$, where each $M_{d}(F)$ is a homogeneous Macaulay matrix defined as follows: For each $d$ with $d\geq d_{0}$, the degree-$d$ homogeneous Macaulay matrix $M_{d}(F)$ of $F$ has columns indexed by the terms of $R_{d}$ sorted, from left to right, according to the chosen order $\prec$.
-The rows of $M_{d}(F)$ are indexed by the polynomials $m_{i,j}f_{j}$, where $m_{i,j}\in R$ is a term such that $\deg(m_{i,j}f_{j})=d$.
+with $d_{0}:=\min\{\deg(f_{j}):1\leq j\leq m\}$, where each $M_{d}(F)$ is a homogeneous Macaulay matrix defined as follows: For each $d$ with $d\geq d_{0}$, the degree-$d$ homogeneous Macaulay matrix $M_{d}(F)$ of $F$ has columns indexed by the monomials of $R_{d}$ sorted, from left to right, according to the chosen order $\prec$.
+The rows of $M_{d}(F)$ are indexed by the polynomials $m_{i,j}f_{j}$, where $m_{i,j}\in R$ is a monomial such that $\deg(m_{i,j}f_{j})=d$.
+For a monomial $t\in\mathcal{T}_{d}$ of degree $d$, the $(m_{i,j}f_{j},t)$-entry of $M_{d}(F)$ is defined as the coefficient of $t$ in $m_{i,j}f_{j}$.
 Theorem 3.1.2.
@@
   1. (1)
-    [2, Proposition 1] The RREF of $M_{\leq D}(F)$ is computed in
+    [3, Proposition 1] The RREF of $M_{\leq D}(F)$ is computed in
     $O\left(mD\binom{n+D-1}{D}^{\omega}\right).$ (3.1.1)
-    Hence, if $\prec$ is a graded order, then the complexity of the Gröbner basis computation for $F$ is upper-bounded by (3.3.2) for $D=\mathrm{max.GB.deg}_{\prec}(F)$.
+    Hence, if $\prec$ is a graded ordering, then the complexity of the Gröbner basis computation for $F$ is upper-bounded by (3.1.1) for $D=\mathrm{max.GB.deg}_{\prec}(F)$.
   2. (2)
@@
     $O\left(m\binom{n+D-1}{D}^{\omega}\right).$ (3.1.2)
-    Hence, if $\prec$ is a graded order, and if $\mathrm{max.GB.deg}_{\prec}(F)=D=O(n)$, then the complexity of the Gröbner basis computation for $F$ is upper-bounded by (3.3.2).
+    Hence, if $\prec$ is a graded ordering, and if $\mathrm{max.GB.deg}_{\prec}(F)=D=O(n)$, then the complexity of the Gröbner basis computation for $F$ is upper-bounded by (3.1.2).
 Proof.
@@
 It follows from $D=O(n)$ that $c$ is upper-bounded by a constant.
-By $10<\varepsilon<$ 1, it is clear that $\frac{c+\varepsilon}{c+1}$ is also upper-bounded by a positive constant smaller than $1$, whence
-$\frac{1}{1-\left(\frac{c+\varepsilon}{c+1}\right)^{\omega}}=O(1)$
-as $n\to\infty$. We have proved the assertion (2). ∎
+By $10<\varepsilon<$ 1, it is clear that $\frac{c+\varepsilon}{c+1}$ is also upper-bounded by a positive constant smaller than $1$, whence $1/{(1-(\frac{c+\varepsilon}{c+1})^{\omega})}=O(1)$ as $n\to\infty$.
+We have proved the assertion (2).
+∎
 Remark 3.1.3.
@@
 Since the complexity of (2) is $(2\ell)^{\omega}$, the total complexity to obtain $B$ is $q\cdot(2\ell)^{\omega}\leq(k/\ell)\cdot 2^{\omega}\ell^{\omega}=O(k\ell^{\omega-1})$, as desired.
-Proof of Theorem 1 (3):
-For the graded reverse lexicographic ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$, it follows from Theorem 1 (2) that $D_{\rm new}$ gives an upper bound on $\mathrm{max.GB.deg}_{\prec}(F)$.
+Proof of Theorem 1.1 (2): For the graded reverse lexicographical ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$, it follows from Theorem 1.1 (1) that $D^{(n,m)}$ gives an upper bound on $\mathrm{max.GB.deg}_{\prec}(F)$.
 Therefore, a Gröbner basis of $\langle F\rangle_{R}$ is obtained by computing the RREF of $M_{\leq D}(F)$ for $D=D^{(n,m)}$.
-Here, the value $D^{(n,m)}$ of (1.0.4) is maximized at $m=n$: In this case, $D^{(n,m)}$ is equal to the Lazard’s bound $\sum_{j=1}^{n}(d_{j}-1)+1$, which is $O(n)$ for fixed $d_{1},\ldots,d_{m}$.
-Therefore, it follows from Theorem 3.3.1 (2) that the RREF of $M_{\leq D}(F)$ for $D=D^{(n,m)}$ is computed in the complexity (1.0.5), as desired.∎
+Here, the value $D^{(n,m)}$ of (1.4) is maximized at $m=n$: In this case, $D^{(n,m)}$ is equal to the Lazard’s bound $\sum_{j=1}^{n}(d_{j}-1)+1$, which is $O(n)$ for fixed $d_{1},\ldots,d_{m}$.
+Therefore, it follows from Theorem 3.1.2 (2) that the RREF of $M_{\leq D}(F)$ for $D=D^{(n,m)}$ is computed in the complexity (1.5), as desired.∎
+Example 3.1.4.
+The bound $D^{(n,m)}$ in (1.3) is optimal for a homogeneous ideal $I$ satisfying the assumptions of Theorem 1.1.
+For example, considering the case where $K=\mathbb{F}_{31}$ and $(n,m)=(3,4)$, and letting $F=\{f_{1},f_{2},f_{3},f_{4}\}$ be a generating set for $I$ with
+$f_{1}$ $=$ $3x_{1}^{2}+20x_{1}x_{2}+x_{2}^{2}+7x_{1}x_{3}+15x_{2}x_{3},$
+$f_{2}$ $=$ $18x_{1}^{2}+15x_{1}x_{2}+14x_{2}^{2}+24x_{1}x_{3}+4x_{2}x_{3},$
+$f_{3}$ $=$ $6x_{1}^{2}+16x_{1}x_{2}+x_{2}^{2}+x_{1}x_{3}+8x_{2}x_{3},$
+$f_{4}$ $=$ $21x_{1}^{2}+x_{1}x_{2}+21x_{2}^{2}+24x_{1}x_{3}+6x_{2}x_{3},$
+we have $\mathrm{Krull.dim}(R/I)=1$ and
+$\mathrm{HS}_{R/(I+\langle x_{3}\rangle)}(z)=1+2z,\quad\mathrm{HS}_{R/I}(z)=1+3z+2z^{2}+z^{3}+z^{4}+z^{5}+\cdots$
+with $d_{\rm reg}(I+\langle x_{3}\rangle)=2$ and $\widetilde{d}_{\rm reg}(I)=3$, and therefore $(f_{1},f_{2},f_{3},f_{4},x_{3})$ (resp.
+$(f_{1},f_{2},f_{3},f_{4})$) is cryptographic semi-regular (resp.
+generalized cryptographic semi-regular).
+One can easily verify $D^{(3,4)}=3$.
+The reduced Gröbner basis of $I$ with respect to the graded reverse lexicographical ordering $\prec$ with $x_{3}\prec x_{2}\prec x_{1}$ is
+$\{x_{2}x_{3}^{2},x_{1}^{2}+27x_{2}x_{3},x_{1}x_{2}+14x_{2}x_{3},x_{2}^{2}+30x_{2}x_{3},x_{1}x_{3}+26x_{2}x_{3}\},$
+whence $\mathrm{max.GB.deg}_{\prec}(I)=3$. Namely, we have
+$3=\mathrm{max.GB.deg}_{\prec}(I)=\max\{d_{\rm reg}(I+\langle x_{n}\rangle),\widetilde{d}_{\rm reg}(I)\}=\widetilde{d}_{\rm reg}(I)=D^{(n,m)}.$
+On the other hand, $\mathrm{max}\{d_{\rm reg}(I+\langle x_{n}\rangle),\widetilde{d}_{\rm reg}(I)\}$ in (1.3) can be strictly less than $D^{(n,m)}$, e.g., when $F:=\{f_{1}:=23x_{1}^{2}+26x_{1}x_{2}+13x_{2}^{2},f_{2}:=29x_{1}^{2}+x_{1}x_{2}+x_{2}^{2}\}$ with $n=m=2$ and $K=\mathbb{F}_{31}$, a straightforward computation shows that $I=\langle F\rangle$ satisfies the assumptions of Theorem 1.1 with $\mathrm{max.GB.deg}_{\prec}(I)=2$ and $d_{\rm reg}(I+\langle x_{2}\rangle)=\widetilde{d}_{\rm reg}(I)=2<D^{(2,2)}=3$.
+(In this case, we can replace $D^{(n,m)}$ by $D^{(n,m)}-1$ in (1.5)).
+3.2 Proof of Theorem 1.2
+As in the previous subsection, let $\prec$ be the graded reverse lexicographical ordering on $R=K[x_{1},\ldots,x_{n}]$ with $x_{n}\prec\cdots\prec x_{1}$.
+We also denote the restriction of $\prec$ on $K[x_{1},\ldots,x_{n-1}]$ by the same notation $\prec$.
+In the following, we shall prove Theorem 1.2, namely the inequality (1.2) becomes an equality if the initial ideal ${\rm in}_{\prec}(I)=\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic.
+Here, a weakly reverse lexicographic ideal is a monomial ideal $J$ such that if $x^{\alpha}$ is one of the minimal generators of $J$ then every monomial of the same degree which preceeds $x^{\alpha}$ must belong to $J$ as well (see also e.g., [43, Section 4] for the definition).
+As in [30] and [31] (see also [35] and [36]), we consider the elimination ideal
+${\color[rgb]{0,0,0}I^{\prime}:=(I+\langle x_{n}\rangle_{R})\cap K[x_{1},\ldots,x_{n-1}]=\langle f_{1}|_{x_{n}=0},\ldots,f_{m}|_{x_{n}=0}\rangle_{K[x_{1},\ldots,x_{n-1}]}.}$
+Since $R/(I+\langle x_{n}\rangle)$ and $K[x_{1},\ldots,x_{n-1}]/I^{\prime}$ are isomorphic as graded rings, they have the same Hilbert series, and $d_{\rm reg}(I^{\prime})=d_{\rm reg}(I+\langle x_{n}\rangle_{R})$.
+Put $D:=d_{\rm reg}(I^{\prime})=d_{\rm reg}(I+\langle x_{n}\rangle_{R})$.
+By comparing Hilbert series, it can be also easily checked that, the sequence $(f_{1},\ldots,f_{m},x_{n})\in R^{m+1}$ is cryptographic semi-regular if and only if $(f_{1}|_{x_{n}=0},\ldots,f_{m}|_{x_{n}=0})\in K[x_{1},\ldots,x_{n-1}]^{m}$ is cryptographic semi-regular, equivalently these sequences are both $D$-regular.
+Let $G$ (resp.
+$G^{\prime}$) be the reduced Gröbner basis of $I$ (resp.
+$I^{\prime}$) with respect to $\prec$.
+Note that, since $I+\langle x_{n}\rangle_{R}=I^{\prime}R+\langle x_{n}\rangle_{R}$, the reduced Gröbner basis of $I+\langle x_{n}\rangle_{R}$ is given by $G^{\prime}\cup\{x_{n}\}$, whence $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=\mathrm{max.GB.deg}_{\prec}(I+\langle x_{n}\rangle_{R})$.
+To prove Theorem 1.2, we start with extending some results in [35] to our case.
+Lemma 3.2.1 (cf. [35, Corollary 2]).
+With notation as above, we assume that $(f_{1},\ldots,f_{m},x_{n})\in R^{m+1}$ is cryptographic semi-regular (i.e., $(f_{1}|_{x_{n}=0},\ldots,f_{m}|_{x_{n}=0})\in K[x_{1},\ldots,x_{n-1}]^{m}$ is cryptographic semi-regular), namely both the sequences are $D$-regular with $D:=d_{\rm reg}(I^{\prime})=d_{\rm reg}(I+\langle x_{n}\rangle_{R})$.
+Then, we have
+$(\langle\mathrm{LM}_{\prec}(I)\rangle_{R})_{d}=(\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{R})_{d}$ (3.2.1)
+for any $d$ with $d<D$.
+Proof.
+We first prove $(\langle\mathrm{LM}_{\prec}(I)\rangle_{R})_{d}\supset(\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{R})_{d}$ for any non-negative integer $d$.
+For this, it suffices to show that $\mathrm{LM}_{\prec}(f)\in\mathrm{LM}_{\prec}(I)$ for any $f\in I_{d}^{\prime}\smallsetminus\{0\}$.
+We may assume that $f$ is homogeneous, and then it is written as $f=g+x_{n}h$ for some $g\in I_{d}$ and $h\in R_{d-1}$.
+Here, $g$ is not divisible by $x_{n}$, since otherwise $f\notin K[x_{1},\ldots,x_{n-1}]$, a contradiction.
+Therefore, we have $\mathrm{LM}_{\prec}(f)=\mathrm{LM}_{\prec}(g)$ with $g\in I$, as desired.
+Since $(f_{1},\ldots,f_{m},x_{n})$ is $D$-regular, it follows from [15, Proposition 2 a)] (cf. [35, Corollary 1]), its sub-sequence $(f_{1},\ldots,f_{m})$ is also $D$-regular. Hence, we have
+${\rm HS}_{R/I}(z)$ $\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})\cdot(1-z)}{(1-z)^{n}}\cdot\frac{1}{1-z}$
+$\equiv{\rm HS}_{R/(I+\langle x_{n}\rangle)}(z)\cdot\frac{1}{1-z}\equiv{\rm HS}_{R/(I+\langle x_{n}\rangle)}(z)\cdot(1+z+z^{2}+\cdots)$
+$\equiv{\rm HS}_{K[x_{1},\ldots,x_{n-1}]/I^{\prime}}(z)\cdot(1+z+z^{2}+\cdots)\pmod{z^{D}}.$
+Thus, for any $d$ with $d<D=d_{\rm reg}(I^{\prime})$, one has ${\rm HF}_{R/I}(d)=\sum_{i=0}^{d}\mathrm{HF}_{K[x_{1},\ldots,x_{n-1}]/I^{\prime}}(i)$, so that
+$\dim_{K}R_{d}-\dim_{K}I_{d}=\sum_{i=0}^{d}(\dim_{K}K[x_{1},\ldots,x_{n-1}]_{i}-\dim_{K}I^{\prime}_{i}).$ (3.2.2)
+Since $(JR)_{d}=x_{n}^{d}J_{0}\oplus x_{n}^{d-1}J_{1}\oplus\cdots\oplus x_{n}J_{d-1}\oplus J_{d}$ for any of $J=I^{\prime}$ and $J=K[x_{1},\ldots,x_{n-1}]$ (note that this holds also for $d\geq D$), the right hand side of (3.2.2) is equal to $\dim_{K}R_{d}-\dim_{K}(I^{\prime}R)_{d}$, so that $\dim_{K}(I^{\prime}R)_{d}=\dim_{K}I_{d}$.
+Here, it follows from [29, Theorem 5.2.6] that $\dim_{K}(\langle\mathrm{LM}_{\prec}(I)\rangle_{R})_{d}=\dim_{K}(\langle\mathrm{LM}_{\prec}(I^{\prime}R)\rangle_{R})_{d}$.
+Since $\langle\mathrm{LM}_{\prec}(I^{\prime}R)\rangle_{R}=\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{R}$, we finally obtain (3.2.1).
+∎
+Lemma 3.2.2 (cf. [35, Lemma 2]).
+Under the same setting as in Lemma 3.2.1, for each degree $d<D$, we have
+${\rm LM}_{\prec}(G)_{d}={\rm LM}_{\prec}(G^{\prime})_{d}.$ (3.2.3)
+Proof.
+This can be proved just by translating the proof of [35, Lemma 2]: Replace $R^{\prime}$, $R$, $\langle\bm{F}^{h}\rangle_{R^{\prime}}$, $\langle\bm{F}^{\rm top}\rangle_{R}$, $G_{\rm hom}$, and $G_{\rm top}$ in the proof by $R$, $K[x_{1},\ldots,x_{n-1}]$, $I$, $I^{\prime}$, $G$, and $G^{\prime}$ in our case, respectively.
+∎
+Lemma 3.2.3.
+Under the same setting as in Lemma 3.2.1, we also suppose that $R/(I+\langle x_{n}\rangle)$ is Artinian (equivalently $K[x_{1},\ldots,x_{n-1}]/I^{\prime}$ is Artinian), namely $D<\infty$.
+Furthermore, we assume that $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic.
+Then we have $\mathrm{max.GB.deg}_{\prec}(I)\geq D$ and $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=D$.
+Proof.
+Put $d:=\mathrm{max.GB.deg}_{\prec}(I)$.
+Assume for a contradiction that $d<D$.
+Since $\mathrm{Krull.dim}(R/(I+\langle x_{n}\rangle))=0$, it follows from Lemma 2.1.5 that there is an element $g$ in $G$ with ${\rm LM}_{\prec}(g)=x_{n-1}^{d^{\prime}}$ for some $d^{\prime}\leq d$.
+Then, any monomial $t\in R_{d^{\prime}}$ with $t\succeq x_{n-1}^{d^{\prime}}$ (namely $t$ is any monomial in $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}$) belongs to the initial ideal ${\rm in}_{\prec}(I)=\langle{\rm LM}_{\prec}(I)\rangle_{R}=\langle{\rm LM}_{\prec}(G)\rangle_{R}$ by the weakly reverse lexicographicness.
+Here, by $d<D$ together with (3.2.3), we have $\langle{\rm LM}_{\prec}(G)\rangle_{R}=\langle{\rm LM}_{\prec}(G^{\prime})_{\leq d}\rangle_{R}$.
+This implies that $t$ also belongs to $\langle{\rm LM}_{\prec}(G^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$, and hence $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}/I^{\prime}_{d^{\prime}}=0$.
+Thus, we have $d\geq d^{\prime}\geq d_{\rm reg}(I^{\prime})$, a contradiction.
+By the same argument as above, we can show that $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=d_{\rm reg}(I^{\prime})$ as follows.
+Note that $\mathrm{max.GB.deg}_{\prec}(I^{\prime})\leq d_{\rm reg}(I^{\prime})$ follows from a general fact written in the second paragraph of Remark 3.2.4 below.
+Thus, we suppose to the contrary that $d:=\mathrm{max.GB.deg}_{\prec}(I^{\prime})<D:=d_{\rm reg}(I^{\prime})$.
+Then, it follows from (3.2.3) that ${\rm LM}_{\prec}(G^{\prime})={\rm LM}_{\prec}(G)_{\leq d}$, and the initial ideal ${\rm in}_{\prec}(\langle G^{\prime}\rangle)=\langle{\rm LM}_{\prec}(G^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$ has the weak reverse lexicographicness up to $d$.
+Since there is an element $g$ in $G^{\prime}$ with ${\rm LM}_{\prec}(g)=x_{n-1}^{d^{\prime}}$ for some $d^{\prime}\leq d$, any monomial in $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}$ belongs to $\langle{\rm LM}_{\prec}(G^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$ and so $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}/I^{\prime}_{d^{\prime}}=0$, which implies $d\geq d^{\prime}\geq D$, a contradiction.
+∎
+Remark 3.2.4.
+In Lemma 3.2.3, when $R/(I+\langle x_{n}\rangle)$ is Artinian (or equivalently $K[x_{1},\ldots,x_{n-1}]/I^{\prime}$ is Artinian), we can easily prove $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=D$ if ${\rm in}_{\prec}(I^{\prime})=\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$ is a weakly reverse lexicographic ideal, assuming neither that $(f_{1},\ldots,f_{m},x_{n})$ is cryptographic semi-regular nor that ${\rm in}_{\prec}(I)=\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic.
+In general, for any homogeneous ideal $J\subset R$ with $\mathrm{Krull.dim}(R/J)=0$, we have $\mathrm{max.GB.deg}_{\prec}(J)\leq\mathrm{reg}(J)=d_{\rm reg}(J)$ by [5, Corollary 2.5] (see also [30, Proposition 4.15]), and it is straightforward that the equality holds if the initial ideal $\langle\mathrm{LM}_{\prec}(J)\rangle_{R}$ of $J$ is a weakly reverse lexicographic.
+Proof of Theorem 1.2: Put $D:=d_{\rm reg}(I+\langle x_{n}\rangle_{R})=d_{\rm reg}(I^{\prime})$ and $D^{\prime}:=\widetilde{d}_{\rm reg}(I)$.
+From our assumption that $R/(I+\langle x_{n}\rangle)$ is Artinian, we have the inequality (1.2), i.e., $\mathrm{max.GB.deg}(I)\leq\mathrm{max}\{D,D^{\prime}\}$.
+The case where $\mathrm{Krull.dim}(R/I)=0$ is clear from Remark 3.2.4, so we assume that $\mathrm{Krull.dim}(R/I)=1$.
+Note that, for the reduced Gröbner basis $G$ of $I$, the elements of $\mathrm{LM}_{\prec}(G)$ are the minimal generators of the monomial ideal $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}=\langle\mathrm{LM}_{\prec}(G)\rangle_{R}$.
+First we consider the case where $D^{\prime}\geq D$, so that $\max\{D,D^{\prime}\}=D^{\prime}$.
+Put $d:=\mathrm{max.GB.deg}_{\prec}(I){\leq D^{\prime}}$.
+Then, it follows from Macaulay’s basis theorem (cf.
+[33, Theorem 1.5.7]) that $R_{d}/I_{d}$ has a $K$-linear basis of the form
+$\{t\in R_{d}:\mbox{$t$ is a monomial and }{t\notin\langle{\rm LM}_{\prec}(I)\rangle_{R}=\langle\mathrm{LM}_{\prec}(G)\rangle_{R}}\},$
+which is called the standard monomial basis.
+We set $r:=\dim_{K}R_{d}/I_{d}$, and write this standard monomial basis as $\{t_{1},\ldots,t_{r}\}$ with $t_{1}\succ\cdots\succ t_{r}$.
+In the following, we prove that $\{t_{1}x_{n}^{s},\ldots,t_{r}x_{n}^{s}\}$ is a $K$-linear basis of $R_{d+s}/I_{d+s}$ for any $s$ with $s\geq 1$, from which we have $\dim_{K}R_{d}/I_{d}=\dim_{K}R_{d+s}/I_{d+s}$ for any positive integer $s$, so that $d\geq D^{\prime}$ and therefore $d=D^{\prime}$.
+By Lemma 3.2.3, we have $d\geq D$, and thus it follows from Remark 2.1.12 that the multiplication-by-$x_{n}^{s}$ map from $R_{d}/I_{d}$ to $R_{d+s}/I_{d+s}$ is surjective (since $R/(I+\langle x_{n}\rangle)$ is Artinian).
+Therefore the set $B_{s}:=\{t_{1}x_{n}^{s},\ldots,t_{r}x_{n}^{s}\}$ generates $R_{d+s}/I_{d+s}$ over $K$.
+Suppose to the contrary that $B_{s}$ is not a basis.
+In this case, there exists an $i$ such that $t_{i}x_{n}^{s}$ is divisible by $\mathrm{LM}_{\prec}(g)$ for some $g\in G$.
+Indeed, since the elements of $B_{s}$ are linearly dependent over $K$, we obtain $f:=a_{1}t_{1}x_{n}^{s}+\cdots+a_{r}t_{r}x_{n}^{s}\in I_{d+s}$ for some $(a_{1},\ldots,a_{r})\in K^{r}\smallsetminus\{(0,\ldots,0)\}$.
+Thus, for the least $i$ with $a_{i}\neq 0$, the leading monomial of $f$ is $t_{i}x_{n}^{s}$, which is divisible by $\mathrm{LM}_{\prec}(g)$ for some $g\in G$.
+Putting $u_{i}=\mathrm{GCD}(t_{i},\mathrm{LM}_{\prec}(g))$ and $s_{i}=t_{i}/u_{i}$, we have $u_{i}s_{i}x_{n}^{s}=t_{i}x_{n}^{s}$.
+Since $t_{i}=s_{i}u_{i}$ is not divisible by $\mathrm{LM}_{\prec}(g)$ and since $u_{i}=\mathrm{GCD}(t_{i},\mathrm{LM}_{\prec}(g))$, we can write $u_{i}x_{n}^{s^{\prime}}=\mathrm{LM}_{\prec}(g)$ for some $s^{\prime}\leq s$.
+(We note that $\deg s_{i}=d-\deg u_{i}\geq\deg{\rm LM}_{\prec}(g)-\deg u_{i}=s^{\prime}$.) Note that $s^{\prime}\geq 1$ since otherwise $t_{i}$ is divisible by $\mathrm{LM}_{\prec}(g)$.
+Take an arbitrary monomial $s^{\prime}_{i}$ such that $\deg s^{\prime}_{i}=s^{\prime}$ and $s_{i}^{\prime}$ divides $s_{i}$.
+Then, by ${\rm LM}_{\prec}(g)=u_{i}x_{n}^{s^{\prime}}\preceq u_{i}s^{\prime}_{i}$ together with the weakly reverse lexicographicness of the initial ideal ${\rm in}_{\prec}(I)=\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$, the monomial $s^{\prime}_{i}u_{i}$ should belong to $\langle{\rm LM}_{\prec}(G)\rangle_{R}$.
+Moreover, since $s^{\prime}_{i}u_{i}$ divides $t_{i}=s_{i}u_{i}$, the monomial $t_{i}=s_{i}u_{i}$ also belongs to $\langle{\rm LM}_{\prec}(G)\rangle_{R}$, which is a contradiction.
+Finally, consider the remaining case, i.e., $D>D^{\prime}$, so that $\max\{D,D^{\prime}\}=D$. In this case, it follows from Lemma 3.2.3 that $d\geq D$, whence $d=D$.∎
+4 Results on genericness with conjectures
+Let $n$ be a positive integer, and let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$.
+We may assume that $K$ is infinite.
+Let $m$, $d_{1},\ldots,d_{m}$ be positive integers.
+As in Section 2.3, let $V_{n,m,(d_{1},\ldots,d_{m})}=R_{d_{1}}\times\cdots\times R_{d_{m}}$, and naturally endow it with the Zariski topology.
+Recall Definition 2.3.1 for the definition of the genericness of a property for $\bm{F}=(f_{1},\ldots,f_{m})\in V_{n,m,(d_{1},\ldots,d_{m})}$.
+For $\bm{F}=(f_{1},\ldots,f_{m})\in V_{n,m,(d_{1},\ldots,d_{m})}$, we denote by $\langle\bm{F}\rangle_{R}$ (or by $\langle\bm{F}\rangle$ simply) the ideal generated by $f_{1},\ldots,f_{m}$.
+4.1 Restricted genericness and our conjectures
+For $n$, $m$, $(d_{1},\ldots,d_{m})$, and a non-negative integer $k$, we set
+$\Lambda_{n,m,(d_{1},\ldots,d_{m})}^{{\rm dim}=k}$ :=$:=\{\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}:\mathrm{Krull.dim}(R/\langle\bm{F}\rangle)=k\},$
+$\Lambda_{n,m,(d_{1},\ldots,d_{m})}^{{\rm dim}\geq k}$ :=$:=\{\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}:\mathrm{Krull.dim}(R/\langle\bm{F}\rangle)\geq k\},$
+[…]
+[…]
+It follows from Lemma 2.2.12 that $\Lambda_{n,m,(d_{1},\ldots,d_{m})}^{\rm gcsr}\cap\Lambda_{n,m,(d_{1},\ldots,d_{m})}^{{\rm dim}=0}\subset\Lambda_{n,m,(d_{1},\ldots,d_{m})}^{\rm csr}$ and the equality holds if and only if $m\geq n$.
+Here, as a variant of Conjecture 2.3.7, we start with raising the following conjecture (cf. [36, Conjecture 4.3.4] as its strict version):
+Conjecture 4.1.1.
+Assume that $K$ is an infinite field.
+For any $n$, $m$, and $(d_{1},\ldots,d_{m})$, the property that $\bm{F}=(f_{1},\ldots,f_{m})\in V_{n,m,(d_{1},\ldots,d_{m})}$ is generalized cryptographic semi-regular is generic.
+We see that Conjectures 2.3.7 and 4.1.1 are equivalent to each other. For this, we first prove the following lemma:
+Lemma 4.1.2.
+For any $n$, $m$, and $(d_{1},\ldots,d_{m})$ with $m\geq n$, the property that $R/\langle f_{1},\ldots,f_{m}\rangle$ is Artinian is generic.
+Proof.
+The case $m=n$ is well-known to hold: More strongly, $\Lambda^{{\rm dim}=0}_{n,n,(d_{1},\ldots,d_{n})}$ is a non-empty Zariski-open set of $V_{n,n,(d_{1},\ldots,d_{n})}$.
+Assume $m>n$, fix $n$, $m$, and $(d_{1},\ldots,d_{m})$.
+Then, the preimage of $\Lambda^{{\rm dim}=0}_{n,n,(d_{1},\ldots,d_{n})}$ by a natural projection sending an element of $V_{n,m,(d_{1},\ldots,d_{m})}$ to the first $n$-entries is included in $\Lambda^{{\rm dim}=0}_{n,m,(d_{1},\ldots,d_{m})}$.
+Since $\pi$ is continuous, the preimage is Zariski-open in $V_{n,m,(d_{1},\ldots,d_{m})}$, and it is clearly non-empty, whence the assertion holds.
+∎
+Proposition 4.1.3.
+Assume that $K$ is an infinite field. For $m\geq n$, Conjecture 4.1.1 is equivalent to Fröberg’s conjecture (Conjecture 2.3.7).
+Proof.
+Assume that there exists a non-empty Zariski-open set $U_{1}\subset V_{n,m,(d_{1},\ldots,d_{m})}$ with $U_{1}\subset\Lambda^{\rm gcsr}_{n,m,(d_{1},\ldots,d_{m})}$.
+By Lemma 4.1.2, there exists a non-empty Zariski-open set $U_{2}\subset V_{n,m,(d_{1},\ldots,d_{m})}$ such that $U_{2}\subset\Lambda^{{\rm dim}=0}_{n,m,(d_{1},\ldots,d_{m})}$.
+Since both $U_{1}$ and $U_{2}$ are dense, we have $\emptyset\neq U_{1}\cap U_{2}\subset\Lambda^{\rm gcsr}_{n,m,(d_{1},\ldots,d_{m})}\cap\Lambda^{{\rm dim}=0}_{n,m,(d_{1},\ldots,d_{m})}\subset\Lambda^{\rm csr}_{n,m,(d_{1},\ldots,d_{m})}$, whence Fröberg’s conjecture holds for $(n,m,d_{1},\ldots,d_{m})$.
+The converse is obvious by $\Lambda^{\rm csr}_{n,m,(d_{1},\ldots,d_{m})}\subset\Lambda^{\rm gcsr}_{n,m,(d_{1},\ldots,d_{m})}$ (cf. Lemma 2.2.12). ∎
+Despite of the equivalence between two conjectures, those cannot guarantee that $\Lambda_{n,m,(d_{1},\ldots,d_{m})}^{\rm gcsr}\cap\Lambda_{n,m,(d_{1},\ldots,d_{m})}^{{\rm dim}\geq 1}$ includes a non-empty open set in the subspace $\Lambda_{n,m,(d_{1},\ldots,d_{m})}^{{\rm dim}\geq 1}$ equipped with the induced topology.
+To discuss this, we define the restricted genericness as follows:
+Definition 4.1.4 (Restricted genericness).
+Let $V^{\prime}$ be a non-empty subset of $V_{n,m,(d_{1},\ldots,d_{m})}$.
+For given $n$, $m$, and $(d_{1},\ldots,d_{m})$, a property $\mathcal{P}$ for $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ is said to be generic on $V^{\prime}$ if it holds on a non-empty open set of $V^{\prime}$ with respect to the induced topology, namely there exists a non-empty Zariski open set $U$ in $V_{n,m,(d_{1},\ldots,d_{m})}$ such that $\emptyset\neq U\cap V^{\prime}\subset\{\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}:\mbox{$\bm{F}$ satisfies $\mathcal{P}$}\}$.
+Namely, it is important whether the property that $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ is generalized cryptographic semi-regular is generic on $\Lambda_{n,m,(d_{1},\ldots,d_{m})}^{{\rm dim}\geq 1}$ or not.
+We will also discuss the genericness on
+$Z_{\bm{p}}:=\{\bm{H}\in V_{n,m,(d_{1},\ldots,d_{m})}:\bm{p}\in V_{\overline{K}}(\bm{H})\}$ (4.1.1)
+for each point $\bm{p}\in\mathbb{P}^{n-1}(K)$.
+Remark 4.1.5.
+If $K$ is an infinite field and if the property that $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ is generalized cryptographic semi-regular is generic on $\Lambda_{n,m,(d_{1},\ldots,d_{m})}^{{\rm dim}\geq 1}$, then ‘almost’ all sequences $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ such that $\mathrm{Krull.dim}(R/\langle\bm{F}\rangle)\geq 1$ are expected to be generalized cryptographic semi-regular.
+As the first discussion about the restricted genericness of the generalized cryptographic semi-regularity on $\Lambda_{n,m,(d_{1},\ldots,d_{m})}^{{\rm dim}\geq 1}$, we consider the following sub-case, where $\mathrm{Krull.dim}(R/\langle\bm{F}\rangle)=1$ and $\dim_{K}R/\langle\bm{F}\rangle_{D}=1$ for $D=\widetilde{d}_{\rm reg}(\langle\bm{F}\rangle)<\infty$.
+In this case, $\langle\bm{F}\rangle$ has a unique projective zero $\bm{p}$ in $\mathbb{P}^{n-1}(K)$.
+First we show that it suffices to consider the case of the point $\bm{o}:=(0:\cdots:0:1)$ as $\bm{p}$, by the following lemma:
+Lemma 4.1.6.
+If the property that $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ is generalized cryptographic semi-regular is generic on $Z_{\bm{o}}$, then it is also generic on $Z_{\bm{p}}$ for any $\bm{p}\in\mathbb{P}^{n-1}(K)$.
+Proof.
+For each $\bm{p}=(p_{1}:\ldots:p_{n})\in\mathbb{P}^{n-1}(K)$, let $\phi_{\bm{p}}$ be an invertible projective linear transformation sending $\bm{p}$ to $\bm{o}$.
+We may assume $p_{n}\neq 0$ and so $p_{n}=1$ without loss of generality, so that we can take $\phi_{\bm{p}}$ as
+${(x_{1}:\cdots:x_{n})\to(x_{1}-p_{1}x_{n}:\cdots:x_{n-1}-p_{1}x_{n-1}:x_{n})}.$
+Then, the induced polynomial map $\phi_{\bm{p}}^{\ast}:R\to R$ is an isomorphism substituting $x_{i}-p_{i}x_{n}$ to $x_{i}$ in $f\in R$ for $11\leq i\leq n-$ 1 and stabilizing $x_{n}$.
+Here, we consider a map on $V_{n,m,(d_{1},\ldots,d_{m})}$ defined by
+$\Phi_{\bm{p}}:V_{n,m,(d_{1},\ldots,d_{m})}\to V_{n,m,(d_{1},\ldots,d_{m})}\ ;\ (f_{1},\ldots,f_{m})\mapsto(\phi_{\bm{p}}^{\ast}(f_{1}),\ldots,\phi_{\bm{p}}^{\ast}(f_{m})).$
+It is straightforward that this is an invertible linear (and hence homeomorphic) map on $V_{n,m,(d_{1},\ldots,d_{m})}$.
+Moreover, it is easily checked that $\phi_{\bm{p}}^{*}$ does not change LM for any $h\in R$ with respect to the graded reverse lexicographical ordering $\prec$ such that $x_{1}\succ\cdots\succ x_{n}$.
+Thus, for any subset $S$ of $R$, we have ${\rm LM}(S)={\rm LM}(\phi_{\bm{p}}^{*}(S))$, from which ${\rm in}_{\prec}(\langle\bm{F}\rangle)={\rm in}_{\prec}(\langle\Phi_{\bm{p}}(\bm{F})\rangle)$ for $\bm{F}$ in $V_{n,m,(d_{1},\ldots,d_{m})}$.
+This also implies that ${\rm HS}_{R/\langle\bm{F}\rangle}(z)={\rm HS}_{R/{\rm in}(\langle\bm{F}\rangle)}(z)={\rm HS}_{R/{\rm in}(\langle\Phi_{\bm{p}}(\bm{F})\rangle)}(z)={\rm HS}_{R/\langle\Phi_{\bm{p}}(\bm{F})\rangle}(z)$.
+From our assumption, we have a non-empty Zariski-open set $U_{\bm{o}}$ of $V_{n,m,(d_{1},\ldots,d_{m})}$ intersecting with $Z_{\bm{o}}$ such that $U_{\bm{o}}\cap Z_{\bm{o}}\subset\Lambda_{n,m,(d_{1},\ldots,d_{m})}^{\rm gcsr}\cap Z_{\bm{o}}$.
+Then $\Phi_{\bm{p}}(U_{\bm{o}})$ is a non-empty Zariski-open set of $V_{n,m,(d_{1},\ldots,d_{m})}$ intersecting with $\Phi_{\bm{p}}(Z_{\bm{o}})=Z_{\bm{p}}$.
+Moreover, as any $\bm{F}\in U_{\bm{o}}\cap Z_{\bm{o}}$ is generalized cryptographic semi-regular, $\Phi_{\bm{p}}(\bm{F})$ is also generalized cryptographic semi-regular, as ${\rm HS}_{R/\langle\bm{F}\rangle}(z)={\rm HS}_{R/\langle\Phi_{\bm{p}}(\bm{F})\rangle}(z)$.
+Thus, letting $U_{\bm{p}}=\Phi_{\bm{p}}(U_{\bm{o}})$, we have $U_{\bm{p}}\cap Z_{\bm{p}}\subset\Lambda_{n,m,(d_{1},\ldots,d_{m})}^{\rm gcsr}\cap Z_{\bm{p}}$.
+∎
+Note that
+$Z_{\bm{o}}=\{(f_{1},\ldots,f_{m})\in V_{n,m,(d_{1},\ldots,d_{m})}:\mbox{the $x_{n}^{d_{j}}$-coefficient of $f_{j}$ is zero ($1\leq\forall j\leq m$)}\}.$
+Definition 4.1.7 (Generic sequence for $Z_{\bm{o}}$).
+For checking the genericness of the generalized cryptographic semi-regularity of a sequence in $Z_{\bm{o}}$, we introduce a generic sequence $(g_{1},\ldots,g_{m})$ for $Z_{\bm{o}}$ with $\deg(g_{j})=d_{j}$ as follows: Let $\Omega$ be an extension field of $K$, and let $\mathcal{C}:=\{c_{j,t}:1\leq j\leq m,\ t\in\mathcal{T}_{d_{j}}\smallsetminus\{x_{n}^{d_{j}}\}\}$ be a set of elements in $\Omega$ that are algebraically independent over the prime field $K_{0}$ of $K$.
+Then we set $L:=K(\mathcal{C})$ and $L_{0}:=K_{0}(\mathcal{C})$.
+Putting $g_{j}=\sum_{t\in\mathcal{T}_{d_{j}}\smallsetminus\{x_{n}^{d_{j}}\}}c_{j,t}t$ for each $1\leq j\leq m$, we call a sequence $\bm{G}=(g_{1},\ldots,d_{m})$ in $K_{0}[\mathcal{C}][X]\subset L_{0}[X]\subset L[X]$ a generic sequence for $Z_{\bm{o}}$.
+We also set $\mathcal{F}_{j}:=\left\{\sum_{t\in\mathcal{T}_{d_{j}}\smallsetminus\{x_{n}^{d_{j}}\}}a_{j,t}t:a_{j,t}\in K\right\}$ for $1\leq j\leq m$, so that $Z_{\bm{o}}=\mathcal{F}_{d_{1}}\times\cdots\times\mathcal{F}_{d_{m}}$.
+Then we state an analogue of Lemma 2.3.2:
+Lemma 4.1.8.
+For any $\bm{F}=(f_{1},\ldots,f_{m})\in Z_{\bm{o}}\subset V_{n,m,(d_{1},\ldots,d_{m})}$, we have the coefficient-wise inequality
+$\mathrm{HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}(z)=\mathrm{HS}_{R_{L}/\langle f_{1},\ldots,f_{m}\rangle_{R_{L}}}\geq\mathrm{HS}_{R_{L}/\langle g_{1},\ldots,g_{m}\rangle_{R_{L}}}(z),$ (4.1.2)
+where $(g_{1},\ldots,g_{m})$ is the generic sequence for $Z_{\bm{o}}$ defined in Definition 4.1.7, and where $R_{L}=L[X]$ with $L=K(\mathcal{C})$.
+Proof.
+This can be proved similarly to Lemma 2.3.2, see [23, Lemma 1] for a proof by Fröberg. ∎
+Proposition 4.1.9.
+Assume that $K$ is an infinite field. Then, for given $n$, $m$, and $(d_{1},\ldots,d_{m})$, the following conditions (1), (2), and (3) are equivalent to each other:
+  1. (1)
+    There exists a generalized cryptographic semi-regular sequence $\bm{F}\in Z_{\bm{o}}$ with $D:=\widetilde{d}_{\rm reg}(\langle\bm{F}\rangle)<{\infty}$ such that $\dim_{K}(R/\langle\bm{F}\rangle)_{D}=1$.
+  2. (2)
+    The property that $\bm{F}\in Z_{\bm{o}}$ is generalized cryptographic semi-regular with $D=\widetilde{d}_{\rm reg}(\langle\bm{F}\rangle)<{\infty}$ satisfying $\dim_{K}(R/\langle\bm{F}\rangle)_{D}=1$ is generic on $Z_{\bm{o}}$ (in terms of Definition 4.1.4).
+  3. (3)
+    For $\mathcal{C}$, $L$, $R_{L}$, and $\bm{G}=(g_{1},\ldots,g_{m})$ as in Definition 4.1.7, we have
+    $\mathrm{HS}_{R_{L}/\langle g_{1},\ldots,g_{m}\rangle}(z)\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{D^{\prime}}}$ (4.1.3)
+    and $\dim_{L}(R_{L}/\langle\bm{G}\rangle)_{D^{\prime}}=1$ for $D^{\prime}=\widetilde{d}_{\rm reg}(\langle\bm{G}\rangle)<{\infty}$.
+Proof.
+The implication (2) $\Rightarrow$ (1) is obvious.
+To see (1) $\Rightarrow$ (3), by Lemmas 2.2.6 and 4.1.8, we have the same inequalities as in (2.3.3).
+Since $\bm{F}$ is generalized cryptographic semi-regular (i.e., $\widetilde{d}_{\rm reg}(\langle\bm{F}\rangle)$-regular), $\bm{G}$ is also $\widetilde{d}_{\rm reg}(\langle\bm{F}\rangle)$-regular.
+Moreover, since $\mathrm{Krull.dim}(R_{L}/\langle\bm{G}\rangle)\geq 1$, the inequality $1=\mathrm{HF}_{R/\langle\bm{F}\rangle}(d)\geq\mathrm{HF}_{R_{L}/\langle\bm{G}\rangle}(d)$ becomes an equality for any $d\geq D$, so that $\mathrm{HS}_{R/\langle\bm{F}\rangle}(z)=\mathrm{HS}_{R_{L}/\langle\bm{G}\rangle}(z)$ and $\widetilde{d}_{\rm reg}(\langle\bm{G}\rangle)=D<{\infty}$.
+The implication (3) $\Rightarrow$ (2) follows from Theorem 2.3.4. ∎
+From Proposition 4.1.9 and Lemma 4.1.6, if there exists a generalized cryptographic semi-regular sequence $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ with $\dim_{K}(R/\langle\bm{F}\rangle)_{D}=1$ for $D=\widetilde{d}_{\rm reg}(\langle\bm{F}\rangle)<{\infty}$, it follows that $Z_{\bm{p}}\cap\Lambda^{\rm gcsr}_{n,m,(d_{1},\ldots,d_{m})}$ includes a non-empty Zariski-open set in $Z_{\bm{p}}$ equipped with the induced topology for any $\bm{p}\in\mathbb{P}^{n-1}(K)$, which implies the following conjecture:
+Conjecture 4.1.10 (cf. [36, Conjecture 4.3.4]).
+Assume that $K$ is an infinite field.
+For any $n$, $m$, and $(d_{1},\ldots,d_{m})$ and for any $\bm{p}\in\mathbb{P}^{n-1}(K)$, the property that $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ is generalized cryptographic semi-regular is generic on $Z_{\bm{p}}$.
+Let us raise a more general conjecture:
+Conjecture 4.1.11.
+Assume that $K$ is an infinite field.
+For any $n$, $m$, and $(d_{1},\ldots,d_{m})$, the property that $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ is generalized cryptographic semi-regular is generic on $V_{n,m,(d_{1},\ldots,d_{m})}^{\dim\geq 1}$.
+4.2 Genericness on cryptographic semi-regularity and weakly reverse lexicographicness
+In this subsection, let $K$ be an infinite field.
+Here we give a proof of Theorem 1.5 and Theorem 1.6 using the notion of “parametric Gröbner basis” for the ideal generated by a “generic sequence” $\bm{G}=(g_{1},\ldots,g_{m})$ under the setting in Section 2.3 and Definition 4.1.7.
+We fix a monomial ordering $\prec$ on the set $\mathcal{T}$ of monomials in $X=\{x_{1},\ldots,x_{n}\}$ as the graded reverse lexicographical ordering with $x_{1}\succ x_{2}\succ\cdots\succ x_{n}$.
+As additional notations, we simply write ${\rm LM(*)}$ and ${\rm in(*)}$ for ${\rm LM}_{\prec}(*)$ and ${\rm in}_{\prec}(*)$, respectively, and we define the following:
+For each element $\bm{a}\in\mathcal{A}_{(d_{1},\ldots,d_{m})}:=\mathbb{A}^{N}(K)=\prod_{j=1}^{m}\mathbb{A}^{N_{j}}(K)$ with $N_{j}=\binom{n+d_{j}-1}{d_{j}}$, we define a map $\pi_{\bm{a}}$ from $K[\mathcal{C}][X]$ to $K[X]$ by substituting each $c_{j,t}$ with $\bm{a}_{j,t}$.
+The map $\pi_{\bm{a}}$ is generalized to a map from $K[\mathcal{C}][X]^{m}$ to $K[X]^{m}$, for which we use the same symbol.
+Thus, for a sequence $\bm{F}$ in $V_{n,m,(d_{1},\ldots,d_{m})}$, there exists $\bm{a}$ in $\mathcal{A}$ uniquely such that $\bm{F}=\pi_{\bm{a}}(\bm{G})$, where $\bm{G}$ is a generic sequence in terms of Fröberg, see the second paragraph of Section 2.3 for the definition.
+We also note that, for a polynomial $h$ in $L[X]$ with $L=K(\mathcal{C})$, if $\pi_{\bm{a}}$ does not annihilate any of the denominators of its coefficients, $\pi_{\bm{a}}(h)$ can be defined.
+Then, for each sequence $\bm{F}$ in $V_{n,m,(d_{1},\ldots,d_{m})}$, let ${\rm GB}(\bm{F})$ denote the reduced Gröbner basis of the ideal $\langle\bm{F}\rangle_{R}$ with respect to $\prec$.
+Moreover, for the generic sequence $\bm{G}$, let ${\rm GB}(\bm{G})$ denote the reduced Gröbner basis of the ideal $\langle\bm{G}\rangle_{L_{0}[X]}$ with respect to $\prec$.
+Then, it follows from $\bm{G}\subset L_{0}[X]\subset L[X]$ that ${\rm GB}(\bm{G})$ is also the reduced Gröbner basis of the ideal $\langle\bm{G}\rangle_{L[X]}$ with respect to $\prec$.
+Therefore, in the following, for deciding the shape of ${\rm GB}(\bm{G})$, we assume that all elements in $\mathcal{C}$ are algebraically independent over $K$.
+(See Remark 2.3.3.) Recall that ${\rm LM}({\rm GB}(\bm{F}))$ forms the minimal generating set of the initial ideal ${\rm in}(\langle\bm{F}\rangle_{R}):=\langle{\rm LM}(\langle\bm{F}\rangle_{R})\rangle_{R}$.
+We simply write $\langle\bm{G}\rangle$ and $\langle\bm{F}\rangle$ for $\langle\bm{G}\rangle_{L[X]}$ and $\langle\bm{F}\rangle_{K[X]}$, respectively.
+Theorem 4.2.1.
+If there exists a cryptographic semi-regular sequence $\bm{F}$ in
+$V_{n,m,(d_{1},\ldots,d_{m})}$, that is, $\bm{F}=\pi_{\bm{a}}(\bm{G})$ for some element $\bm{a}\in\mathcal{A}_{(d_{1},\ldots,d_{m})}$, such that ${\rm in}(\langle\bm{F}\rangle)$ is weakly reverse lexicographic, then we have ${\rm GB}(\bm{F})=\pi_{\bm{a}}({\rm GB}(\bm{G}))$ and ${\rm LM}({\rm GB}(\bm{F}))={\rm LM}({\rm GB}(\bm{G}))$.
+Thus, ${\rm in}(\langle\bm{G}\rangle)$ is also weakly reverse lexicographic.
+Then Theorem 2.3.4 can be applied directly to $\bm{G}$, from which it follows that there exists a non-empty Zariski-open set $U_{\mathcal{A}}$ of $\mathcal{A}_{(d_{1},\ldots,d_{m})}$ such that ${\rm LM}({\rm GB}(\bm{F}))={\rm LM}({\rm GB}(\pi_{\bm{b}}(\bm{G})))$ and ${\rm in}(\langle\pi_{\bm{b}}(\bm{G})\rangle)$ is weakly reverse lexicographic for any $\bm{b}\in U_{\mathcal{A}}$.
+We remark that, for each homogeneous ideal $I$, since the set of the leading monomials of the reduced Gröbner basis of $I$ forms the minimal generating set of the initial ideal ${\rm in}(I)$, the property of “weakly reverse lexicographicness” depends on the leading monomials of the reduced Gröbner basis.
+Considering the subset $U$ of $V_{n,m,(d_{1},\ldots,d_{m})}$ corresponding to $U_{\mathcal{A}}$ in Theorem 4.2.1, we have the following:
+Corollary 4.2.2.
+If there exists a cryptographic semi-regular sequence $\bm{F}$ in $V_{n,m,(d_{1},\ldots,d_{m})}$ such that ${\rm in}(\langle\bm{F}\rangle)$ is weakly reverse lexicographic, then there also exists a non-empty Zariski-open set $U$ of $V_{n,m,(d_{1},\ldots,d_{m})}$ such that $\bm{H}$ is cryptographic semi-regular and ${\rm in}(\langle\bm{H}\rangle)$ is weakly reverse lexicographic for any $\bm{H}$ in $U$.
+Namely, Fröberg’s conjecture and Moreno-Socias’ one are both true for $(n,m,d_{1},\ldots,d_{m})$.
+Proof of Theorem 4.2.1. It suffices to show the following by induction argument on a non-negative integer $d$:
+${\rm GB}(\bm{F})_{d}=\pi_{\bm{a}}({\rm GB}(\bm{G})_{d})\mbox{ with }{\rm LM}({\rm GB}(\bm{F})_{d})={\rm LM}({\rm GB}(\bm{G})_{d}).$ (4.2.1)
+For $d=0$, it is clear as ${\rm GB}(\bm{F})_{0}={\rm GB}(\bm{G})_{0}=\emptyset$.
+Thus, we show that the condition (4.2.1) holds for $d>0$ under the assumption on induction that it holds for $d-1$.
+Put ${\rm GB}(\bm{F})_{d}=\{h_{1},\ldots,h_{v}\}$ and ${\rm GB}(\bm{G})_{d}=\{\bar{h}_{1},\ldots,\bar{h}_{\bar{v}}\}$ with $v=\#({\rm GB}(\bm{F})_{d})$ and $\bar{v}=\#({\rm GB}(\bm{G})_{d})$, where we order $h_{1},\ldots,h_{v}$ and $\bar{h}_{1},\ldots,\bar{h}_{\bar{v}}$ so that ${\rm LM}(h_{1})\succ{\rm LM}(h_{2})\succ\cdots\succ{\rm LM}(h_{v})$ and ${\rm LM}(\bar{h}_{1})\succ{\rm LM}(\bar{h}_{2})\succ\cdots\succ{\rm LM}(\bar{h}_{\bar{v}})$, respectively.
+By the induction hypothesis, we have ${\rm LM}(\mathrm{GB}(\bm{F})_{<d})={\rm LM}(\mathrm{GB}(\bm{G})_{<d})$ and so ${\rm LM}(\langle\bm{F}\rangle_{<d})={\rm LM}(\langle\bm{G}\rangle_{<d})$.
+We let $Y:={\rm LM}(\langle\bm{F}\rangle_{<d})$.
+In the following we show $v=\bar{v}$ and $h_{i}=\pi_{\bm{a}}(\bar{h}_{i})$ with ${\rm LM}(h_{i})={\rm LM}(\bar{h}_{i})$ for each $i$.
+Claim 1: $\bar{v}=v$ holds.
+Let $W:=\{ts:s\in Y,\ t\in\mathcal{T}_{d-\deg(s)}\}$. By the theory of Gröbner basis, the following holds:
+${\rm LM}(\langle\bm{G}\rangle_{d})=W\cup{\rm LM}({\rm GB}(\bm{G})_{d}),\quad{\rm LM}(\langle\bm{F}\rangle_{d})=W\cup{\rm LM}({\rm GB}(\bm{F})_{d})$
+Moreover, ${\rm LM}(\langle\bm{G}\rangle_{d})$ (resp.
+${\rm LM}(\langle\bm{F}\rangle_{d})$) forms a linear basis of ${\rm in}(\langle\bm{G}\rangle)_{d}$ (resp.
+${\rm in}(\langle\bm{F}\rangle)_{d}$), respectively.
+Here we note ${\rm HS}_{K[X]/\langle\bm{F}\rangle}(z)={\rm HS}_{K[X]/{\rm in}(\langle\bm{F}\rangle)}(z)$ and ${\rm HS}_{L[X]/\langle\bm{G}\rangle}(z)={\rm HS}_{L[X]/{\rm in}(\langle\bm{G}\rangle)}(z)$.
+On the other hand, as $\bm{F}$ is cryptographic semi-regular, so is $\bm{G}$ by (1) $\Rightarrow$ (4) of Proposition 2.3.6, whence ${\rm HS}_{K[X]/\langle\bm{F}\rangle}(z)={\rm HS}_{L[X]/\langle\bm{G}\rangle}(z)$.
+Therefore, their $d$-th coefficients, namely $\dim_{K}K[X]_{d}/\langle\bm{F}\rangle_{d}$ and $\dim_{L}L[X]_{d}/\langle\bm{G}\rangle_{d}$, coincide with each other, from which we have $v=\#({\rm GB}(\bm{F})_{d})=\#({\rm GB}(\bm{G})_{d})=\bar{v}$.
+Claim 2: For each $i$, there is an element ¯$\bar{\ell}_{i}$ in $\langle\bm{G}\rangle\cap K[\mathcal{C}][X]$ with ${\rm supp}(\bar{\ell}_{i})\cap W=\emptyset$ such that $\pi_{\bm{a}}(\bar{\ell}_{i})$ is defined and is a non-zero constant multiple of $h_{i}$.
+Here we denote by ${\rm supp}(h_{i})$ the support of $h_{i}$, that is, the set of monomials appearing in $h_{i}$ with non-zero coefficient.
+For each $i$, as $h_{i}\in\langle\bm{F}\rangle$, there are polynomials $q_{i,1},\ldots,q_{i,m}\in K[X]$ such that
+$h_{i}=\sum_{j=1}^{m}q_{i,j}f_{j}=\sum_{j=1}^{m}q_{i,j}\pi_{\bm{a}}(g_{j})=\pi_{\bm{a}}\left(\sum_{j=1}^{m}q_{i,j}g_{j}\right).$
+Let $\ell_{i}:=\sum_{j=1}^{m}q_{i,j}g_{j}$ for each $i$.
+Then $\ell_{i}$ belongs to $K[\mathcal{C}][X]\cap\langle\bm{G}\rangle$ as $g_{i}\in K[\mathcal{C}][X]$.
+Moreover, as ${\rm LM}(\pi_{\bm{a}}(\ell_{i}))={\rm LM}(h_{i})$, we have ${\rm LM}(\ell_{i})\succeq{\rm LM}(h_{i})$.
+There are many choices on $q_{i,j}$, and $\ell_{i}$ may have some monomial in $W$ with non-zero coefficient.
+Thus, to eliminate such a monomial in $W$ from $\ell_{i}$, we apply the following procedure to ${\ell}_{i}$, which outputs a desired ~$\tilde{\ell}_{i}$:
+(Step 1) Check if $\ell_{i}$ has no monomial in $W$, that is, ${\rm supp}(\ell_{i})\cap W=\emptyset$ or not:
+(Step 2: Case where ${\rm supp}(\ell_{i})\cap W=\emptyset$)
+We set ~$\tilde{\ell}_{i}:=\ell_{i}$, and quite the procedure.
+(Step 3: Case where ${\rm supp}(\ell_{i})\cap W\not=\emptyset$)
+Let $t$ be the greatest monomial in ${\rm supp}(\ell_{i})\cap W$ with respect to $\prec$.
+Then we compute $\ell^{\prime}_{i}$ from $\ell_{i}$ by the following monomial reduction, by which the term with the monomial $t$ is eliminated:
+$\ell^{\prime}_{i}=\ell_{i}-b_{t}(\ell_{i})\frac{t}{{\rm LM}(\bar{p})}\bar{p},$
+where $b_{t}(\ell_{i})$ is the coefficient of $t$ in $\ell_{i}$ and $t$ is divided by ${\rm LM}(\bar{p})$ of some $\bar{p}\in{\rm GB}(\bm{G})_{<d}$.
+As $\pi_{\bm{a}}(\ell_{i})$ is defined, $\pi_{\bm{a}}(b_{t}(\ell_{i}))$ is also defined.
+Moreover, by the assumption of induction, $\pi_{\bm{a}}(\bar{p})$ is defined and ${\rm LM}(\pi_{\bm{a}}(\bar{p}))={\rm LM}(\bar{p})$.
+Therefore, $\pi_{\bm{a}}(\ell^{\prime}_{i})$ is defined and we have the following image of the monomial reduction:
+$\pi_{\bm{a}}(\ell^{\prime}_{i})=\pi_{\bm{a}}(\ell_{i})-\pi_{\bm{a}}(b_{t}(\ell_{i}))\frac{t}{{\rm LM}(\bar{p})}\pi_{\bm{a}}(\bar{p})=h_{i}-\pi_{\bm{a}}(b_{t}(\ell_{i}))\frac{t}{{\rm LM}(\pi_{\bm{a}}(\bar{p}))}\pi_{\bm{a}}(\bar{p})$
+As $h_{i}$ is an element in ${\rm GB}(\bm{F})_{d}$, we have ${\rm supp}(\pi_{\bm{a}}(\ell_{i}))\cap W={\rm supp}(h_{i})\cap W=\emptyset$, from which $\pi_{\bm{a}}(b_{t}(\ell_{i}))=0$ and $\pi_{\bm{a}}(\ell^{\prime}_{i})=h_{i}$.
+In this case, if ${\rm supp}(\ell^{\prime}_{i})\cap W$ is still not empty, its greatest monomial is smaller than $t$.
+Replacing $\ell_{i}$ with $\ell^{\prime}_{i}$, we go back to Step 1.
+Since $W$ is a finite set and the greatest monomial of ${\rm supp}(\ell_{i})\cap W$ is decreasing, the procedure terminates in finitely many steps.
+Then, we convert the output polynomial ~$\tilde{\ell}_{i}$ to a polynomial ¯$\bar{\ell}_{i}$ in $K[\mathcal{C}][X]$ by multiplying the least common multiple (LCM) of the denominators of coefficients of ~$\tilde{\ell}_{i}$ to itself.
+Then ¯$\bar{\ell}_{i}$ satisfies the condition of Claim 2.
+Claim 3: For each $i$, the image $\pi_{\bm{a}}(\bar{h}_{i})$ is defined and it coincides with $h_{i}$.
+Let ${\rm vect}(\bar{\ell}_{i})$ be the coefficient vector of ¯$\bar{\ell}_{i}$, where components are indexed by monomials in $\mathcal{T}_{d}$ in decreasing order.
+As ¯$\bar{\ell}_{i}$ does not have any monomial in $W$, we consider a sub-vector ${\rm vect}_{0}(\bar{\ell}_{i})$ which is obtained from ${\rm vect}(\bar{\ell}_{i})$ by eliminating all coefficients of monomials in $W$.
+Let $t_{1}\succ t_{2}\succ\cdots\succ t_{w}$ are indices of vectors.
+(Thus, $\{t_{1},\ldots,t_{w}\}=\mathcal{T}_{d}\smallsetminus W$.) Since ${\rm in}(\langle\bm{F}\rangle)$ is weakly reverse lexicographic, it can be shown easily that $t_{1}={\rm LM}(h_{1})$, …, $t_{v}={\rm LM}(h_{v})$.
+Because, any monomial of degree $d$ not smaller than ${\rm LM}(h_{v})$ should belong to $W\cup\{{\rm LM}(h_{1}),\ldots,{\rm LM}(h_{v})\}$, which implies that ${\rm LM}(h_{1}),\ldots,{\rm LM}(h_{v})$ are top $v$ monomials in $\mathcal{T}_{d}\smallsetminus W$.
+Letting ¯$\bar{\ell}_{i}=a_{i,t_{1}}t_{1}+\cdots+a_{i,t_{v}}t_{v}+a_{i,t_{v+1}}t_{v+1}+\cdots+a_{i,t_{w}}t_{w}$ with $a_{i,t_{j}}\in K[\mathcal{C}]$ for each $i$, a sub-Macaulay matrix $M$ over $K[\mathcal{C}]$ is constructed from vectors ${\rm vect}_{0}(\bar{\ell}_{1}),\ldots,{\rm vect}_{0}(\bar{\ell}_{v})$ as follows:
+$M:=\begin{pmatrix}a_{1,t_{1}}&\ldots&a_{1,t_{v}}&a_{1,t_{v+1}}&\ldots&a_{1,t_{w}}\\ \vdots&\ddots&\vdots&\vdots&&\vdots\\ a_{v,t_{1}}&\ldots&a_{v,t_{v}}&a_{v,t_{v+1}}&\ldots&a_{v,t_{w}}\end{pmatrix}$
+Then, we apply $\pi_{\bm{a}}$ naturally to $M$ by $\pi_{\bm{a}}(M)_{i,t}=\pi_{\bm{a}}(M_{i,t})$ for each $(i,t)$-entry.
+As $\pi_{\bm{a}}(\bar{\ell}_{i})$ is a non-zero constant multiple of $h_{i}$, we have
+$\pi_{\bm{a}}(M)=\begin{pmatrix}b_{1,t_{1}}&0&\ldots&0&b_{1,t_{v+1}}&\ldots&b_{1,t_{w}}\\ 0&{b_{2,t_{2}}}&\ldots&0&{b_{2,t_{v+1}}}&\ldots&{b_{2,t_{w}}}\\ \vdots&\vdots&\ddots&\vdots&\vdots&&\vdots\\ 0&0&\ldots&b_{v,t_{v}}&b_{v,t_{v+1}}&\ldots&b_{v,t_{w}}\end{pmatrix},$
+where $b_{i,t_{j}}=\pi_{\bm{a}}(a_{i,t_{j}})$ for each $i$ and $j$, and $b_{i,t_{i}}\not=0$ for $1\leq i\leq v$.
+Let $M_{0}$ be the square matrix consisting of the first $v$ left columns of $M$ and $\overline{M_{0}}$ the adjugate matrix of $M_{0}$.
+Namely we have
+$M_{0}:=\begin{pmatrix}a_{1,t_{1}}&\ldots&a_{1,t_{v}}\\ \vdots&\ddots&\vdots\\ a_{v,t_{1}}&\ldots&a_{v,t_{v}}\end{pmatrix},$
+and $\overline{M_{0}}$ is a matrix over $K[\mathcal{C}]$.
+Then, $\pi_{\bm{a}}(M_{0})$ is a diagonal matrix with non-zero diagonal entries $b_{1,t_{1}},\ldots,b_{v,t_{v}}$ and $\det(\pi_{\bm{a}}(M_{0}))=\pi_{\bm{a}}(\det(M_{0}))\not=0$.
+Hence we have $\det(M_{0})\not=0$ as an element in $K[\mathcal{C}]$.
+Since $\overline{M_{0}}M_{0}=\det(M_{0})E_{v}$, where $E_{v}$ denotes the identity matrix of degree $v$, we have
+${\color[rgb]{0,0,0}\overline{M_{0}}M=(\det(M_{0})E_{v},\overline{M_{0}}V_{v+1},\ldots,\overline{M_{0}}V_{w}),}$
+where $V_{j}$ denotes the $j$-th column of $M$ for each $j$.
+For each $i$, let $\bar{h}^{\prime}_{i}$ be a polynomial in $K[\mathcal{C}][X]$ converted from the $i$-th low of $\overline{M_{0}}M_{0}$.
+Then, since each ¯$\bar{\ell}_{i}$ belongs to $\langle\bm{G}\rangle\cap K[\mathcal{C}][X]$ and the left-multiplication of $\overline{M}_{0}$ induces low-operations on $M$ over $K[\mathcal{C}]$, each $\bar{h}^{\prime}_{i}$ belongs to $\langle\bm{G}\rangle\cap K[\mathcal{C}][X]$.
+Moreover, we have $t_{i}={\rm LM}(\bar{h}^{\prime}_{i})\in{\rm in}(\langle\bm{G}\rangle)$ and ${\rm LC}(\bar{h}^{\prime}_{i})=\det(M_{0})$.
+Since $t_{i}$ cannot be divided by any monomial in ${\rm LM}({\rm GB}(\bm{F})_{<d})={\rm LM}({\rm GB}(\bm{G}))_{<d})$, the monomials $t_{1},\ldots,t_{v}$ should belong to the minimal generating set ${\rm LM}({\rm GB}(\bm{G}))$ of ${\rm in}(\langle\bm{G}\rangle)$.
+Comparing the number $v=\#{\rm LM}(\bm{G}_{d})$, we have $\{t_{1},\ldots,t_{v}\}={\rm LM}(\bm{G}_{d})$.
+Now let $\hat{h}_{i}=\frac{\bar{h}^{\prime}_{i}}{{\rm LC}(\bar{h}^{\prime}_{i})}=\frac{\bar{h}^{\prime}_{i}}{\det(M_{0})}$ for each $i$.
+Then, we have ${\rm LC}(\hat{h}_{i})=1$ for each $i$.
+Since ${\rm GB}(\bm{G})_{<d}\cup\{\hat{h}_{1},\ldots,\hat{h}_{v}\}$ forms a set of inter-reduced elements, $\{\hat{h}_{1},\ldots,\hat{h}_{v}\}$ coincides with ${\rm GB}(\bm{G})_{d}$, that is, $\bar{h}_{i}=\hat{h}_{i}$ for each $i$.
+Finally we show $\pi_{\bm{a}}(\bar{h}_{i})=h_{i}$ for each $i$.
+Since $\pi_{\bm{a}}(\det(M_{0}))\not=0$, the image $\pi_{\bm{a}}(\bar{h}_{i})$ is defined for each $i$.
+Moreover, ${\rm LM}(\pi_{\bm{a}}(\bar{h}_{i}))=t_{i}={\rm LM}(h_{i})$, ${\rm LC}(\pi_{\bm{a}}(\bar{h}_{i}))=1$, and $\bar{h}_{i}$ is reduced with respect to ${\rm GB}(\bm{G})_{<d}$ and $t_{j}$ for $j\not=i$.
+This implies $\pi_{\bm{a}}(\bar{h}_{i})=h_{i}$ for each $i$ by the uniqueness of the reduced Gröbner basis with respect to a fixed monomial ordering.
+∎
+Finally we consider a generalized cryptographic semi-regular sequence belonging to $Z_{\bm{o}}$, where ${Z}_{\bm{o}}$ is defined in Section 4.1 (see (4.1.1)).
+In this case, the set $\mathcal{C}$ of parameters is $\{c_{j,t}:1\leq j\leq m,\ t\in\mathcal{T}_{d_{j}}\smallsetminus\{x_{n}^{d_{j}}\}\}$.
+Moreover, we set $\mathcal{B}_{d}:=\{(a_{t})_{t\in\mathcal{T}_{d}\smallsetminus\{x_{n}^{d}\}}:a_{t}\in K\}$ for each positive integer $d$, and set $\mathcal{B}_{(d_{1},\ldots,d_{m})}:=\prod_{j=1}^{m}\mathcal{B}_{d_{j}}$.
+Then, for each element $\bm{a}$ in $\mathcal{B}_{(d_{1},\ldots,d_{m})}$, we define a map $\pi_{\bm{a}}$ from $K[\mathcal{C}][X]$ to $K[X]$ by substituting each $c_{j,t}$ with $a_{j,t}$.
+The map $\pi_{\bm{a}}$ is generalized to a map from $K[\mathcal{C}][X]^{m}$ to $K[X]^{m}$, for which we use the same symbol.
+Let $\bm{G}$ be a generic sequence for $Z_{\bm{o}}$ defined in Definition 4.1.7.
+Under this setting, we can apply the same proof in Theorem 4.2.1 to obtain the following results corresponding to Theorem 4.2.1 and Corollary 4.2.2.
+Proposition 4.2.3.
+Let $\bm{G}$ be a generic sequence for $Z_{\bm{o}}$.
+If there exists a generalized cryptographic semi-regular sequence $\bm{F}$ in $Z_{\bm{o}}$, that is, $\bm{F}=\pi_{\bm{a}}(\bm{G})$ for some element $\bm{a}\in\mathcal{B}_{(d_{1},\ldots,d_{m})}$, such that $\dim_{K}(K[X]/\langle\bm{F}\rangle)_{D}=1$ for $D=\widetilde{d}_{\rm reg}(\langle\bm{F}\rangle)<\infty$ and ${\rm in}(\langle\bm{F}\rangle)$ is weakly reverse lexicographic, then ${\rm GB}(\bm{F})=\pi_{\bm{a}}({\rm GB}(\bm{G}))$ and ${\rm LM}({\rm GB}(\bm{F}))={\rm LM}({\rm GB}(\bm{G}))$.
+In this case, $\bm{G}$ is generalized cryptographic semi-regular with $\dim_{L}(L[X]/\langle\bm{G}\rangle)_{D}=1$ for $D=\widetilde{d}_{\rm reg}(\langle\bm{G}\rangle)<\infty$ and ${\rm in}(\langle\bm{G}\rangle)$ is weakly reverse lexicographic.
+Moreover, in this case, there exists a non-empty Zariski-open set $U$ of $Z_{\bm{o}}$ with respect to the induced topology such that $\bm{H}$ is generalized cryptographic semi-regular and ${\rm in}(\langle\bm{H}\rangle)$ is weakly reverse lexicographic for any $\bm{H}$ in $U$.
+Proof.
+By Proposition 4.1.9, we have
+${\rm HS}_{K[X]/\langle\bm{F}\rangle}(z)\equiv{\rm HS}_{L[X]/\langle\bm{G}\rangle}(z)\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{D}},$
+where $D=\widetilde{d}_{\rm reg}(\langle\bm{F}\rangle)=\widetilde{d}_{\rm reg}(\langle\bm{G}\rangle)$ and $\dim_{L}(L[X]/\langle\bm{G}\rangle)_{D}=1$.
+In this case, as $\bm{F}$ and $\bm{G}$ have $(0:\cdots:0:1)$ as their projective zero, $x_{n}^{D}$ cannot belong neither to $\langle\bm{F}\rangle_{D}$ nor to $\langle\bm{G}\rangle_{D}$.
+This implies that $(K[X]/\langle\bm{F}\rangle)_{D}$ and $(L[X]/\langle\bm{G}\rangle)_{D}$ have a common linear basis $\{x_{n}^{D}\}$, which also implies that ${\rm max.GB.deg}(\langle\bm{F}\rangle):={\rm max}\{\deg(h):h\in{\rm GB}(\bm{F})\}$ and ${\rm max.GB.deg}(\langle\bm{G}\rangle)$ are both smaller than $D$.
+Thus, all arguments in the proof of Theorem 4.2.1 can be applied to the case of Proposition 4.2.3, from which we obtain ${\rm GB}(\bm{F})=\pi_{\bm{a}}({\rm GB}(\bm{G}))$ and ${\rm LM}({\rm GB}(\bm{F}))={\rm LM}({\rm GB}(\bm{G}))$.
+Then, by applying Theorem 2.3.4 to $\bm{G}$, there exists a non-empty Zariski-open set $U_{\mathcal{B}}$ of $\mathcal{B}_{(d_{1},\ldots,d_{m})}$ such that ${\rm LM}({\rm GB}(\bm{F}))={\rm LM}({\rm GB}(\pi_{\bm{b}}(\bm{G})))$ for any $\bm{b}$ in $U_{B}$, that is, ${\rm in}(\langle\bm{F}\rangle)={\rm in}(\langle\pi_{\bm{b}}(\bm{G})\rangle))$.
+This implies that the subset $U$ of $Z_{\bm{o}}$ corresponding to $U_{\mathcal{B}}$ is also a non-empty Zariski-open set such that $\bm{H}$ is generalized cryptographic semi-regular and ${\rm in}(\langle\bm{H}\rangle)$ is weakly reverse lexicographic for any $\bm{H}\in U_{\mathcal{B}}$.
+∎
+Proposition 4.2.4.
+Suppose that the conditions of Proposition 4.2.3 are satisfied.
+Then, for any $\bm{p}\in\mathbb{P}^{n-1}(K)$, there exists a non-empty Zariski-open set $U_{\bm{p}}$ of $Z_{\bm{p}}$ with respect to the induced topology such that $\bm{H}$ is generalized cryptographic semi-regular and ${\rm in}(\langle\bm{H}\rangle)$ is weakly reverse lexicographic for any $\bm{H}$ in $U_{\bm{p}}$.
+Proof.
+We show the assertion by using the arguments in the proof of Lemma 4.1.6 Let $\bm{p}:=(p_{1}:\ldots:p_{n})$ be a point in $\mathbb{P}^{n-1}(K)$.
+Without loss of generality, we may assume that $p_{n}\not=0$ and moreover $p_{n}=1$, and consider the following maps, where $\mathcal{V}:=V_{n,m,(d_{1},\ldots,d_{m})}$:
+$\phi_{\bm{p}}:(a_{1}:\cdots:a_{n-1}:a_{n})\mapsto(a_{1}-p_{1}a_{n}:\cdots:a_{n-1}-p_{n-1}a_{n}:a_{n}),$
+$\phi_{\bm{p}}^{*}:K[X]\ni h(x_{1},x_{2},\ldots,x_{n})\mapsto h(x_{1}-p_{1}x_{n},\ldots,x_{n-1}-p_{n-1}x_{n},x_{n})\in K[X],$
+$\Phi_{\bm{p}}:\mathcal{V}\in(h_{1},\ldots,h_{m})\mapsto(\phi_{\bm{p}}^{*}(h_{1}),\ldots,\phi_{\bm{p}}^{*}(h_{m}))\in\mathcal{V}.$
+Then $\phi_{\bm{p}}^{*}$ is a ring homomorphism and $\Phi_{\bm{p}}$ is an invertible linear map on $\mathcal{V}$ and its restriction on $Z_{\bm{o}}$ induces a bijection between $Z_{\bm{o}}$ and $Z_{\bm{p}}$.
+We also let $\prec$ be the graded reverse lexicographical ordering with $x_{1}\succ x_{2}\succ\cdots\succ x_{n}$.
+Now we suppose that the conditions of Proposition 4.2.3 hold.
+Then the generic sequence $\bm{G}$ for $Z_{\bm{o}}$ is generalized cryptographic semi-regular sequence with $\dim_{L}(L[X]/\langle\bm{G}\rangle)_{D}=1$ for $D=\widetilde{d}_{\rm reg}(\langle\bm{G}\rangle)<\infty$ and ${\rm in}(\langle\bm{G}\rangle)$ is weakly reverse lexicographic, and there is a non-empty Zariski-open set $U_{\bm{o}}$ in $Z_{\bm{o}}$ such that, for any $\bm{H}$ in $U_{\bm{o}}$, $\bm{H}$ is generalized cryptographic semi-regular and ${\rm in}(\langle\bm{H}\rangle)$ is weakly reverse lexicographic.
+Meanwhile, since $\phi_{\bm{p}}^{*}$ does not change LM for any polynomial in $L[X]$ and ${\rm in}(\langle\bm{F}\rangle)={\rm in}(\langle\Phi_{\bm{p}}(\bm{F})\rangle)$ for any $\bm{F}$ in $U_{\bm{o}}$, it can be shown easily that, for any $\bm{H}$ in $U_{\bm{o}}$, $\Phi_{\bm{p}}(\bm{H})$ is generalized cryptographic semi-regular and ${\rm in}(\langle\Phi_{\bm{p}}(\bm{H})\rangle)$ is weakly reverse lexicographic.
+Let $U_{\bm{p}}=\Phi_{\bm{p}}(U_{\bm{o}})$.
+Then, $U_{\bm{p}}$ is also a non-empty Zariski-open set of $\Phi_{\bm{p}}(Z_{\bm{o}})=Z_{\bm{p}}$ (see the proof of Lemma 4.1.6).
+Thus, $U_{\bm{p}}$ satisfies the assertion of Proposition 4.2.4.
+∎
+Remark 4.2.5.
+It is clear that the converse of Proposition 4.2.4 also holds, that is, if there is a generalized cryptographic semi-regular sequence $\bm{H}$ in $Z_{\bm{p}}$ such that $\dim_{K}(K[X]/\langle\bm{H}\rangle)_{D}=1$ for $D=\widetilde{d}_{\rm reg}(\langle\bm{H}\rangle)<\infty$ and ${\rm in}(\langle\bm{H}\rangle)$ is weakly reverse lexicographic, then the sequence $\bm{F}=(\Phi_{\bm{p}})^{-1}(\bm{H})$ belongs to $Z_{\bm{o}}$ and it satisfies the condition of Proposition 4.2.3.
 Acknowledgement